### v2

# Recent Advances in OOD Detection: Problems and Approaches

Abstract.
Out-of-distribution (OOD) detection aims to detect test samples outside the training category space, which is an essential component in building reliable machine learning systems. Existing reviews on OOD detection primarily focus on method taxonomy, surveying the field by categorizing various approaches. However, many recent works concentrate on non-traditional OOD detection scenarios, such as test-time adaptation, multi-modal data sources and other novel contexts. In this survey, we uniquely review recent advances in OOD detection from the problem scenario perspective for the first time. According to whether the training process is completely controlled, we divide OOD detection methods into training-driven and training-agnostic. Besides, considering the rapid development of pre-trained models, large pre-trained model-based OOD detection is also regarded as an important category and discussed separately. Furthermore, we provide a discussion of the evaluation scenarios, a variety of applications, and several future research directions. We believe this survey with new taxonomy will benefit the proposal of new methods and the expansion of more practical scenarios. A curated list of related papers is provided in the Github repository: https://github.com/shuolucs/Awesome-Out-Of-Distribution-Detection
Trustworthy Machine Learning, Out-of-distribution Detection
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Trustworthy Machine Learning Out-of-distribution Detection
1. Introduction
Machine learning methods  (LeCun et al., 2015) have made significant progress under the closed-world assumption, where test data is drawn from the same distribution as the training set, known as in-distribution (ID). However, in the real world, models inevitably encounter test samples that do not belong to any training set category, commonly referred to as out-of-distribution (OOD) data. OOD detection  (Hendrycks and Gimpel, 2017) aims to identify and reject OOD samples rather than make overconfident predictions arbitrarily  (Nguyen et al., 2015) while maintaining accurate classification for ID data. Models with superior OOD detection capabilities are more reliable and have important applications in numerous security-critical scenarios. For instance, in medical diagnosis systems, a model that cannot detect OOD samples will misjudge unknown diseases and cause serious misdiagnosis  (Schlegl et al., 2017). Similarly, autonomous driving algorithms  (Geiger et al., 2012) should detect unknown scenarios and resort to human control to avoid accidents caused by arbitrary judgment.
Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.  Yang et al. (2021) discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based, and reconstruction-based methods.  Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods. Additionally,  Lang et al. (2023) offer a review of OOD detection methods in natural language processing. However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios. Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method. Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022), there is an urgent need for a survey that incorporates the latest technologies.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=13em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ OOD Detection ,fill=hidden-yellow!70,ver [ Problem: Training-driven
OOD detection(§ 3),fill=hidden-red!70, align=center, text width=12em [ Approaches with only
ID Data (§ 3.1),fill=hidden-blue!70,align=center [ Reconstruction-based, fill=hidden-blue!70 [ MoodCat  (Yang et al., 2022a), RAE (Zhou, 2022), MOOD  (Li et al., 2023a), MOODv2  (Li et al., 2024a), PRE  (Osada et al., 2023), LMD  (Liu et al., 2023a) DiffGuard  (Gao et al., 2023a),
DenoDiff (Graham et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Probability-based, fill=hidden-blue!70 [ HVCM (Li et al., 2023b), DDR (Huang et al., 2022), LID (Kamkari et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Logits-based, fill=hidden-blue!70 [ LogitNorm  (Wei et al., 2022a), UE-NL  (Huang and Qiao, 2023), DML  (Zhang and Xiang, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ OOD Synthesis, fill=hidden-blue!70 [ ConfiCali (Lee et al., 2018a), CODEs  (Tang et al., 2021), CMG  (Wang et al., 2022a), VOS  (Du et al., 2022a), NPOS  (Tao et al., 2023), SHIFT  (Kwon et al., 2023), ATOL  (Zheng et al., 2023),
SSOD  (Pei, 2024), SEM  (Yang et al., 2023b) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Prototype-based, fill=hidden-blue!70 [ SIREN  (Du et al., 2022b), CIDER  (Ming et al., 2023), PALM  (Lu et al., 2024), ReweightOOD  (Regmi et al., 2024), Step  (Zhou et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ A special case:
Long-tail ID data, fill=hidden-blue!70 [ OLTR  (Liu et al., 2019), II-Mixup (Mehta et al., 2022), AREO  (Sapkota and Yu, 2023), IDCP (Jiang et al., 2023a), Open-Sampling  (Wei et al., 2022b), COOD  (Hogeweg et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Approaches with both ID
and OOD Data (§ 3.2),fill=hidden-blue!70,align=center [ Boundary Regularization,fill=hidden-blue!70 [ OE  (Hendrycks et al., 2019), ELOC (Vyas et al., 2018), Why-ReLU (Hein et al., 2019), SSL-GOOD (Mohseni et al., 2020), EnergyOE  (Liu et al., 2020), MixOE  (Zhang et al., 2023a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Outlier Mining,fill=hidden-blue!70 [ BD-Resamp (Li and Vasconcelos, 2020), POEM  (Ming et al., 2022a), DAOL  (Wang et al., 2023a), DOE (Wang et al., 2023b), MixOE  (Zhang et al., 2023a), DivOE  (Zhu et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Imbalanced ID,fill=hidden-blue!70 [ PASCAL  (Wang et al., 2022b), COCL  (Miao et al., 2024), BERL (Choi et al., 2023), EAT  (Wei et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Training-agnostic
OOD detection (§ 4),fill=hidden-red!70,align=center,text width=12em [ Post-hoc Approaches (§ 4.1),fill=hidden-blue!70,align=center [ Output-based,fill=hidden-blue!70 [ MSP  (Hendrycks and Gimpel, 2017), MaxLogits  (Hendrycks et al., 2022), Energy  (Liu et al., 2020), GEN  (Liu et al., 2023b), ZODE  (Xue et al., 2024), LogicOOD  (Kirchheim et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Distance-based,fill=hidden-blue!70 [ Mahalanobis  (Lee et al., 2018b), NNGuide  (Park et al., 2023a), KNN  (Sun et al., 2022), SSD  (Sehwag et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Gradient-based,fill=hidden-blue!70 [ Grad  (Lee and AlRegib, 2020), GradNorm  (Huang et al., 2021), GradOrth  (Behpour et al., 2023), GAIA  (Chen et al., 2023), OPNP  (Chen et al., 2024a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Feature-based,fill=hidden-blue!70 [ ODIN  (Liang et al., 2018a), ReAct  (Sun et al., 2021), VRA  (Xu et al., 2023), SHE  (Zhang et al., 2022), ViM  (Wang et al., 2022c), Neco  (Ammar et al., 2024), ASH  (Djurisic et al., 2023), NAC  (Liu et al., 2024a)
Optimal-FS (Zhao et al., 2024), BLOOD  (Jelenić et al., 2024), SCALE  (Xu et al., 2024), DDCS  (Yuan et al., 2024), LINe  (Yong Hyun Ahn, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Density-based,fill=hidden-blue!70 [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Test-time Adaptive Approaches
(§ 4.2) ,fill=hidden-blue!70,align=center [ Model-update needed,fill=hidden-blue!70 [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023b), SAL  (Du et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Model-update free,fill=hidden-blue!70 [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024), RTL  (Fan et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Large pre-trained
model-based OOD detection
(§ 5),fill=hidden-red!70,align=center,text width=12em [ Zero-shot Approaches
(§ 5.1),fill=hidden-blue!70,align=center [ VLM-based,fill=hidden-blue!70 [ CLIPScope  (Fu et al., 2024), RONIN  (Nguyen et al., 2024), ZOC  (Esmaeilpour et al., 2022), MCM  (Ming et al., 2022b), CLIPN  (Wang et al., 2023c), NegLabel  (Jiang et al., 2023b),
NegPrompt-C (Nie et al., 2024), LAPT  (Zhang et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ LLM-based,fill=hidden-blue!70 [ WK-LLM (Dai et al., 2023), ODPC  (Huang et al., 2024), LLM-OOD (Liu et al., 2024b) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Few-shot Approaches
(§ 5.2),fill=hidden-blue!70,align=center [ Fine-tuning based, fill=hidden-blue!70 [ CLIP-OS  (Sun et al., 2024), EOK-Prompt  (Zeng et al., 2024), GalLop  (Lafon et al., 2024), ID-like  (Bai et al., 2024), DSFG  (Dong et al., 2023), LoCoOp  (Miyai et al., 2023a),
NegPrompt  (Li et al., 2024b), TagOOD  (Li et al., 2024c), CRoFT (Zhu et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Meta-learning based, fill=hidden-blue!70 [ OOD-MAML  (Jeong and Kim, 2020), HyperMix  (Mehta et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Full-shot Approaches
(§ 5.3),fill=hidden-blue!70,align=center [ NPOS  (Tao et al., 2023), PT-OOD (Miyai et al., 2023b), TOE  (Park et al., 2023b) , leaf, text width=48.5em,fill=hidden-blue!70,text width=50.6em ] ] ] ]
Figure 1. Taxonomy of OOD detection problem scenarios and solutions.
In this survey, we for the first time review the recent advances in OOD detection with a problem-oriented taxonomy, as illustrated in Fig. 1. Based on whether the method needs to control the pre-training process, we categorize OOD detection algorithms into training-driven and training-agnostic methods. Considering the rapid development of large pre-trained models nowadays, we also regard large pre-trained model-based OOD detection as a separate section. In detail, training-driven methods achieve high detection capability by designing the optimization process of the training stage. They are further classified and discussed according to whether OOD data is used in training. Training-agnostic methods distinguish OOD data from ID ones based on a well-trained model, skipping the time-consuming and expensive pre-training process in practice. According to whether utilizing test samples to further improve OOD detection performance, we categorize them into post-hoc and test-time methods. Large pre-trained model-based OOD detection methods focus on models such as vision language models or large language models, which are pre-trained on vast datasets and excel in a wide array of tasks. We discuss them in terms of whether they have access to a few examples, including zero-shot, few-shot and full-shot scenarios.
The remainder of this survey is organized as follows. We recap the related work of OOD detection in Sec. 2. Next, we summarize training-driven OOD detection approaches in Sec. 3, and introduce the training-agnostic OOD detection methods in Sec. 4. Then, in Sec. 5, we introduce large pre-trained model-based OOD detection. An overview of the evaluation metrics, experimental protocols, and applications is presented in Sec. 6. Following that, we discuss promising trends and open challenges in Sec. 7 to shed light on underexplored and potentially critical avenues. Finally, we conclude this survey in Sec. 8.
2. Related work
Anomaly Detection. Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009). These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003), or system failures  (Du et al., 2017). The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations. Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency. While both AD and OOD detection aim to identify unusual or unexpected data points, AD focuses on deviations within a single distribution (typically the training set), whereas OOD detection addresses differences between the training distribution and new, unseen inputs.
Novelty Detection. Novelty detection (ND) focuses on identifying new or unknown data points that a model has not seen during training  (Pimentel et al., 2014). This process is essential in situations where the system needs to adapt to evolving conditions or when it’s crucial to flag new, unseen scenarios. Unlike AD, which seeks to find patterns that deviate from the norm, novelty detection aims to discover entirely new patterns. Applications include identifying new trends in social media  (Li et al., 2017), discovering new species in biological data  (Pimentel et al., 2014), or detecting new topics in document streams  (Aksoy et al., 2012). Essentially, ND deals with surprises in familiar contexts, while OOD detection deals with data from unfamiliar contexts.
Open Set Recognition. Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016). This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that were not present during training. Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully. OSR is concerned with identifying new classes that are not present during training, which is generally achieved by dividing the categories of the same dataset into base classes and new classes, which means that new classes and base classes usually come from the same domain. In contrast, OOD detection focuses on identifying any data that is different from the training distribution, regardless of whether it belongs to a new class or a different domain altogether. Essentially, OSR is a subset of OOD detection, specifically aimed at classifying new class types within the same domain.
Outlier Detection. Outlier detection (OD) is the process of identifying data points that are significantly different from the majority of the data  (Singh and Upadhyaya, 2012). It is similar to AD but focuses more on the identification of individual data points rather than patterns. Outliers can arise due to variability in measurement, experimental errors, or genuinely novel variations. Applications include fraud detection, fault detection, and removing anomalous data from datasets to improve model accuracy  (Pawar et al., 2014). Compared to OOD detection, OD is more of a transductive scenario, as it inherently has access to outliers. However, classical OOD detection only encounters outliers at the time of deployment.
Zero-shot learning. Zero-shot learning is a paradigm in machine learning where the goal is to recognize objects without having seen examples of these objects during the training phase  (Romera-Paredes and Torr, 2015; Wang et al., 2019; Pourpanah et al., 2022). The fundamental challenge in zero-shot learning is to effectively transfer knowledge from seen to unseen classes  (Romera-Paredes and Torr, 2015). It has been primarily tackled by learning semantic relationships between classes through attributes or by embedding both seen and unseen classes into a shared semantic space. While both deal with unknowns during inference, zero-shot learning tries to classify new categories  (Xian et al., 2018), whereas OOD detection flags data that is anomalous or unfamiliar to the training data distribution  (Lee et al., 2018b).
Selective Classification. Selective classification, also known as reject option classification, provides a mechanism for models to abstain from making a decision when they are not sufficiently confident in their predictions  (Geifman and El-Yaniv, 2017; Chang et al., 2024). Selective classification involves a model deciding when to make a prediction based on its confidence level, effectively choosing not to predict when uncertain  (Wiener and El-Yaniv, 2011). On the other hand, OOD detection identifies data points that differ from the training distribution, aiming to flag them as unfamiliar to the model. While both methods manage uncertainty, selective classification deals with confidence in making predictions on ID data (El-Yaniv et al., 2010; Chang et al., 2024), and OOD detection focuses on recognizing and handling data that is not represented in the training set.
3. Problem: Training-driven OOD Detection
In the training-driven OOD detection problem, researchers design the pre-training process to obtain models with superior OOD detection capabilities. Based on whether OOD data is accessible during training, we further divide methods under this scenario into two folds: training with only ID data and training with both ID and OOD data, as shown in Fig. 2.
3.1. OOD Detection Approaches with only ID Data
Overview. Given the ID data, approaches in this section train a model on them and aim to utilize the model for detecting OOD test samples while ensuring accurate classification of the ID data. Approaches with only ID data focus specifically on mining information from ID data, without explicitly relying on other information from real-world OOD data. We further differentiate these methods into the following four categories: Reconstruction-based, Probability-based, Logits-based, OOD Synthesis, and Prototype-based. Considering real-world requirements, we also delve into a specific scenario: Long-tail ID data.
Figure 2. Illustration of training-driven OOD detection approaches. Dashed borders indicate that they are not used in the specific phase. OOD images are excluded in the “Approaches with only ID data” on the left but are included in the “Approaches with both ID and OOD data” on the right. In both cases, OOD labels are not utilized.
Reconstruction-based. The reconstruction-based methodology offers a new research avenue by scrutinizing the discrepancy between sample representations before and after reconstruction, which relies heavily on the reconstruction performance of models. Fundamentally, the objective of the reconstruction task is to recuperate the inherent semantic content of the dataset, guided by a designated supervisory signal. This premise rests on the notion that OOD data intrinsically possess semantic characteristics that are incongruent with those of ID label information. By quantitatively assessing the extent of this semantic deviation, the model becomes equipped to discern OOD data with a high degree of precision. The approximate process of such methods is shown in Fig.  3.
MoodCat  (Yang et al., 2022a) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process. Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data. Following this, MOOD  (Li et al., 2023a) and MOODv2  (Li et al., 2024a) leverage pretraining tasks based on masked image modeling, exhibiting significant advantages in learning the internal distribution of data. PRE  (Osada et al., 2023) introduces the concept of normalized flow, combined with a penalty based on typicality to constrain reconstruction errors, which discerns differences between OOD and ID data well.
Figure 3. Illustration of reconstruction-based OOD detection approaches. After extracting the feature representation $F(x)$ of the image through a neural network, it is then fed into a reconstruction network (such as VAE or DDPM) to obtain the reconstructed image representation $F^{\prime}(x)$. By comparing the dissimilarity between the original and reconstructed images, we can identify OOD data, as OOD samples typically exhibit greater dissimilarity before and after reconstruction.
Recently, diffusion models have made remarkable strides in both training stability and the quality of generated images. Leveraging diffusion models to detect OOD data has become a new research direction.  Graham et al. (2023) address this issue by introducing DDPM, utilizing it for the reconstruction of noise-disturbed images. The model employs multidimensional reconstruction error for the identification of OOD data. The information bottleneck of this model can be externally regulated. Similarly, LMD  (Liu et al., 2023a) utilizes diffusion models for OOD validation. LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold. DiffGuard  (Gao et al., 2023a) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
Probability-based. Research in probability-based direction aims to establish probability models to describe the distribution of training data. Through these probability models, suitable scoring functions are developed to compute the scores of test samples within the ID distribution, which reflect whether these samples belong to the ID distribution. In this domain,  Li et al. (2023b) model each ID category during training using multiple Gaussian mixture models, and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.  Huang et al. (2022) address this issue by introducing two regularization constraints. The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples. Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models. It measures whether data should be classified as in-distribution by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
Logits-based. These algorithms primarily focus on the predictions of neural networks, particularly on Logits, which are the results of the neural network’s output layer. Logits typically represent the model’s confidence or probability for each category. LogitNorm  (Wei et al., 2022a) proposes a method using logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence. Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores. It adjusts the learning intensity between samples during training, rendering the learning process robust. DML  (Zhang and Xiang, 2023) challenges the potential hindrance to OOD detection performance. Drawing on empirical insights, DML enhances OOD detection performance by decoupling logits and balancing individual components, thereby mitigating the impact of each attribute on the results. It enhances OOD detection performance by decoupling logits and scenario MaxNorm as a parameter, thereby balancing the influence of each attribute on the outcome.
OOD Synthesis. In the task of OOD detection, it is theoretically believed that incorporating the features of OOD data during model training can enhance the OOD detection performance. Due to the challenges in acquiring distribution information for OOD samples, some methods employ ID data to estimate the distribution of OOD data. This is done to simulate the scenario in which a model encounters OOD data in real-world situations.
Figure 4. Illustration of OOD synthesis approaches. In the absence of real-world OOD data, utilizing extrapolation from ID data to identify suitable OOD data and then training the model using the method of exposing anomalies enhances the capability of OOD detection. The orange dots in the figure represent virtual OOD data synthesized using ID data.
Lee et al. (2018a) add additional regularization terms by incorporating two supplementary terms into the original loss function. The first term compels the classifier to increase uncertainty for OOD samples, while the second term encourages the GAN to generate appropriate OOD samples. Specifically, a joint training scheme is designed to select suitable boundary samples in the low-density regions of ID samples. Similarly, VOS  (Du et al., 2022a) employs adaptive synthesis of anomaly values. It samples anomaly values from the low-likelihood regions of class-conditional distributions in the feature space, eliminating the need for reliance on external data. This enables the model to synthesize OOD data for predictions. It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention. Noticing the additional distributional assumptions posed by previous methods for OOD data, Some researches approach the problem from another perspective. For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality. it selects boundary points through non-parametric density estimation based on nearest neighbors, with these boundary points situated between ID and OOD data. Additionally, SSOD  (Pei, 2024) introduces self-supervised sampling to implicitly generate OOD data. It directly samples natural OOD signals from the backgrounds of ID data images, thus overcoming limitations introduced by biases during the OOD synthesis stage.
In contrast to the implicit generation methods of OOD samples mentioned above, some approaches take a different route by directly leveraging explicit construction of OOD samples from ID samples for learning. To acquire OOD samples, CODEs  (Tang et al., 2021) initially generate seed OOD examples by slicing and stitching samples from internal distributions of different categories. These examples are then inputted into Chamfer GAN for distribution transformation, yielding high-quality OOD samples. CMG  (Wang et al., 2022a) then generates pseudo OOD data by providing class embeddings mixed as conditions to a Conditional Variational Autoencoder (CVAE), and subsequently utilizes this data to fine-tune the classifier for OOD detection. SEM  (Yang et al., 2023b) introduces a problem setting for full-spectrum OOD detection, where negative samples are generated using Mixup operations during training. It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data. SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples. It achieves this by employing CLIP to eliminate ID object regions within the training samples. The latent diffusion model is then utilized to substitute these regions with authentic features, all the while taking into account the contextual background. This methodology thereby establishes the model’s capability for rejection. However, ensuring the quality of generated data is often challenging. Furthermore, the quality of the generated data itself may have inherent flaws. ATOL  (Zheng et al., 2023) introduces an auxiliary task, wherein both auxiliary ID and auxiliary OOD data coexist. In a low-dimensional latent space, distinct regions are manually sought for both auxiliary ID and auxiliary OOD data, ensuring non-overlapping characteristics in the input space. Subsequently, data generation by the generator guarantees the non-overlapping properties of the input space. Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances.
Prototype-based. During the model training process, prototype-based OOD detection methods aim to model the ID data using prototypes to learn the common distribution characteristics of the ID data. In the testing phase, the model measures the differences between the sample and class-level prototypes to determine the category of the sample. The general procedure of these methods is roughly illustrated in Fig.  5.
Figure 5. Illustration of prototype-based OOD detection approaches. The algorithm for OOD detection based on prototype learning roughly consists of three network structures: 1) Encoder. This structure is responsible for propagating sample data through a neural network to obtain feature representations $h$. 2) Projector. Structures of this type project the image features encoded in $h$ to a new representation space. 3) Prototype Learning. In the new representation space, prototypes representing each class are learned, with each class learning its corresponding prototype center. During testing, the likelihood of anomalous data is measured by comparing the similarity between test samples and prototypes of each ID class.
SIREN  (Du et al., 2022b) first models the distribution of ID data using the von Mises-Fisher(vMF) model, which enables representing each class as a compact cluster, i.e., a class prototype. Generally, the vMF distribution modeling formula can be represented as: $p_{D}(z;p_{k},k)=Z_{D}(k)exp(kp_{k}^{T}z)$, where $p_{k}$ is the $k$-th prototype with unit norm, $\kappa\geq 0$ represents the concentration around the mean, and $Z_{D}(\kappa)$ is the normalization factor. In the prototype-based approach, an embedding vector $z$ is assigned to class $c$ with the following normalized probability:
(1) $p(y=c|z;(P_{j},k_{j})_{j=1}^{C})=\frac{Z_{D}(k_{c})exp(k_{c}\mu_{c}^{T}z)}{{\textstyle\sum_{j=1}^{C}Z_{d}(k_{j})exp(k_{j}\mu_{j}^{T}z)}},$
where $c\in{\{1,2,...C\}}$. Additionally, within the loss function, it enforces alignment between the embedding vectors of ID samples and class prototypes, constraining each ID class sample. This parametric OOD score can be directly obtained after training, without requiring separate estimation. CIDER  (Ming et al., 2023) builds upon SIREN by jointly optimizing two losses to enhance data discriminability, encouraging the maximization of angular distances between prototypes of different classes and the internal compactness of prototypes within the same class. The optimization process during model training pertains to the prototypes of classes. ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores. To address this, they propose a re-weighting optimization strategy to balance the significance of different losses. Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning.
However, PALM  (Lu et al., 2024) observes that modeling each class with a single prototype may fail to capture the diversity of internal information within the data. PALM introduces a mixed-prototype strategy to optimize the modeling process, utilizing a strategy of multiple prototypes to learn informative representations for each class. By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function.
A special case: Long-tail ID data. ID data in real scenarios may present a long-tailed distribution due to the difficulty of collection and frequency of occurrence. This imbalance will strongly affect the performance of OOD detection. Many approaches are proposed to address the challenge of ID imbalance and enhance OOD detection capabilities.
OLTR  (Liu et al., 2019) addresses the robustness of tail recognition by associating visual concepts between the head and tail embeddings. It accomplishes this by dynamically adjusting embedding norms through the utilization of visual memory. The approach involves two components to enhance the robustness of long-tailed data: one utilizes direct features from standard embeddings, while the other employs memory features that store discriminative centroids derived from direct features, thereby augmenting the direct features of the images.  Mehta et al. (2022) propose a tailored complex subset mixing strategy designed specifically for handling medium-to-tail classes. This method creates mixed samples by pairing two independent samples from different tail categories and calculates the distance between the mixed sample and prototypes of specific classes. The mixing strategy is ultimately combined with prototype learning to effectively address the challenges posed by long-tail scenarios. AREO  (Sapkota and Yu, 2023) introduces a method to quantify sample uncertainty through evidential learning, akin to confidence. The entire training process is governed by an innovative multi-scheduler learning mechanism, dynamically adjusting training parameters based on the importance of different classes to ensure the model focuses on the features of minority classes, thus balancing the disparities between majority and minority classes. Overall, AREO identifies higher uncertainty in certain samples during training and dynamically adjusts parameters to mitigate overfitting to these samples.
Existing OOD detection methods predominantly assume a uniform distribution of probabilities from OOD to ID. In the presence of class imbalances,  Jiang et al. (2023a) propose an alternative strategy to enhance current OOD detection methodologies. This involves recalibrating the OOD score based on the class prior distribution of the ID and the KL divergence from the output of a pre-trained model. This approach signifies the potential to augment the robustness of OOD detection beyond the achievements of previously optimized methodologies, further fortifying the overall performance. Additionally, Open-Sampling  (Wei et al., 2022b) leverages noisy labels from the OOD dataset to rebalance the class priors of the ID training dataset. These labels are sampled from a predefined distribution that complements the original class prior distribution. COOD  (Hogeweg et al., 2024) uses the supervised model to combine individual OOD measures into a single ensemble, similar to the concept of random forests. This approach addresses the limitations of individual OOD methods and can also overcome issues related to data imbalance.
3.2. OOD Detection Approaches with Both ID and OOD Data
Overview. In some known deployment scenarios, real OOD data can be easily collected at a low cost. Some methods based on this assumption focus on how to use OOD data for better detection performance. Differing from methods involving OOD Synthesis, in these research directions, models have access to real-world OOD data during the training phase. The primary focus in such problems is on optimizing the model’s decision boundary, rather than the OOD data itself. Due to the introduction of real OOD information, the boundary of these two categories will be accurately calculated.
Boundary Regularization. The Boundary Regularization class of methods belongs to the traditional Outlier Exposure (OE) approaches. The central idea of  Hendrycks et al. (2019) and  Hein et al. (2019) are to fully leverage OOD data to optimize the model’s decision boundary, thus achieving OOD detection. Proponents of this concept can utilize auxiliary anomaly datasets to enhance the OOD detector, enabling it to generalize and detect anomalous information not encountered during training. The central idea of this method can be grasped from Fig.  6.
Figure 6. Illustration of boundary regularization approaches. The key to model optimization strategy lies in the decision boundary of the model classification. Boundary regularization utilizes existing known ID data (the colored dots) and some OOD data (black dots) to optimize the potential decision boundary, compacting its classification boundary with abnormal data as tightly as possible without affecting the classification of ID data.
Specifically, given a model $f$ and the original loss function, the model training process aims to minimize the objective:
(2) $\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{in}}}\left[\mathcal{L}(f(x),y)+\lambda\mathbb{E}_{x^{\prime}\sim\mathcal{D}_{\text{out}}^{\text{OE}}}\left[\mathcal{L}_{\text{OE}}(U,f(x))\right]\right],$
over the parameters of $f$, where $x^{\prime}$ represents auxiliary anomaly data and $U$ denotes the uniform distribution. $\mathcal{L}_{\text{OE}}$ represents the cross-entropy loss with respect to $U$. The fundamental purpose is to compel the model to optimize the OOD data distribution to a uniform distribution, a principle that is universal in OE-type approaches. The specific design of $\mathcal{L}_{\text{OE}}$ can depend on other task requirements and the chosen OOD score. This design can utilize a maximum softmax probability baseline  (Hendrycks and Gimpel, 2017) detector to detect anomalous data. Compared to traditional softmax scores, EnergyOE  (Liu et al., 2020) builds upon OE by leveraging energy scores for better discrimination between ID and OOD samples, and it is less prone to issues of overconfidence. Specifically, its calculation formula:
(3) $E(x;f)=-T\cdot log\displaystyle\sum_{i}^{K}{e}^{{f}_{i}(x)/T},$
where the temperature coefficient $T$ is used and $f(x)$ denotes the discriminative neural classifier $f(x):\mathbb{R}^{D}\rightarrow\mathbb{R}^{K}$, which maps an input $x\in\mathbb{R}^{D}$ to $K$ real-valued logits.
Mohseni et al. (2020) train a model using a self-supervised approach, optimizing the objective function for unlabeled OOD samples using pseudo-labeling to generalize OOD detection capabilities.  Vyas et al. (2018) similarly employ self-supervised training of the classifier. Unlike the OE approach, its aim is to find a gap between the average entropy of OOD and ID samples. MixOE  (Zhang et al., 2023a) takes into account the beneficial effect of subtle OOD samples on enhancing the generalization ability of OOD detection. Its main idea is to mix ID and OOD data samples to broaden the generalization of OOD data. Training the model with these outliers can linearly decrease the prediction confidence with inputs from ID to OOD samples, explicitly optimizing the generalization ability of the decision maker.
Outlier Mining. The traditional OE concept assumes the existence of ID input $\mathcal{D}_{\text{in}}$ and OOD input $\mathcal{D}_{\text{out}}$, both independently and heterogeneously distributed, originating from different sources. However, this premise cannot be fully guaranteed in the current training process due to potential noise in the training OOD data. Outlier Mining differs slightly from the traditional OE approach in that, although it also utilizes real-world OOD samples to address the issue, it focuses on identifying the optimal selection within the existing OOD data. The main process is depicted in Fig.  7.
Figure 7. Illustration of outlier mining approaches. Due to the limitation of whether real OOD data exists, methods like anomaly mining allow the model to access real OOD data during training. The black dots in the figure represent real OOD data sampled from other datasets, which do not overlap with the ID categories. Utilizing such OOD data to extrapolate more representative OOD data (orange dots) is a solution to this type of problem.
POEM  (Ming et al., 2022a) raises methodology leans towards excavating anomalies that carry more representative characteristics.  Li and Vasconcelos (2020) propose a method of data resampling to acquire representative outlier data for training, assigning higher priority score reweighting to hard negative instances. Iterative optimization, leveraging adversarial principles, selects target data. POEM employs posterior sampling to unearth anomalies with elevated boundary scores from an extensive auxiliary dataset, facilitating a nuanced comprehension of the intricacies within OOD samples. DAOL  (Wang et al., 2023a) takes the disparity between authentic data and auxiliary OOD data as a starting point to formulate an OOD dataset. Utilizing the Wasserstein ball, it models the distribution of all OOD data, selecting the most challenging OOD data within the ball for training.
Beyond solely relying on raw data, another direction in addressing this issue involves synthesizing representative outlier data by utilizing authentic OOD data through information extrapolation. DivOE  (Zhu et al., 2023) introduces a novel learning objective to alleviate challenges associated with limited auxiliary OOD datasets. It achieves this by adaptively inferring and learning information from surrogate OOD data through the maximization of differences between generated OOD data and original data, given the specified anomalies. This adaptive inference extends to a broader spectrum, addressing the limitations imposed by a finite auxiliary OOD dataset. Moreover, DOE  (Wang et al., 2023b) introduces a Min-Max learning strategy to identify the most challenging OOD data for a synthetic model. Through model perturbation, the data is implicitly transformed, and the model continues learning from this perturbed data to improve its robustness.
Imbalance. Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information. The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two. To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024) incorporates a learnable tail class prototype during the training process. This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.  Choi et al. (2023) take a different approach, suggesting that factors affecting the performance of OOD detection may be related to the imbalance in the cross-class distribution of auxiliary OOD data. Consequently, it proposes an energy regularization loss, specifically regularizing auxiliary samples from the majority classes to address the issue of class imbalance in OOD data. EAT  (Wei et al., 2024) utilizes dynamically assigned virtual labels during the model training process to train OOD data, thereby expanding the classification space.
Despite the success and considerable attention received by methods like Outlier Exposure in the research community, there are voices questioning the essence of allowing access to OOD data during training. Nevertheless, concerns are raised that the superior classification performance observed in certain datasets may not necessarily translate to competitiveness in real-world deployment, challenging the original intention of OOD detection.
4. Problem: Training-agnostic OOD Detection
Training-agnostic OOD detection places a primary emphasis on adaptation strategies during test time, as opposed to the focus on classifier performance seen in training-driven OOD detection. Based on whether they rely on dependencies among test data, methods are categorized into two types: post-hoc and test-time adaptive approaches, as illustrated in Fig. 8. Post-hoc methods compute results for individual samples independently, unaffected by changes in other samples. In contrast, test-time adaptive methods exhibit dependencies among test samples.
Figure 8. Illustration of training-agnostic OOD detection approaches. Both methods require access to a pre-trained model. Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing. “A’/B’ ” means that the original features are deformed; Bank means that some samples are stored; Clustering means that clustering is performed for ID images.
4.1. Post-hoc OOD Detection Approaches
Overview. Given a well-trained model, this problem scenario involves utilizing only the intermediate results computed by the trained model during testing, without modifying any parameters of the model, to accomplish the OOD detection task. Post-hoc methods are favored for their lightweight nature, low computational costs, and the fact that they require minimal modifications to the model and objectives. Its main objective is to construct an effective scoring function that can accurately reflect the behavior of ID data. These characteristics make them highly desirable for convenient and straightforward deployment in practical scenarios.
Post-hoc approaches are categorized into five types: Output-based, Distance-based, Gradient-based, Feature-based and Density-based. Recent work on this type of problem has some recent progress. A summary of the key factors involved in such methods is given in Table  1.
Table 1. Comparison of key components in OOD detection methods.
Method Type Space
feature logit gradient probability
MSP (Hendrycks and Gimpel, 2017) Output-Based $\checkmark$
Maxlogits (Hendrycks et al., 2022) $\checkmark$
Energy (Liu et al., 2020) $\checkmark$
GEN (Liu et al., 2023b) $\checkmark$ $\checkmark$
Mahalanobis (Lee et al., 2018b) Distance-Based $\checkmark$
NNGuide (Park et al., 2023a) $\checkmark$
KNN (Sun et al., 2022) $\checkmark$
SSD (Sehwag et al., 2021) $\checkmark$
Grad (Lee and AlRegib, 2020) Gradient-Based $\checkmark$
GradNorm (Huang et al., 2021) $\checkmark$ $\checkmark$ $\checkmark$
GradOrth (Behpour et al., 2023) $\checkmark$ $\checkmark$
GAIA (Chen et al., 2023) $\checkmark$ $\checkmark$
OPNP (Chen et al., 2024a) $\checkmark$ $\checkmark$
ODIN (Liang et al., 2018a) Feature-Based $\checkmark$ $\checkmark$ $\checkmark$
ReAct (Sun et al., 2021) $\checkmark$
VRA (Xu et al., 2023) $\checkmark$
SHE (Zhang et al., 2022) $\checkmark$
Vim (Wang et al., 2022c) $\checkmark$ $\checkmark$
Neco (Ammar et al., 2024) $\checkmark$ $\checkmark$
ASH (Djurisic et al., 2023) $\checkmark$
NAC (Liu et al., 2024a) $\checkmark$
Optimal-FS (Zhao et al., 2024) $\checkmark$ $\checkmark$
BLOOD (Jelenić et al., 2024) $\checkmark$
SCALE (Xu et al., 2024) $\checkmark$
ConjNorm (Peng et al., 2024) Density-based $\checkmark$ $\checkmark$
GEM (Morteza and Li, 2022) $\checkmark$ $\checkmark$ $\checkmark$
Output-based. Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others. MSP  (Hendrycks and Gimpel, 2017) was the first to employ the maximum softmax value to validate OOD detection effectiveness. For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category. In contrast to the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by comparing the maximum logit value in the logits vector output by the neural network. Logits, representing the model’s confidence in each category, are the neural network’s output before the softmax layer, without undergoing softmax transformation. Meanwhile, Energy  (Liu et al., 2020) introduces the Helmholtz Free Energy, which theoretically aligns with the input probability density and is less susceptible to overconfidence issues. GEN  (Liu et al., 2023b) introduces the concept of generalized entropy and directly utilizes Bregman divergence to compute the statistical distance between the model’s probability output and uniform distribution, aiming to identify OOD data. Additionally, leveraging sufficient prior knowledge might be a viable solution. ZODE  (Xue et al., 2024) perform predictions on samples across multiple pre-trained models simultaneously to determine whether multiple models can identify OOD samples, using this as a basis to distinguish between data. LogicOOD  (Kirchheim et al., 2024) present a novel approach that uses first-order logic for knowledge representation to perform OOD detection. This reasoning system uses prior knowledge to infer whether an input is consistent with the prior knowledge about the training distribution. It is particularly user-friendly in terms of interpretability, as it allows for comparing the output of samples against a knowledge base to determine whether they belong to the ID data.
Distance-based. Another approach in OOD detection research focuses on measuring statistical distance metrics. Mahalanobis  (Lee et al., 2018b) is typically computed by calculating the distance between the feature vector and its mean. Specifically, for each class, we compute the mean and covariance matrix of its feature vectors. During testing, it calculates the Mahalanobis distance between the feature vector and the mean of each class. SSD  (Sehwag et al., 2021) essentially utilizes the Mahalanobis distance. After being trained on unlabeled ID data via self-supervised representation learning, it employs the Mahalanobis distance as a statistical measure for classification using the pre-trained model. In comparison, while the Mahalanobis distance makes strong distributional assumptions about the data, KNN  (Sun et al., 2022) explores the effectiveness of non-parametric nearest-neighbor distance for OOD detection. By measuring the k-nearest neighbor distance between input embeddings and training set embeddings, a threshold is designed to determine whether the data belongs to the ID data. NNGuide  (Park et al., 2023a) takes a step further in the direction of granularity by combining the idea of KNN. It assigns weights before the traditional OOD Score, depending on the nearest neighbor distance between the sample and the embeddings in the training set.
Gradient-based. Grad  (Lee and AlRegib, 2020) research indicates that gradient-based methods also contribute to OOD detection by quantifying model uncertainty through the gradients propagated in reverse during backpropagation. If the input sample is an ID sample, the gradients of the model with respect to these samples tend to be relatively small and stable. Conversely, the gradients are typically larger or more irregular for OOD samples. GradNorm  (Huang et al., 2021) posits that the gradient magnitude for ID data surpasses that of OOD data. Leveraging this observation, it employs gradient vector norms, computing KL divergence through softmax output and uniform distribution backpropagation to detect OOD data. In contrast, GradOrth  (Behpour et al., 2023) adopts an alternative perspective, recognizing that crucial features of OOD data reside in a low-rank subspace. Consequently, it shifts focus to calculating gradient projection norms in this subspace to identify OOD data. GAIA  (Chen et al., 2023) employs a combination of gradient anomaly checks and aggregation. It enables the model to interpret indeterminacy attributively, introducing channel-wise average abnormality and zero-deflation abnormality without prior knowledge to gauge the extent of data distribution variations. The OPNP  (Chen et al., 2024a) method discovers that the OOD detection capability of the model is highly sensitive to parameters and activated neurons that are close to zero. Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability.
Feature-based. Another avenue of research into OOD detection involves the impact of neural network intermediate variables relative to the final prediction. Inspired by adversarial examples, ODIN  (Liang et al., 2018a) experimentally discovered that adding small perturbations to the input image features can more effectively detect OOD data. The perturbation formula is as follows:
(4) $\tilde{x}=x+\varepsilon sign(\bigtriangledown_{x}logmax_{c}p_{c}(x)),$
where the parameter $\varepsilon$ is the perturbation magnitude. For a given input $x$, compute its logit output $p_{c}(x)$. Therefore, ReAct  (Sun et al., 2021) focuses on the high activation values in the intermediate results of the model. These activation values do not affect the model’s classification, but truncating high activation values can significantly improve OOD detection performance. VRA  (Xu et al., 2023), an extended iteration of ReAct  (Sun et al., 2021), builds upon the premise of ReAct which truncates only high activation values. However, VRA posits that this might not be the optimal solution and, therefore, employs a variational approach to seek an optimal solution. It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data. SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison. It employs a storage mechanism to calculate the average logits for each class, preparing this simple storage pattern for subsequent OOD detection tasks. Drawing inspiration from experience, DDCS  (Yuan et al., 2024) adapts to select suitable channels for data classification after correcting each channel in the neural network. These channels are evaluated based on inter-class similarity and variance to measure their discriminative power for ID data. LINe  (Yong Hyun Ahn, 2023) similarly emphasizes neuron outputs at the feature level. It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs.
Feature shaping is emerging as a prominent area of research focus. This methodology entails the refinement of intermediate representations—specifically, the intermediate features—within the forward propagation phase of the model. The salient advantage of this approach lies in its non-disruptive nature concerning the original classification results, coupled with its simplicity and efficacy. Building upon this foundation, ViM  (Wang et al., 2022c) simultaneously considers the roles of features, logits, and probabilities, constructing virtual logits to aggregate information for decision-making. By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection. Its computation formula can be expressed in the following form:
(5) $-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
where $\alpha$ is a scaling constant, computed by the model. Here $z=z^{P}+z^{P^{\perp}}$ and $z^{P^{\perp}}$ is the projection of $z$ to $P^{\perp}$. And it have $Wz^{P^{\perp}}=0$. Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020), and $f(z)$ represents the logit output of the model. The first term here represents virtual logits, while the second term represents the score of the energy function. Neco  (Ammar et al., 2024) subsequently reveals the prevalent phenomenon of neural collapse in contemporary neural networks, impacting OOD detection performance. The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data. ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted. NAC  (Liu et al., 2024a) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data. By quantifying this statistical property, NAC aims to distinguish between ID and OOD data. Based on the existing literature on feature shaping,  Zhao et al. (2024) utilize a piecewise constant shaping function to partition the feature domain into disjoint intervals, estimating a scalar within each interval as an approximation. As the interval width approaches zero, an approximation of the maximum logits can be obtained. The goal of BLOOD  (Jelenić et al., 2024) is to smooth the representations of intermediate layers in neural networks for predicting OOD data. This study found that, compared to OOD data, ID data exhibits smoother variations in the intermediate layer representations of neural networks. Leveraging this characteristic, new statistical measures can be designed to discriminate against anomalous data. In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH. SCALE  (Xu et al., 2024) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data. Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.
Density-based. Recent advancements in density-based OOD detection models have demonstrated substantial performance gains, predicated on the models’ ability to accurately capture and understand the intrinsic characteristics of the true data distribution. For example, GEM  (Morteza and Li, 2022) models the feature space of ID data as a class-conditional multivariate Gaussian distribution. Under this assumption, it designs new statistical metrics to validate the model’s performance. Using a modeling approach based on Gaussian mixture models, GEM Score is aligned with the true log-likelihood for capturing OOD uncertainty. However, while GEM strictly relies on Gaussian assumptions, recent works ConjNorm  (Peng et al., 2024) introduce a novel framework based on Bregman divergence, extending the consideration of data distributions to encompass an exponential family of distributions. By redefining density functions for data, this model has a broader range of applications.
4.2. Test-Time Adaptive OOD Detection Approaches
Overview. Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize test data, either the complete test set or a series of unlabeled mini-batches, to enhance OOD detection performance through model adaptation. Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge is impossible. These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free. Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
Model-optimization-based. Model optimization-based methods enhance the trained model by leveraging unlabeled data during the post-training phase.
A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection. The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them. Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024) explains how they help OOD detection from the lens of separability and learnability. Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model. In addition, the extra training requirement in WOODS, thought necessary, incurs significant costs and is generally undesirable.
Another line of model-optimization-based approaches draws inspiration from the semi-supervised-learning (SSL) techniques  (Zhu, 2005), aiming for a more efficient, lightweight training process during the post-training phase. Pseudo-labeling  (Arazo et al., 2020) is a simple but effective way to label the test data, which enhances learning from test unlabeled data. The method proposed by  Yang et al. (2023a), termed AUTO, employs only pseudo-OOD data to refine the model. The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification. In contrast, ATTA  (Gao et al., 2023b) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model. SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
Model-optimization-free. Modifying the original trained model is infeasible in certain security-sensitive scenarios. Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” (MOF) techniques, are increasingly garnering interest. These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself. ETLT observes a linear correlation between the feature representation and the OOD score of a test input. In other words, for a given image, the pair consisting of its feature and OOD score (feature, OOD score) exhibits a linear correlation. Furthermore, the aforementioned pairs of ID and OOD data is linearly separable. Based on the observation, ETLT proposes to learn a linear regression module trained from the (feature, OOD score) pair. The availability of complete test datasets may not always be feasible, therefore,  Fan et al. (2022) also provide an online variant to ensure safer deployment. Similarly, GOODAT develops an add-on named graph masker, designed specifically for graph data. It integrates GIB-boosted losses and employs it as the metric for OOD scoring. In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach. The core principle behind AdaOOD is the maintenance of a memory bank, similar to the approach taken by AUTO.
Online v.s. offline. Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment. In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023b) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
More challenging scenario. In the context of test-time OOD detection scenarios, some scholars have proposed more challenging configurations that demand a higher level of capability from the models. MOL (Wu et al., 2023) introduces a more realistic problem scenario, namely Continuous Adaptive Out-of-Distribution (CAOOD) detection, aimed at addressing the challenge of constantly changing ID and OOD distributions in the real world. The meta-learning approach is employed to swiftly adapt models in response to the complexities encountered in various scenarios in MOL.
Figure 9. Overview and trends in LPM-based OOD Detection Methods.
5. Problem: Large Pre-trained Model-based OOD Detection
Large pre-trained models have showcased remarkable performance in numerous downstream ID classification tasks, but their potential in OOD detection tasks remains a less-explored area. Recent research  (Vaze et al., 2022) highlights a correlation between higher ID classification accuracy and better OOD detection performance. Consequently, large pre-trained model-based OOD detection problem comes naturally. In recent years, large pre-trained models of various types, including single-modal (ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), Diffusion (Ho et al., 2020)), visual language models (VLMs) (CLIP (Radford et al., 2021), multi-modal Diffusion (Saharia et al., 2022) ALIGN (Jia et al., 2021), ), and large language models (LLMs) (GPT3 (Brown et al., 2020)), have been increasingly utilized for OOD detection tasks, as shown in Fig. 9. Leveraging the powerful representational capabilities of large pre-trained models has further relaxed the constraints of OOD detection tasks, leading to a focus on more challenging and realistic scenarios, which has emerged as a new hotspot. Given the number of ID shots exposed to the large pre-trained model, large pre-trained model-based OOD detection can be classified into Zero-shot, Few-shot, and Full-shot OOD detection, as shown in Fig. 10. The performance evaluations of several relevant competitive methods are summarized in Table  2 to provide an understanding of the performance level of OOD detection in this area.
Figure 10. Illustration of large pre-trained model-based OOD detection approaches. In the training phase, zero-shot approaches require only category labels of ID classes. Few-shot approaches need a subset of images of each ID class along with the category labels (indicated by dashed lines). Full-shot approaches utilize both the category labels and all images of each ID class. None of these approaches use labels or images of OOD categories.
5.1. Zero-shot OOD Detection Approaches
Overview. Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data. Note that we do not need to have access to ID images and rely solely on textual category knowledge. It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training.
VLMs-based. Existing research on zero-shot OOD detection using VLMs can be categorized into two main approaches based on which VLMs are utilized as a basic backbone. Diffusion-based. Traditional generative approaches, such as diffusion models  (Ho et al., 2020), capture the ID distribution effectively. Subsequently, these methods are used to identify OOD samples by assessing the likelihood that a given test input is derived from an OOD source. In a recent study in RONIN  (Nguyen et al., 2024), the diffusion model has been applied to achieve OOD object detection, which utilizes the diffusion model to generate ID inpainting, while the CLIP model is employed to calculate the similarity score. CLIP-based. Given the remarkable proficiency of CLIP in correlating texts and images, a substantial number of researchers have ventured into zero-shot OOD detection leveraging CLIP. Typically, in the zero-shot OOD detection setup, add-ons are added to the pre-trained model to better adapt it to the OOD detection task. ZOC  (Esmaeilpour et al., 2022) achieves zeto-shot OOD detection by training an image description generator on a large image captioning dataset  (Lin et al., 2014), enabling the model to generate candidate unseen labels. It’s essential to note that ZOC treats CLIP merely as a feature extractor and doesn’t impart OOD detection capabilities to CLIP itself. The subsequent work CLIPN  (Wang et al., 2023c) empowers CLIP with the ability to say “no” by a “no-prompt” encoder. Other approaches alleviate the need for additional training and focus on enhancing the performance of post-hoc zero-shot OOD detection. A simple baseline of clip-based zero-shot OOD detection is to use the normalized text-image similarity as an OOD score.  Ming et al. (2022b) further substitute the score with the maximum concept matching (MCM) score and offer a comprehensive theoretical explanation of this modification. Moreover, NegLabel  (Anonymous, 2024), LAPT (Zhang et al., 2024) and CLIPScope (Fu et al., 2024) enriches the text information through a predetermined corpus. Concretely, NegLabel selects negative words (labels) from the corpus, which are dissimilar to ID labels, to enhance OOD detection performance. Futhermore, CLIPScope uses those labels to revise the original score by applying Bayesian rules. In contrast, LAPT leverages text-to-image generation models or image retrieval models to obtain images corresponding to the given ID and Negative ID labels, followed by prompt tuning. It is noteworthy that ZOC, CLIPN, NegLabel and CLIPScope each devise unique OOD scores for integrating into their respective models. Conversely, the MCM score stands out for its generalizability across different models. The calculation of MCM is as follows:
(6) $S_{MCM}(x^{\prime};\mathcal{Y}_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}},$
where $x^{\prime}$ is the test input image; $\mathcal{Y}_{in}=\{y_{1},y_{2},\ldots,y_{K}\}$ represents the label set; $y_{i}$ denotes the $i$-th label in the label set; $T(t_{i})$ corresponds to the concept vector for the text prompt $t_{i}$; $s_{i}(x^{\prime})$ is the cosine similarity score between the image feature $I(x^{\prime})$ and the concept vector $T(t_{i})$, computed as $s_{i}(x^{\prime})=\frac{I(x^{\prime})\cdot T(t_{i})}{\|I(x^{\prime})\|\|T(t_{i})\|}$; $e^{s_{i}(x^{\prime})/\tau}$ is the exponentiated similarity score normalized by the temperature parameter $\tau$; $K$ is the number of labels in the label set; $\tau$ is the temperature parameter of the softmax function; and $S_{MCM}$ is the maximum concept matching score. Some works have explored variations of zero-shot OOD detection or more challenging scenarios, such as detecting ID objects (Miyai et al., 2023c) or handling labels with noise (Ding and Pang, 2024), which will be discussed in the Section.7.
LLM-based. With the flourishing development of LLMs, new opportunities have emerged in the field of OOD detection with the application of LLMs. The strength of LLMs resides in their extensive world knowledge, enabling them to furnish comprehensive details about ID labels. Nevertheless, the issue of generating false or misleading information, known as hallucinations, poses a significant challenge when employing LLMs for OOD detection tasks. Existing methods (Huang et al., 2024; Dai et al., 2023) are based on an observation: two categories of images may have very similar visual features but different properties in the semantic space.  Dai et al. (2023) introduce a consistency-based calibration method with an object detector to mitigate hallucinations and utilize LLMs’ world knowledge for describing ID classes. However,  Huang et al. (2024) propose ODPC to generate peer-class based on ID labels and OOD samples within text and image modality. It should be noted that ODPC requires ID images for training MLP, indicating that it is not a zero-shot method. Regarding the OOD score,  Dai et al. (2023) adopt MSP, however, ODPC uses a KNN-based score. Later, a study  (Liu et al., 2024b) on LLMs was proposed to study three problems with LLMs doing OOD detection: the propensity of LLMs for near-OOD and far-OOD, the effect of different fine-tuning methods, and the choice of OOD score.  Liu et al. (2024b) find that LLMs are natural far-OOD detectors, generative fine-tuning is better than discriminative fine-tuning, and cosine distance is sufficient as OOD score because the embedding space of LLMs exhibits anisotropy.
Remark. As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels. However,  Fort et al. (2021) argue that in deployment scenarios, labels for OOD are either readily available. They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance. However, access to OOD labels is rare in real-life scenarios and seems less reasonable.
Table 2. Performance evaluation of some competitive methods using the ImageNet-1K dataset as the ID dataset and iNaturalist, SUN, Places, and Textures as OOD datasets. The best result is emphasized in bold.
Scenario Method iNaturalist SUN Places Textures Average
AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ PFR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$
Zero-shot ZOC  (Esmaeilpour et al., 2022) 86.09 87.30 81.20 81.51 83.39 73.06 76.46 98.90 81.79 85.19
MCM  (Ming et al., 2022b) 94.59 32.20 92.25 38.80 90.31 46.20 86.12 58.50 90.82 43.93
CLIPN  (Wang et al., 2023c) 95.27 23.94 93.93 26.17 92.28 33.45 90.93 40.83 93.10 31.10
NegLabel  (Anonymous, 2024) 99.49 1.91 95.49 20.53 91.64 35.59 90.22 43.56 94.21 25.40
Few-shot CoOp  (Zhou et al., 2022a) 93.77 29.81 93.29 40.83 90.58 40.11 89.47 45.00 91.78 51.68
LoCoOp  (Miyai et al., 2023a) 96.86 16.05 95.07 23.44 91.98 32.87 90.19 42.28 93.52 28.66
NegPrompt  (Li et al., 2024b) 98.73 6.32 95.55 22.89 93.34 27.60 91.60 35.21 94.81 23.01
5.2. Few-shot OOD Detection Approaches
Overview. Given the large pre-trained model and a few ID data, we can adapt the model using the ID data and subsequently detect OOD test data. Zero-shot OOD detection does not necessitate any training images, making it suitable for scenarios with high-security requirements. However, it may face challenges related to domain gaps with ID downstream data, which can limit the performance of zero-shot methods. Therefore, there are many few-shot methods employed in OOD detection, and their effectiveness is often superior to that of zero-shot OOD detection.
Studies. The direct way of adapting a large pre-trained model with several ID samples is fine-tuning.  Ming and Li (2024) and  Dong et al. (2023) conduct studies on the impact of fine-tuning on OOD detection within the context of VLMs, with a special focus on CLIP.  Ming and Li (2024) pay more attention to the role of parameter-efficient fine-tuning (PEFT) methods and the OOD score. The MCM score, introduced by  Ming et al. (2022b) as an innovative measure, alongside methods based on prompt learning for detecting OOD instances, is recognized for its effectiveness. Similarly,  Dong et al. (2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  ming2023does. Following this,  Kim et al. (2024) believe that Finetune Like You Pretrain (FLYP), a fine-tuning method, deserves attention due to its good performance on classification tasks. Specifically, FLYP is to mimic contrastive language-image pre-training as CLIP. After comparing the performance of FLYP and PEFT methods on zero-shot OOD detection,  Kim et al. (2024) find FLYP yields better OOD detection performance than PEFT. It should be noted that while  Fort et al. (2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
Fine-tuning based. A prevalent method for fine-tuning CLIP using limited ID data involves prompt learning (Zhou et al., 2022b, 2023), where the fine-tuning focuses on the prompt’s contextual words, maintaining the pre-trained parameters unchanged. To distinguish between ID and OOD data, both LoCoOp  (Miyai et al., 2023a) and ID-like Prompting Learning  (Bai et al., 2024) are designed to enhance the learning of context vectors from the near-ID feature perspective. LoCoOp employs entropy maximization to distance ID-irrelevant local features (such as the backgrounds in ID images) from the textual embeddings of ID classes. Similarly, ID-like Prompting Learning generates ID-like data—outliers within the vicinity of ID samples—to refine the context vector. Additionally, it incorporates a diversity loss to enhance the variance among the sampled OOD data. However, recent work EOK-Prompt (Zeng et al., 2024) and GalLop (Lafon et al., 2024) argues that optimizing only a prompt using numerous features from an image wastes valuable information. Instead, EOK proposes an orthogonal method to optimize the novel local prompt, making better use of the limited image data, while GalLop optimizes global prompts and local prompts by leveraging global and local features, respectively. Furthermore, based on the mentioned studies, DSFG  (Dong et al., 2023) posits that traditional fine-tuning approaches applied to CLIP may inadvertently lead to a loss of critical OOD knowledge. To address this issue, DSFG adopts a strategy to retain broad OOD knowledge by merging the original features with those modified through fine-tuning, followed by the training of a classifier. DSFG’s plug-and-play capability makes it seamlessly compatible with all fine-tuning-based approaches, enhancing its practicality and value. The aforementioned few-shot tuning methods inevitably suffer from overfitting on limited-shot instances. Chen et al. (2024b) proposed a training-free few-shot OOD detection method, Dual-Adapter(Chen et al., 2024b), which constructs two types of adapters by extracting positive and negative features to aid in OOD detection. The currently mentioned few-shot methods lack OOD supervision, so CLIP-OS (Sun et al., 2024) tries to find such supervision and also achieves stunning results
Meta-learning based. Meta-learning aims to devise a learning approach that enables rapid adaptation to new challenges  (Vilalta and Drissi, 2002; Hospedales et al., 2021). OOD-MAML  (Jeong and Kim, 2020) adapt model-agnostic meta-learning (MAML) for few-shot OOD detection. It generates OOD samples and incorporates them along with ID data for the adapted N-way K-shot task, which is divided into N sub-tasks, each focusing on K-shot OOD detection. The decision on whether test data is OOD is based on the outcomes of these fast and simple N sub-tasks. In contrast, HyperMix  (Mehta et al., 2024) advocates for employing a hypernetwork-based method to enhance sample augmentation without the necessity for extra outliers. This is because classes not included in a specific meta-training task can act as OOD samples.
5.3. Full-shot OOD Detection Approaches
Overview. This setup is generally less realistic than the first two (zero-shot and few-shot). However, we list them separately to ensure a comprehensive review of existing methods across the spectrum. Given the full set of ID data and corresponding labels, VLMs can enhance OOD detection significantly by fine-tuning. Moreover, a novel task called “PT-OOD” detection is introduced.
Fine-tuning based. With access to the complete dataset, then more data can be used to fine-tune the large pre-trained model or the data can be used to better simulate the ID distribution, facilitating the differentiation of OOD data. NPOS  (Tao et al., 2023) proposes a non-parametric outlier synthesis technique to distinguish ID and OOD data by fine-tuning CLIP with complete ID data. In contrast, TOE  (Park et al., 2023b), while also using CE loss to constrain the model during fine-tuning, builds on the ideas of OE by focusing on textual outliers within the CLIP framework to control the model’s recognition capabilities, which differs significantly from directly using OOD images.
PT-OOD Detection. “PT-OO” samples are OOD samples with overlap in pretraining data. After investigating and elucidating the effects of various pre-training methodologies (supervised, self-supervised) on PT-OOD detection,  Miyai et al. (2023b) observe the low linear separability in feature space significantly degrades the PT-OOD detection performance. They suggest using distinctive features for each instance to distinguish between ID and OOD samples.
6. Evaluation and Application
6.1. Evaluation metrics
In the vast majority of OOD detection tasks in the visual domain, the following evaluation metrics are commonly used:
AUROC (Area Under the Receiver Operating Characteristic curve). This metric quantifies the likelihood that a classifier will assign higher scores to ID samples compared to OOD samples. An elevated AUROC value is indicative of superior model performance, signifying an enhanced ability to distinguish between ID and OOD instances. Consequently, a higher value is desirable.
AUPR (Area under the Precision-Recall curve). This metric is pertinent when the ID class is considered the positive class and is particularly valuable in the context of imbalanced class distributions. It assesses the balance between precision and recall, with a higher AUPR value indicating superior model performance. Therefore, a greater AUPR value is desirable.
FPR@95 (False Positive Rate at 95% True Positive Rate). This metric delineates the false positive rate (FPR) at the juncture where the true positive rate (TPR) reaches 95%. It essentially gauges the proportion of OOD samples erroneously identified as ID, thus providing insight into the model’s propensity for false alarms at a high sensitivity threshold. A reduced FPR@95% TPR is indicative of a model’s enhanced specificity in correctly flagging OOD samples while maintaining high sensitivity towards ID samples. Therefore, a lower value is desirable.
These metrics provide comprehensive insights into the OOD detection performance, considering different aspects such as the ability to distinguish between ID and OOD samples, handling imbalanced class sensitivity towards ID samples, and controlling the false positive rate at a specific true positive rate threshold.
6.2. Experimental Protocols
In the traditional experimental protocol for OOD detection, test data is exclusively classified as either ID or OOD. However, as the field has advanced, there is now a more nuanced distinction between OOD and ID data, which has led to variations in the evaluation process.
Subsequently, OOD data is categorized into near-OOD and far-OOD based on the degree of covariate shift from ID data. This categorization corresponds to dividing OOD detection tasks into near-OOD and far-OOD detection. It is evident the near-OOD detection task is more challenging, however, numerous methods  (Fort et al., 2021; Ming et al., 2022b; Bai et al., 2024) have demonstrated excellent performance in this area.
Recently,  Yang et al. (2023b); Bai et al. (2023) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously. This is crucial to prevent the loss of model generalization. The samples mentioned earlier are termed as cs-ID data, an abbreviation for “covariate shift ID” data. Consequently, a new experimental protocol has been explored, called full-spectrum OOD detection. During the testing phase, the model is expected to identify near-OOD and far-OOD instances. Additionally, it should refuse to provide predictions for the OOD data and accurately predict ID and cs-ID data.
6.3. Application
Table 3. Summary of Datasets. The CARLA System is a simulation platform designed for evaluating OOD Detection in the field of autonomous driving, hence its entire row is filled with dashes. Other “-” symbols represent numbers that vary depending on usage. More detailed descriptions can be found in the code repo.
TASK Dataset Name Data Type # Classes # Samples Papers
Image Classification CIFAR-10 Images 10 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
CIFAR-100 Images 100 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
MNIST Images 10 70,000 (Hendrycks and Gimpel, 2017; Zhao et al., 2024)
ImageNet-1K Images 1,000 1,431,167 (Ming et al., 2022b; Wang et al., 2023c)
iNaturalist Images 5,089 675,170 (Ming et al., 2022b; Wang et al., 2023c)
SUN Images 397 108,754 (Ming et al., 2022b; Wang et al., 2023c)
Places Images >205 >2,500,000 (Ming et al., 2022b; Wang et al., 2023c)
Textures Images 47 5,640 (Ming et al., 2022b; Wang et al., 2023c)
Semantic Segmentation Cityscapes Images - 25,000 (Gao et al., 2023b)
Road Anomaly Dataset Images - 100 (Gao et al., 2023b)
Object Detection PASCAL VOC Images 20 2,913 (Nguyen et al., 2024; Du et al., 2022b)
BBD100K Video->Image Variable 100,000 (Nguyen et al., 2024; Du et al., 2022b)
Autonomous Driving CARLA System - - - (Mao et al., 2024)
Medical Image Analysis Kvasir-Capsul Images 4,741,621 (Tan et al., 2024)
Text Category News Category Text - 210,000 (Arora et al., 2021)
SST-2 Text - 215,154 (Arora et al., 2021)
Intent Detection CLINC150 Text 150 22,500 (Zhan et al., 2022)
Banking Text 77 13,083 (Zhan et al., 2022)
StackOverflow Text 20 20,000 (Zhan et al., 2022)
Audio MSCW Audio - >23400000 (Bukhsh and Saeed, 2023)
Vocalsound Audio - 21,024 (Bukhsh and Saeed, 2023)
Graph data TU Graph data - Variable (Wang et al., 2024)
OGB Graph data - Variable (Wang et al., 2024)
6.3.1. Computer Vision
Most of the efforts in OOD detection have been devoted to the field of computer vision, we list extensive vision-related tasks as follows:
  • Image Classification. The majority of tasks discussed in this paper focus on OOD detection within the realm of image classification. In such task scenarios, commonly utilized ID datasets include MNIST  (Xiao et al., 2017), CIFAR-10  (Krizhevsky et al., 2009), CIFAR-100  (Krizhevsky et al., 2009), and ImageNet-1K  (Deng et al., 2009). Various OOD datasets are constructed accordingly for evaluating different methods  (Yang et al., 2022b), with the most frequently employed datasets being iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017) and Textures (Cimpoi et al., 2014). Additionally, full-spectrum OOD is typically evaluated on three benchmarks: DIGITS, OBJECTS, and COVID, as proposed by  Yang et al. (2023b).
  • Semantic Segmentation. Recent works  (Gao et al., 2023b) have started delving into the dense OOD detection task, also known as anomaly segmentation. The datasets used for evaluation include the Cityscapes dataset  (Cordts et al., 2016), the Road Anomaly dataset  (Lis et al., 2019), and the recently developed SOOD-ImageNet  (Bacchin et al., 2024).
  • Object Detection. The application of methods related to OOD detection in the field of object detection is relatively nascent, with only a few studies exploring this area  (Du et al., 2022b). Evaluations are commonly performed using datasets like PASCAL-VOC  (Everingham et al., 2015) and Berkeley DeepDrive-100K  (Yu et al., 2018).
  • Autonomous Driving. Autonomous driving has long been a crucial practical application of OOD Detection. Recently, Mao et al. (2024) utilized the CARLA (Dosovitskiy et al., 2017) system to simulate and evaluate the performance of OOD Detection in autonomous driving scenarios.
  • Medical Image Analysis. In the field of medical image analysis, OOD detection is crucial. Depending on the specific category of medical image, OOD detection employs various datasets including CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021).
OOD detection has significant applications across various fields such as human action recognition (Sim et al., 2023) and solar image analysis (Giger and Csillaghy, 2024; Das et al., 2024). For further details, please refer to the accompanying code repository due to space limitations. The same applies to the subsequent paragraphs.
6.3.2. Natural Language Processing
OOD detection is also explored in various tasks across numerous Natural Language Processing applications. The two most common applications are as follows:
  • Intent Detection. Intent Detection is a significant application of OOD detection in NLP. The datasets used for evaluation include CLINC150  (Larson et al., 2019), Banking  (Casanueva et al., 2020), StackOverflow  (Xu et al., 2015), among others.
  • Text Catagory. In text category OOD detection applications, datasets like News Category  (Misra, 2022) and SST-2 (Socher et al., 2013) are commonly utilized to form ID/OOD pairs and assess the models’ detection capabilities.
6.3.3. Beyond Computer Vision and Natural Language Processing
In addition to the two data modalities mentioned above, OOD detection still has many important applications across various types of data.
  • Audio data. In audio OOD detection, MSCW (Micro-EN) (Mazumder et al., 2021) and Vocalsound  (Gong et al., 2022) are usually used as ID datasets, and they also act as OOD for each other.
  • Graph data. Recent studies have proposed various approaches for OOD detection in graph data. The existing graph-level OOD detection benchmark comprises a total of 10 pairs of datasets from TU  (Morris et al., 2020) and OGB datasets (Hu et al., 2020), which has been widely utilized.
  • Reinforcement learning. Currently, there is a rising trend of integrating OOD detection with reinforcement learning  (Leonardos et al., 2023) to bolster model robustness.  Mohammed and Valdenegro-Toro (2021) provide a benchmark and explore methods for crafting tailored reinforcement learning environments that can generate OOD data.
7. Emerging Trends And Open Challenges
Despite rapid advancements in OOD detection, numerous emerging trends and less-explored challenges remain. In this section, we explore emerging trends and open challenges from three distinct perspectives: methodologies, scenarios, and applications.
7.1. Better methodologies of OOD detection
Meta-Learning Adaptation. Faced with the quick adaptation challenge in test-time OOD detection, mete-learning algorithms, which provide a learning-to-learn paradigm to efficiently adapt the model to new test data, may be the solution. Additionally, in addressing the dauntingly vast sample space of potential outliers inherent in training-driven OOD detection, improved sampling methods could offer a pathway to the efficient utilization of outliers  (Ming et al., 2022a).
Theoretically-Driven Score Designing. In the traditional single-modal regime, numerous post-hoc scores have been carefully designed for OOD detection, effectively reducing training costs. However, as the field has progressed into the multi-modal domain, there is an increasing demand for theoretically-driven score designs. MCM  (Ming et al., 2022b) is a notable example, but it is merely an extension of Softmax and does not deeply explore the relationships between text and image. Therefore, more advanced scores are needed.
7.2. More practical scenarios of OOD detection
Under the current trend, there is a growing need for the emergence of more practical scenarios, driven by the limitations of existing impractical restrictions.
Quick Test-Time Adaptation. The advent of test-time scenarios like CAOOD  (Wu et al., 2023) significantly advances the application of OOD detection in real-world contexts, promising improved reliability and adaptability.
Multi-Modal Detection. Exploring multi-modal OOD detection enhances our grasp of data dynamics and boosts model efficacy across varied sensory inputs. Additionally, the forthcoming integration of LLMs into multi-modal OOD detection stands to transform the field, marrying advanced linguistic analysis with other data types to forge more effective and versatile detection systems for complex applications.
Open-Vocabulary Scenario. Existing LPM-based methods assume that ID categories are accessible. However, this assumption does not hold in open-vocabulary scenarios.  Li et al. (2024b) addresses this by learning migratable negative prompts. It uses a subset of ID labels for prompt learning, and these learned prompts can be applied to other unseen labels. OOD detection in open-vocabulary settings further relaxes the requirement for prior access to ID knowledge, paving the way for the development of more generalized models.
Noise Settings. While previous research has largely focused on standard clean label settings, few studies have explored the other side of the coin: noisy label scenarios.  Humblot-Renaux et al. (2024) investigate practical scenarios and examine the impact of varying levels of noise on OOD detection. Then they provide several key takeaways for improving OOD detection in noisy label environments.
7.3. New applications of OOD detection
Additional Modalities. While OOD detection has made strides across diverse sectors, its potential in speech and physiological signal analysis remains largely untapped. In particular, OOD issues are prevalent in emotion-related physiological signals due to the variability in subjects’ emotional states. Therefore, this represents a promising area for further exploration. Some of the current methods  (Dong et al., 2024) also provide benchmarks for reference.
Human-in-the-loop Application. A pivotal future direction, proposed by  Vishwakarma et al. (2024), is the integration of human insight into the detection process. This human-in-the-loop approach, particularly in high-stakes decision-making, will be vital for enhancing the accuracy and responsiveness of OOD detection systems, merging human intuition with algorithmic precision.
Web Image Scraping. To automate the process of scraping images from the Internet,  Miyai et al. (2023c) propose a zero-shot ID detection task, a concept derived from zero-shot OOD detection. An image will be classified as an ID image if it contains ID objects; it will only be considered an OOD image if it lacks any ID objects. It presents a novel perspective on the application of OOD detection and merits further exploration
8. Conclusion
OOD detection is critical for trustworthy machine learning. In this paper, we provide a comprehensive review of recent advances in OOD detection, focusing for the first time on the problem scenario perspective: training-driven, training-agnostic, and large pre-trained model-based OOD detection. We also summarize extensively used evaluation metrics, experimental protocols, and diverse applications. We believe that our novel taxonomy of existing papers and extensive discussion of emerging trends will contribute to a better understanding of the current state of research, assist practitioners in selecting suitable approaches, and inspire new research hotspots.

### v3

@@
-Recent Advances in OOD Detection: Problems and Approaches
+Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances
 Abstract.
@@
 However, many recent works concentrate on non-traditional OOD detection scenarios, such as test-time adaptation, multi-modal data sources and other novel contexts.
-In this survey, we uniquely review recent advances in OOD detection from the problem scenario perspective for the first time.
-According to whether the training process is completely controlled, we divide OOD detection methods into training-driven and training-agnostic.
+In this survey, we uniquely review recent advances in OOD detection from the task-oriented perspective for the first time.
+According to the user’s access to the model, that is, whether the OOD detection method is allowed to modify or retrain the model, we classify the methods as training-driven or training-agnostic.
 Besides, considering the rapid development of pre-trained models, large pre-trained model-based OOD detection is also regarded as an important category and discussed separately.
@@
 1. Introduction
-Machine learning methods  (LeCun et al., 2015) have made significant progress under the closed-world assumption, where test data is drawn from the same distribution as the training set, known as in-distribution (ID).
+Machine learning methods have made significant progress under the closed-world assumption, where test data is drawn from the same category as the training set, known as in-distribution (ID).
 However, in the real world, models inevitably encounter test samples that do not belong to any training set category, commonly referred to as out-of-distribution (OOD) data.
@@
 Similarly, autonomous driving algorithms  (Geiger et al., 2012) should detect unknown scenarios and resort to human control to avoid accidents caused by arbitrary judgment.
+Table 1.
+Comparison of recent survey papers on OOD detection.
+“Type” indicates whether the survey is organized from a methodological perspective (“Method”) or from a broader task-oriented perspective (“Task”).
+“OOD-Only” indicates an exclusive focus on OOD detection; “LPMs” refers to coverage of methods based on large pretrained models; “Application” refers to in-depth analysis of real-world use cases; “Multi-Modal” indicates consideration of multiple data modalities; “Test-Time” indicates surveys that explicitly distinguish test-time adaptive methods from post-hoc approaches and provide dedicated discussion on test-time adaptation.
+Study Year Type OOD-Only LPMs Application Multi-Modal Test-Time
+Salehi et al. (Salehi et al., 2022) 2021 Method
+Yang et al. (Yang et al., 2021) 2021 Method
+Cui et al. (Cui and Wang, 2022) 2022 Method
+Lang et al. (Lang et al., 2023) 2023 Method
+Xu et al. (Xu and Ding, 2024) 2024 Method
+Miyai et al. (Miyai et al., 2024) 2024 Method
+Ours 2024 Task
 Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.
+Salehi et al.
+(2022) offer a comprehensive review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.
 Yang et al.
-(2021) discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based, and reconstruction-based methods.
+(2021) propose a unified framework to discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based and reconstruction-based methods.
+Concurrent with our work,  Miyai et al.
+(2024) discuss CLIP-based methods across related domains, while our survey specifically centers on OOD detection and provides a more comprehensive coverage of recent advances in this area.
 Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods.
 Additionally,  Lang et al.
-(2023) offer a review of OOD detection methods in natural language processing.
+(2023); Xu and Ding (2024) review various OOD detection methods in the context of natural language processing.
 However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios.
 Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method.
-Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022), there is an urgent need for a survey that incorporates the latest technologies.
+Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a; Geng et al., 2023; Gao et al., 2023a; Du et al., 2024a; Fan et al., 2022; Wang et al., 2024a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022; Nguyen et al., 2024; Fu et al., 2024; Dai et al., 2023; Liu et al., 2024a), there is an urgent need for a comprehensive survey that incorporates the latest technologies.
 {forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=13em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ OOD Detection ,fill=hidden-yellow!70,ver [ Problem: Training-driven
-OOD detection(§ 3),fill=hidden-red!70, align=center, text width=12em [ Approaches with only
-ID Data (§ 3.1),fill=hidden-blue!70,align=center [ Reconstruction-based, fill=hidden-blue!70 [ MoodCat  (Yang et al., 2022a), RAE (Zhou, 2022), MOOD  (Li et al., 2023a), MOODv2  (Li et al., 2024a), PRE  (Osada et al., 2023), LMD  (Liu et al., 2023a) DiffGuard  (Gao et al., 2023a),
-DenoDiff (Graham et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Probability-based, fill=hidden-blue!70 [ HVCM (Li et al., 2023b), DDR (Huang et al., 2022), LID (Kamkari et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Logits-based, fill=hidden-blue!70 [ LogitNorm  (Wei et al., 2022a), UE-NL  (Huang and Qiao, 2023), DML  (Zhang and Xiang, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ OOD Synthesis, fill=hidden-blue!70 [ ConfiCali (Lee et al., 2018a), CODEs  (Tang et al., 2021), CMG  (Wang et al., 2022a), VOS  (Du et al., 2022a), NPOS  (Tao et al., 2023), SHIFT  (Kwon et al., 2023), ATOL  (Zheng et al., 2023),
-SSOD  (Pei, 2024), SEM  (Yang et al., 2023b) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Prototype-based, fill=hidden-blue!70 [ SIREN  (Du et al., 2022b), CIDER  (Ming et al., 2023), PALM  (Lu et al., 2024), ReweightOOD  (Regmi et al., 2024), Step  (Zhou et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ A special case:
-Long-tail ID data, fill=hidden-blue!70 [ OLTR  (Liu et al., 2019), II-Mixup (Mehta et al., 2022), AREO  (Sapkota and Yu, 2023), IDCP (Jiang et al., 2023a), Open-Sampling  (Wei et al., 2022b), COOD  (Hogeweg et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Approaches with both ID
-and OOD Data (§ 3.2),fill=hidden-blue!70,align=center [ Boundary Regularization,fill=hidden-blue!70 [ OE  (Hendrycks et al., 2019), ELOC (Vyas et al., 2018), Why-ReLU (Hein et al., 2019), SSL-GOOD (Mohseni et al., 2020), EnergyOE  (Liu et al., 2020), MixOE  (Zhang et al., 2023a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Outlier Mining,fill=hidden-blue!70 [ BD-Resamp (Li and Vasconcelos, 2020), POEM  (Ming et al., 2022a), DAOL  (Wang et al., 2023a), DOE (Wang et al., 2023b), MixOE  (Zhang et al., 2023a), DivOE  (Zhu et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Imbalanced ID,fill=hidden-blue!70 [ PASCAL  (Wang et al., 2022b), COCL  (Miao et al., 2024), BERL (Choi et al., 2023), EAT  (Wei et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Training-agnostic
-OOD detection (§ 4),fill=hidden-red!70,align=center,text width=12em [ Post-hoc Approaches (§ 4.1),fill=hidden-blue!70,align=center [ Output-based,fill=hidden-blue!70 [ MSP  (Hendrycks and Gimpel, 2017), MaxLogits  (Hendrycks et al., 2022), Energy  (Liu et al., 2020), GEN  (Liu et al., 2023b), ZODE  (Xue et al., 2024), LogicOOD  (Kirchheim et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Distance-based,fill=hidden-blue!70 [ Mahalanobis  (Lee et al., 2018b), NNGuide  (Park et al., 2023a), KNN  (Sun et al., 2022), SSD  (Sehwag et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Gradient-based,fill=hidden-blue!70 [ Grad  (Lee and AlRegib, 2020), GradNorm  (Huang et al., 2021), GradOrth  (Behpour et al., 2023), GAIA  (Chen et al., 2023), OPNP  (Chen et al., 2024a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Feature-based,fill=hidden-blue!70 [ ODIN  (Liang et al., 2018a), ReAct  (Sun et al., 2021), VRA  (Xu et al., 2023), SHE  (Zhang et al., 2022), ViM  (Wang et al., 2022c), Neco  (Ammar et al., 2024), ASH  (Djurisic et al., 2023), NAC  (Liu et al., 2024a)
-Optimal-FS (Zhao et al., 2024), BLOOD  (Jelenić et al., 2024), SCALE  (Xu et al., 2024), DDCS  (Yuan et al., 2024), LINe  (Yong Hyun Ahn, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Density-based,fill=hidden-blue!70 [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Test-time Adaptive Approaches
-(§ 4.2) ,fill=hidden-blue!70,align=center [ Model-update needed,fill=hidden-blue!70 [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023b), SAL  (Du et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Model-update free,fill=hidden-blue!70 [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024), RTL  (Fan et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Large pre-trained
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=13em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ OOD Detection ,fill=hidden-red!70, draw=hidden-red!80!black,ver [ Problem: Training-driven
+OOD detection(§ 3),fill=hidden-blue!70, draw=hidden-blue!80!black, align=center, text width=12em [ Approaches with only
+ID Data (§ 3.1),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Reconstruction-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ MoodCat  (Yang et al., 2022), RAE (Zhou, 2022), MOOD  (Li et al., 2023a), MOODv2  (Li et al., 2024a), PRE  (Osada et al., 2023), LMD  (Liu et al., 2023a) DiffGuard  (Gao et al., 2023b),
+DenoDiff (Graham et al., 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Probability-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ HVCM (Li et al., 2023b), DDR (Huang et al., 2022a), LID (Kamkari et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Logits-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ LogitNorm  (Wei et al., 2022a), UE-NL  (Huang and Qiao, 2023), DML  (Zhang and Xiang, 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ OOD Synthesis, fill=hidden-blue!30, draw=hidden-blue!80!black, [ ConfiCali (Lee et al., 2018a), CODEs  (Tang et al., 2021), CMG  (Wang et al., 2022a), VOS  (Du et al., 2022a), NPOS  (Tao et al., 2023), SHIFT  (Kwon et al., 2023), ATOL  (Zheng et al., 2023),
+SSOD  (Pei, 2024), SEM  (Yang et al., 2023b), Forte  (Ganguly et al., 2025), HamOS  (Li and Zhang, 2025), POP  (Gong et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Prototype-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ Step  (Zhou et al., 2021), SIREN  (Du et al., 2022b), CIDER  (Ming et al., 2023), PALM  (Lu et al., 2024), ReweightOOD  (Regmi et al., 2024), AROS  (Mirzaei and Mathis, 2025), PFS  (Wu et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ A special case:
+Long-tail ID data, fill=hidden-blue!30, draw=hidden-blue!80!black, [ OLTR  (Liu et al., 2019), II-Mixup (Mehta et al., 2022), AREO  (Sapkota and Yu, 2023), IDCP (Jiang et al., 2023a), Open-Sampling  (Wei et al., 2022b), COOD  (Hogeweg et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] [ Approaches with both ID
+and OOD Data (§ 3.2),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Boundary Regularization,fill=hidden-blue!30, draw=hidden-blue!80!black, [ OE  (Hendrycks et al., 2019), ELOC (Vyas et al., 2018), Why-ReLU (Hein et al., 2019), SSL-GOOD (Mohseni et al., 2020), EnergyOE  (Liu et al., 2020a), MixOE  (Zhang et al., 2023a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Outlier Mining,fill=hidden-blue!30, draw=hidden-blue!80!black, [ BD-Resamp (Li and Vasconcelos, 2020), POEM  (Ming et al., 2022a), DAOL  (Wang et al., 2023a), DOE (Wang et al., 2023b), MixOE  (Zhang et al., 2023a), DivOE  (Zhu et al., 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Imbalanced ID,fill=hidden-blue!30, draw=hidden-blue!80!black, [ PASCAL  (Wang et al., 2022b), COCL  (Miao et al., 2024a), BERL (Choi et al., 2023), EAT  (Wei et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Training-agnostic
+OOD detection (§ 4),fill=hidden-blue!70, draw=hidden-blue!80!black,align=center,text width=12em [ Post-hoc Approaches (§ 4.1),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Output-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ MSP  (Hendrycks and Gimpel, 2017), MaxLogits  (Hendrycks et al., 2022), Energy  (Liu et al., 2020a), GEN  (Liu et al., 2023b), ZODE  (Xue et al., 2024), LogicOOD  (Kirchheim et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Distance-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ Mahalanobis  (Lee et al., 2018b), NNGuide  (Park et al., 2023a), KNN  (Sun et al., 2022), SSD  (Sehwag et al., 2021),Mahalanobis++ (Mueller and Hein, 2025a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Gradient-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ Grad  (Lee and AlRegib, 2020), GradNorm  (Huang et al., 2021), GradOrth  (Behpour et al., 2023), GAIA  (Chen et al., 2023), OPNP  (Chen et al., 2024a), PRO  (Chen et al., 2025a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Feature-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ODIN  (Liang et al., 2018a), ReAct  (Sun et al., 2021), VRA  (Xu et al., 2023), SHE  (Zhang et al., 2022), ViM  (Wang et al., 2022c), Neco  (Ammar et al., 2024), ASH  (Djurisic et al., 2023), NAC  (Liu et al., 2024b)
+Optimal-FS (Zhao et al., 2024), BLOOD  (Jelenić et al., 2024), SCALE  (Xu et al., 2024a), DDCS  (Yuan et al., 2024), LINe  (Yong Hyun Ahn, 2023), KANs  (Canevaro et al., 2025),
+CADRef  (Ling et al., 2025), ITP  (Xu and Yang, 2025), NCI  (Liu and Qin, 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Density-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] [ Test-time Adaptive Approaches
+(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-update needed,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-update free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
 model-based OOD detection
-(§ 5),fill=hidden-red!70,align=center,text width=12em [ Zero-shot Approaches
-(§ 5.1),fill=hidden-blue!70,align=center [ VLM-based,fill=hidden-blue!70 [ CLIPScope  (Fu et al., 2024), RONIN  (Nguyen et al., 2024), ZOC  (Esmaeilpour et al., 2022), MCM  (Ming et al., 2022b), CLIPN  (Wang et al., 2023c), NegLabel  (Jiang et al., 2023b),
-NegPrompt-C (Nie et al., 2024), LAPT  (Zhang et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ LLM-based,fill=hidden-blue!70 [ WK-LLM (Dai et al., 2023), ODPC  (Huang et al., 2024), LLM-OOD (Liu et al., 2024b) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Few-shot Approaches
-(§ 5.2),fill=hidden-blue!70,align=center [ Fine-tuning based, fill=hidden-blue!70 [ CLIP-OS  (Sun et al., 2024), EOK-Prompt  (Zeng et al., 2024), GalLop  (Lafon et al., 2024), ID-like  (Bai et al., 2024), DSFG  (Dong et al., 2023), LoCoOp  (Miyai et al., 2023a),
-NegPrompt  (Li et al., 2024b), TagOOD  (Li et al., 2024c), CRoFT (Zhu et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Meta-learning based, fill=hidden-blue!70 [ OOD-MAML  (Jeong and Kim, 2020), HyperMix  (Mehta et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Full-shot Approaches
-(§ 5.3),fill=hidden-blue!70,align=center [ NPOS  (Tao et al., 2023), PT-OOD (Miyai et al., 2023b), TOE  (Park et al., 2023b) , leaf, text width=48.5em,fill=hidden-blue!70,text width=50.6em ] ] ] ]
-Figure 1. Taxonomy of OOD detection problem scenarios and solutions.
-In this survey, we for the first time review the recent advances in OOD detection with a problem-oriented taxonomy, as illustrated in Fig.
+(§ 5),fill=hidden-green!70,draw=hidden-green!80!black,align=center,text width=12em [ Zero-shot Approaches
+(§ 5.1),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ VLM-based,fill=hidden-green!20,draw=hidden-green!80!black, [ CLIPScope  (Fu et al., 2024), RONIN  (Nguyen et al., 2024), ZOC  (Esmaeilpour et al., 2022), MCM  (Ming et al., 2022b), CLIPN  (Wang et al., 2023c), NegLabel  (Jiang et al., 2023b),
+NegPrompt-C (Nie et al., 2024), LAPT  (Zhang et al., 2024a) , AdaNeg (Zhang and Zhang, 2024), CSP (Chen et al., 2024b) , SimLabel (Zou et al., 2025) ,  (Jung et al., 2024) ,
+SeTAR (Li et al., 2024b) , OT-DETECTOR (Liu et al., 2025a) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] [ LLM-based,fill=hidden-green!20,draw=hidden-green!80!black, [ WK-LLM (Dai et al., 2023), ODPC  (Huang et al., 2024), LLM-OOD (Liu et al., 2024a), CMA (Lee et al., 2025), ReGuide (Kim et al., 2025) , COOD (Liu et al., 2024c),
+EOE (Cao et al., 2024) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] ] [ Few-shot Approaches
+(§ 5.2),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ Fine-tuning based, fill=hidden-green!20,draw=hidden-green!80!black, [ CLIP-OS  (Sun et al., 2024), Local-Prompt  (Zeng et al., 2025), GalLop  (Lafon et al., 2024), ID-like  (Bai et al., 2024a), DSFG  (Dong et al., 2023), LoCoOp  (Miyai et al., 2023a),
+NegPrompt  (Li et al., 2024c), SUPREME (Wang et al., 2025a) , SCT (Yu et al., 2024a) , GaCoOp (Tong et al., 2025) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] [ Meta-learning based, fill=hidden-green!20,draw=hidden-green!80!black,, [ OOD-MAML  (Jeong and Kim, 2020), HyperMix  (Mehta et al., 2024) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] ] [ Full-shot Approaches
+(§ 5.3),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ NPOS  (Tao et al., 2023), PT-OOD (Miyai et al., 2023b), TOE  (Park et al., 2023b) , leaf, text width=48.5em,fill=hidden-green!20,draw=hidden-green!80!black,,text width=50.6em ] ] ] ]
+Figure 1.
+Taxonomy of OOD detection problem scenarios and solutions.
+The first two categories (in blue) correspond to shallow model-based methods, while the third category (in green) highlights approaches based on foundation models (large pre-trained models).
+Different colors are used to visually distinguish the traditional shallow-model-based solutions from the newly emerging foundation model-based methods, emphasizing the conceptual differences between them.
+In this survey, we for the first time review the recent advances in OOD detection with a task-oriented taxonomy, as illustrated in Fig.
 1.
 Based on whether the method needs to control the pre-training process, we categorize OOD detection algorithms into training-driven and training-agnostic methods.
-Considering the rapid development of large pre-trained models nowadays, we also regard large pre-trained model-based OOD detection as a separate section.
+Considering the rapid development of large pre-trained models nowadays and the evolving definition of OOD under such models, we also regard large pre-trained model-based OOD detection as a separate section.
 In detail, training-driven methods achieve high detection capability by designing the optimization process of the training stage.
@@
 We discuss them in terms of whether they have access to a few examples, including zero-shot, few-shot and full-shot scenarios.
-The remainder of this survey is organized as follows.
-We recap the related work of OOD detection in Sec.
+The structure of this survey is as follows.
+We discuss the related topics of OOD detection in Sec.
 2.
@@
 7 to shed light on underexplored and potentially critical avenues.
-Finally, we conclude this survey in Sec.
-8.
 2. Related work
 Anomaly Detection.
-Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009).
-These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003), or system failures  (Du et al., 2017).
+Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009; Xing et al., 2025; Wang and Nalisnick, 2025).
+These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003) or system failures  (Du et al., 2017).
 The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations.
 Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency.
-While both AD and OOD detection aim to identify unusual or unexpected data points, AD focuses on deviations within a single distribution (typically the training set), whereas OOD detection addresses differences between the training distribution and new, unseen inputs.
+While both AD and OOD detection aim to identify unusual or unexpected data points, compared to OOD detection, AD differs in three key aspects.
+First, regarding supervision, AD methods are typically unsupervised, training only on normal samples without detailed labels and detecting anomalies during testing, whereas OOD detection involves labeled category discrimination.
+Second, AD focuses on perceptual anomalies, where images belong to known categories but contain anomalous regions, while OOD detection targets semantic differences, with test samples from unseen categories.
+Additionally, although both methods commonly use AUROC (Wehler et al., 2001) for evaluation, AD specifically employs AUPRO (Bertoldo et al., 2024) to assess pixel-level detection performance.
 Novelty Detection.
@@
 Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016).
-This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that were not present during training.
+This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that are not present during training.
 Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully.
@@
 While both methods manage uncertainty, selective classification deals with confidence in making predictions on ID data (El-Yaniv et al., 2010; Chang et al., 2024), and OOD detection focuses on recognizing and handling data that is not represented in the training set.
+Misclassification (Failure) Detection.
+Misclassification detection aims to identify incorrect predictions on in-distribution data, exposing a model’s unreliability in its intended domain (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d).
+This is especially important in safety-critical areas such as autonomous driving and healthcare.
+Failure detection generalizes this task to capture a broader range of model breakdowns, including uncertainty and data corruption.
+While both misclassification and OOD detection seek to improve model reliability, the former focuses on errors within the training distribution (Jaeger et al., 2023), whereas the latter targets inputs outside it.
 3. Problem: Training-driven OOD Detection
@@
 Overview.
-Given the ID data, approaches in this section train a model on them and aim to utilize the model for detecting OOD test samples while ensuring accurate classification of the ID data.
-Approaches with only ID data focus specifically on mining information from ID data, without explicitly relying on other information from real-world OOD data.
-We further differentiate these methods into the following four categories: Reconstruction-based, Probability-based, Logits-based, OOD Synthesis, and Prototype-based.
+Given the ID data, approaches in this section train a model on them and aim to utilize the model for detecting OOD test samples while ensuring classification performance of the ID data.
+They focus specifically on mining information from ID data, without explicitly relying on other information from real OOD data.
+We further differentiate these methods into the following five categories: Reconstruction-based, Probability-based, Logits-based, OOD Synthesis, and Prototype-based.
 Considering real-world requirements, we also delve into a specific scenario: Long-tail ID data.
@@
 Reconstruction-based.
-The reconstruction-based methodology offers a new research avenue by scrutinizing the discrepancy between sample representations before and after reconstruction, which relies heavily on the reconstruction performance of models.
-Fundamentally, the objective of the reconstruction task is to recuperate the inherent semantic content of the dataset, guided by a designated supervisory signal.
-This premise rests on the notion that OOD data intrinsically possess semantic characteristics that are incongruent with those of ID label information.
-By quantitatively assessing the extent of this semantic deviation, the model becomes equipped to discern OOD data with a high degree of precision.
-The approximate process of such methods is shown in Fig.
-3.
-MoodCat  (Yang et al., 2022a) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process.
+The reconstruction-based approach detects OOD samples by measuring the difference between representations of data before and after reconstruction, relying on the model’s ability to accurately recover input semantics.
+The core idea is that OOD data contain semantic features that differ from those of ID data.
+By quantifying these differences, the model can effectively distinguish OOD samples.
+An overview is illustrated in Fig.
+3 (a).
+MoodCat  (Yang et al., 2022) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process.
 Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data.
@@
 Figure 3.
-Illustration of reconstruction-based OOD detection approaches.
-After extracting the feature representation $F(x)$ of the image through a neural network, it is then fed into a reconstruction network (such as VAE or DDPM) to obtain the reconstructed image representation $F^{\prime}(x)$.
-By comparing the dissimilarity between the original and reconstructed images, we can identify OOD data, as OOD samples typically exhibit greater dissimilarity before and after reconstruction.
-Recently, diffusion models have made remarkable strides in both training stability and the quality of generated images.
+Illustration of some subclasses within the Training-driven approaches: (a) Reconstruction-based : After extracting the feature representation $F(x)$ of the image, it is fed into a reconstruction network (e.g., VAE or DDPM) to obtain the reconstructed feature representation $F^{\prime}(x)$.
+By comparing the dissimilarity between the original and reconstructed images, OOD data can be identified, as OOD samples typically exhibit higher dissimilarity before and after reconstruction.
+(b) Prototype-based : The model learns more optimal ID prototypes during the training process to distinguish OOD data.The symbol $h$ typically denotes the features of the penultimate layer in the network.
+(c) OOD Synthesis : During training, real OOD data is not available; instead, OOD data is synthesized by extrapolating information from the ID data.
+(d) Outlier Mining : During training, real OOD data is available, but typical OOD samples are generated using this real OOD data.
+(e) Boundary Regularization : The model’s decision boundary is optimized by using real OOD data, typically without involving the synthesis of OOD data.
+Recently, diffusion models (Ho et al., 2020) have made remarkable strides in both training stability and the quality of generated images.
 Leveraging diffusion models to detect OOD data has become a new research direction.
@@
 LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold.
-DiffGuard  (Gao et al., 2023a) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
+DiffGuard  (Gao et al., 2023b) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
 Probability-based.
-Research in probability-based direction aims to establish probability models to describe the distribution of training data.
-Through these probability models, suitable scoring functions are developed to compute the scores of test samples within the ID distribution, which reflect whether these samples belong to the ID distribution.
+Probability-based approaches aim to establish probability models to describe the distribution of training data.
 In this domain,  Li et al.
-(2023b) model each ID category during training using multiple Gaussian mixture models, and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.
+(2023b) model each ID category during training using multiple Gaussian mixture models (Rasmussen, 1999), and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.
 Huang et al.
-(2022) address this issue by introducing two regularization constraints.
+(2022a) address this issue by introducing two regularization constraints.
 The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples.
@@
 Logits-based.
-These algorithms primarily focus on the predictions of neural networks, particularly on Logits, which are the results of the neural network’s output layer.
+These methods detect OOD samples by analyzing the logits (raw outputs before softmax) from the neural network’s final layer.
 Logits typically represent the model’s confidence or probability for each category.
-LogitNorm  (Wei et al., 2022a) proposes a method using logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence.
+LogitNorm  (Wei et al., 2022a) uses logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence.
 Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores.
@@
 OOD Synthesis.
-In the task of OOD detection, it is theoretically believed that incorporating the features of OOD data during model training can enhance the OOD detection performance.
-Due to the challenges in acquiring distribution information for OOD samples, some methods employ ID data to estimate the distribution of OOD data.
-This is done to simulate the scenario in which a model encounters OOD data in real-world situations.
-Figure 4.
-Illustration of OOD synthesis approaches.
-In the absence of real-world OOD data, utilizing extrapolation from ID data to identify suitable OOD data and then training the model using the method of exposing anomalies enhances the capability of OOD detection.
-The orange dots in the figure represent virtual OOD data synthesized using ID data.
+In the task of OOD detection, incorporating the features of OOD data during model training can enhance the OOD detection performance by allowing the model to better recognize and differentiate OOD data.
+Due to the challenges in acquiring distribution information for OOD samples, some methods employ ID data to estimate the distribution of OOD data, simulating real-world scenarios where a model encounters OOD data, as shown in Fig.
+3 (c).
 Lee et al.
@@
 It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention.
-Noticing the additional distributional assumptions posed by previous methods for OOD data, Some researches approach the problem from another perspective.
+Noticing the additional distributional assumptions posed by previous methods for OOD data, some researchers approach the problem from another perspective.
 For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality.
@@
 It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data.
+Forte (Ganguly et al., 2025) utilizes per-point statistics (such as precision, recall, and density) to capture the local geometric properties of the data manifold in the feature space and performs typicality estimation using density-based non-parametric methods (such as Gaussian mixture models).
+This approach helps improve the accuracy of OOD synthetic data detection.
 SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples.
@@
 Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances.
+In recent studies, HamOS (Li and Zhang, 2025) employs Hamiltonian Monte Carlo (Duane et al., 1987) sampling, treating the synthesis process as sampling from a Markov chain.
+In the absence of OOD data, it synthesizes high-quality OOD virtual anomaly points by constructing a Markov chain (Neal, 1993).
+It projects the features onto a hyperspherical feature space and uses K-nearest neighbor distances to guide the generation of anomaly points.
+POP (Gong et al., 2025) does not require the synthesis of specific outliers, but instead directly reshapes the decision boundaries through virtual OOD prototypes, improving the model’s OOD detection capability by penalizing misclassifications.
 Prototype-based.
 During the model training process, prototype-based OOD detection methods aim to model the ID data using prototypes to learn the common distribution characteristics of the ID data.
-In the testing phase, the model measures the differences between the sample and class-level prototypes to determine the category of the sample.
-The general procedure of these methods is roughly illustrated in Fig.
-5.
-Figure 5.
-Illustration of prototype-based OOD detection approaches.
-The algorithm for OOD detection based on prototype learning roughly consists of three network structures: 1) Encoder.
-This structure is responsible for propagating sample data through a neural network to obtain feature representations $h$.
-2) Projector.
-Structures of this type project the image features encoded in $h$ to a new representation space.
-3) Prototype Learning.
-In the new representation space, prototypes representing each class are learned, with each class learning its corresponding prototype center.
-During testing, the likelihood of anomalous data is measured by comparing the similarity between test samples and prototypes of each ID class.
-SIREN  (Du et al., 2022b) first models the distribution of ID data using the von Mises-Fisher(vMF) model, which enables representing each class as a compact cluster, i.e., a class prototype.
+For OOD data, there is a significant difference between the sample features and the ID prototype.
+In the testing phase, the model determines the sample’s category by measuring the difference between the sample and the ID prototype.
+The general procedure of these methods is outlined in Fig.
+3 (b).
+SIREN  (Du et al., 2022b) first models the distribution of ID data using the von Mises-Fisher(vMF) (Mardia and Jupp, 2009) model, which enables representing each class as a compact cluster, i.e., a class prototype.
 Generally, the vMF distribution modeling formula can be represented as: $p_{D}(z;p_{k},k)=Z_{D}(k)exp(kp_{k}^{T}z)$, where $p_{k}$ is the $k$-th prototype with unit norm, $\kappa\geq 0$ represents the concentration around the mean, and $Z_{D}(\kappa)$ is the normalization factor.
@@
 The optimization process during model training pertains to the prototypes of classes.
+AROS (Mirzaei and Mathis, 2025) focuses more on the feature separation between ID and OOD data.
+It proposes using the Lyapunov stability theorem to constrain the stability of the neural network’s embedding space, ensuring that the embeddings of ID and OOD data converge to distinct stable equilibrium points.
+To obtain OOD data, it also employs a technique of sampling from the low-probability regions of the ID data to generate pseudo-OOD.
+CNC (Harun et al., 2025) proposed controlling the degree of neural collapse (NC) at different layers of the DNN network.
+In the Encoder, entropy regularization is introduced to mitigate NC, while in the Projector, the projection head is used to enhance the NC phenomenon, forming more compact prototype and thereby improving OOD detection performance.
 ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores.
@@
 By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function.
+PFS (Wu et al., 2025) also introduces a clustering loss that encourages features of the same ID class to cluster more closely, thereby improving the consistency of ID data representation.
 A special case: Long-tail ID data.
@@
 The primary focus in such problems is on optimizing the model’s decision boundary, rather than the OOD data itself.
-Due to the introduction of real OOD information, the boundary of these two categories will be accurately calculated.
+Due to the introduction of real OOD information, the decision boundary between ID and OOD data is further accurately calculated.
 Boundary Regularization.
@@
 The central idea of this method can be grasped from Fig.
-6.
-Figure 6.
-Illustration of boundary regularization approaches.
-The key to model optimization strategy lies in the decision boundary of the model classification.
-Boundary regularization utilizes existing known ID data (the colored dots) and some OOD data (black dots) to optimize the potential decision boundary, compacting its classification boundary with abnormal data as tightly as possible without affecting the classification of ID data.
+3 (e).
 Specifically, given a model $f$ and the original loss function, the model training process aims to minimize the objective:
@@
 $\mathcal{L}_{\text{OE}}$ represents the cross-entropy loss with respect to $U$.
+Here, $\mathcal{D}_{\text{in}}$ denotes the distribution of ID data, while $\mathcal{D}_{\text{out}}^{\text{OE}}$ denotes the distribution of OOD.
 The fundamental purpose is to compel the model to optimize the OOD data distribution to a uniform distribution, a principle that is universal in OE-type approaches.
@@
 This design can utilize a maximum softmax probability baseline  (Hendrycks and Gimpel, 2017) detector to detect anomalous data.
-Compared to traditional softmax scores, EnergyOE  (Liu et al., 2020) builds upon OE by leveraging energy scores for better discrimination between ID and OOD samples, and it is less prone to issues of overconfidence.
+Compared to traditional softmax scores, EnergyOE  (Liu et al., 2020a) builds upon OE by leveraging energy scores for better discrimination between ID and OOD samples, and it is less prone to issues of overconfidence.
 Specifically, its calculation formula:
@@
 The main process is depicted in Fig.
-7.
-Figure 7.
-Illustration of outlier mining approaches.
-Due to the limitation of whether real OOD data exists, methods like anomaly mining allow the model to access real OOD data during training.
-The black dots in the figure represent real OOD data sampled from other datasets, which do not overlap with the ID categories.
-Utilizing such OOD data to extrapolate more representative OOD data (orange dots) is a solution to this type of problem.
+3 (d).
 POEM  (Ming et al., 2022a) raises methodology leans towards excavating anomalies that carry more representative characteristics.
@@
 Through model perturbation, the data is implicitly transformed, and the model continues learning from this perturbed data to improve its robustness.
-Imbalance.
+Long-tail ID data.
 Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information.
 The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two.
-To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024) incorporates a learnable tail class prototype during the training process.
+To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024a) incorporates a learnable tail class prototype during the training process.
 This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.
@@
 4. Problem: Training-agnostic OOD Detection
-Training-agnostic OOD detection places a primary emphasis on adaptation strategies during test time, as opposed to the focus on classifier performance seen in training-driven OOD detection.
+Training-agnostic OOD detection focuses on adaptation strategies at test time for models that already possess downstream classification capabilities, as opposed to the focus on classifier performance seen in training-driven OOD detection.
 Based on whether they rely on dependencies among test data, methods are categorized into two types: post-hoc and test-time adaptive approaches, as illustrated in Fig.
-8.
+4.
 Post-hoc methods compute results for individual samples independently, unaffected by changes in other samples.
-In contrast, test-time adaptive methods exhibit dependencies among test samples.
+In contrast, test-time adaptive methods leverage information shared among test samples to enhance OOD detection.
 Figure 4.
@@
 Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing.
-“A’/B’ ” means that the original features are deformed; Bank means that some samples are stored; Clustering means that clustering is performed for ID images.
+“A’/B’ ” means that the original features are deformed; “Bank” means that some samples are stored; “Clustering” means that clustering is performed for ID images.
 4.1. Post-hoc OOD Detection Approaches
@@
 Recent work on this type of problem has some recent progress.
-A summary of the key factors involved in such methods is given in Table  1.
+A summary of the key factors involved in such methods is given in Table  2.
 Table 2. Comparison of key components in OOD detection methods.
-Method Type Space
+Type Method Space
 feature logit gradient probability
-MSP (Hendrycks and Gimpel, 2017) Output-Based $\checkmark$
+Output-Based MSP (Hendrycks and Gimpel, 2017) $\checkmark$
 Maxlogits (Hendrycks et al., 2022) $\checkmark$
-Energy (Liu et al., 2020) $\checkmark$
+Energy (Liu et al., 2020a) $\checkmark$
 GEN (Liu et al., 2023b) $\checkmark$ $\checkmark$
-Mahalanobis (Lee et al., 2018b) Distance-Based $\checkmark$
+Distance-Based Mahalanobis (Lee et al., 2018b) $\checkmark$
 NNGuide (Park et al., 2023a) $\checkmark$
@@
 SSD (Sehwag et al., 2021) $\checkmark$
-Grad (Lee and AlRegib, 2020) Gradient-Based $\checkmark$
+Gradient-Based Grad (Lee and AlRegib, 2020) $\checkmark$
 GradNorm (Huang et al., 2021) $\checkmark$ $\checkmark$ $\checkmark$
@@
 OPNP (Chen et al., 2024a) $\checkmark$ $\checkmark$
-ODIN (Liang et al., 2018a) Feature-Based $\checkmark$ $\checkmark$ $\checkmark$
+Feature-Based ODIN (Liang et al., 2018a) $\checkmark$ $\checkmark$ $\checkmark$
 ReAct (Sun et al., 2021) $\checkmark$
 VRA (Xu et al., 2023) $\checkmark$
-SHE (Zhang et al., 2022) $\checkmark$
 Vim (Wang et al., 2022c) $\checkmark$ $\checkmark$
@@
 ASH (Djurisic et al., 2023) $\checkmark$
-NAC (Liu et al., 2024a) $\checkmark$
 Optimal-FS (Zhao et al., 2024) $\checkmark$ $\checkmark$
-BLOOD (Jelenić et al., 2024) $\checkmark$
-SCALE (Xu et al., 2024) $\checkmark$
-ConjNorm (Peng et al., 2024) Density-based $\checkmark$ $\checkmark$
+SCALE (Xu et al., 2024a) $\checkmark$
+Density-Based ConjNorm (Peng et al., 2024) $\checkmark$ $\checkmark$
 GEM (Morteza and Li, 2022) $\checkmark$ $\checkmark$ $\checkmark$
@@
 Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others.
-MSP  (Hendrycks and Gimpel, 2017) was the first to employ the maximum softmax value to validate OOD detection effectiveness.
+MSP  (Hendrycks and Gimpel, 2017) is the first to employ the maximum softmax value to validate OOD detection effectiveness.
 For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category.
-In contrast to the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by comparing the maximum logit value in the logits vector output by the neural network.
-Logits, representing the model’s confidence in each category, are the neural network’s output before the softmax layer, without undergoing softmax transformation.
-Meanwhile, Energy  (Liu et al., 2020) introduces the Helmholtz Free Energy, which theoretically aligns with the input probability density and is less susceptible to overconfidence issues.
+Departing from the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by utilizing the maximum value of the logits vector generated by the neural network.
+These logits constitute the raw, pre-Softmax layer outputs of the neural network, representing the model’s confidence in each category without undergoing Softmax transformation.
+Alternatively, the Energy  (Liu et al., 2020a) method defines an energy score conceptually derived from Helmholtz Free Energy.
+This score can theoretically reflect the input probability density.
+By assigning lower energy scores to ID data, this method distinguishes them from anomalous data.
+Additionally, it demonstrates reduced susceptibility to model overconfidence issues.
 GEN  (Liu et al., 2023b) introduces the concept of generalized entropy and directly utilizes Bregman divergence to compute the statistical distance between the model’s probability output and uniform distribution, aiming to identify OOD data.
-Additionally, leveraging sufficient prior knowledge might be a viable solution.
-ZODE  (Xue et al., 2024) perform predictions on samples across multiple pre-trained models simultaneously to determine whether multiple models can identify OOD samples, using this as a basis to distinguish between data.
-LogicOOD  (Kirchheim et al., 2024) present a novel approach that uses first-order logic for knowledge representation to perform OOD detection.
-This reasoning system uses prior knowledge to infer whether an input is consistent with the prior knowledge about the training distribution.
-It is particularly user-friendly in terms of interpretability, as it allows for comparing the output of samples against a knowledge base to determine whether they belong to the ID data.
+In addition, leveraging sufficient prior knowledge might be a viable solution.
+ZODE  (Xue et al., 2024) performs predictions on samples across multiple pre-trained models simultaneously to determine whether multiple models can identify OOD samples, using this as a basis to distinguish between data.
+LogicOOD  (Kirchheim et al., 2024) presents a novel approach that uses first-order logic for knowledge representation to perform OOD detection.
+This reasoning system uses prior knowledge to infer whether an input is consistent with prior knowledge about the training distribution.
+It is particularly user-friendly in terms of interpretability, as it allows comparing the output of samples against a knowledge base to determine whether they belong to the ID data.
 Distance-based.
@@
 During testing, it calculates the Mahalanobis distance between the feature vector and the mean of each class.
+Mahalanobis++ (Mueller and Hein, 2025b) also uses Mahalanobis distance, but applies L2 norm normalization to the features in the early stages of calculation to eliminate the instability in feature extraction across different models, which affects the distribution assumption they rely on.
 SSD  (Sehwag et al., 2021) essentially utilizes the Mahalanobis distance.
@@
 Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability.
+In recent research, PRO (Chen et al., 2025a) introduces an adversarial score, with the core idea being to perturb the input data using gradient descent to search for the local minimum score near the original input.
+This method effectively lowers the confidence of OOD samples and is compatible with existing MSP detection methods during score calculation, thereby enhancing detection performance.
+S&I (Zhang et al., 2025a) identifies the issue of insufficient sample gradient feature discrimination in this type of method.
+By introducing adversarial samples and integrating gradients along the attribution path, it achieves a more accurate feature distinction strategy.
 Feature-based.
@@
 It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data.
+KANs (Canevaro et al., 2025) also focus on the neuron activation states.
+Due to the local neuroplasticity characteristic of Kolmogorov-Arnold Networks (Liu et al., 2025b), KANs distinguish between ID and OOD data by comparing the activation response differences between the trained and untrained KANs.
 SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison.
@@
 It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs.
+CADRef (Ling et al., 2025) decouples the features by using a strategy based on the sign alignment between the relative features and model weights, dividing the sample features into positive and negative errors.
+Intuitively, the positive and negative error components influence the final output logits, which helps in identifying OOD images from this perspective.
+ITP (Xu and Yang, 2025) optimizes the model by quantifying the contribution of model parameters to ID data prediction, specifically by evaluating the contribution of each parameter through its partial derivatives and removing those with low contribution.
+Additionally, during the testing phase, ITP uses a right-tail Z-score test to assess whether any parameter exhibits overconfidence in the classification of a given test sample, further refining the model’s decision-making.
 Feature shaping is emerging as a prominent area of research focus.
@@
 By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection.
-Its computation formula can be expressed in the following form:
-(5) $-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
+Its computation of this score $S$ can be expressed in the following form:
+(5) $S=-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
 where $\alpha$ is a scaling constant, computed by the model.
@@
 And it have $Wz^{P^{\perp}}=0$.
-Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020), and $f(z)$ represents the logit output of the model.
+Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020a), and $f(z)$ represents the logit output of the model.
 The first term here represents virtual logits, while the second term represents the score of the energy function.
@@
 The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data.
+Similarly, NCI (Liu and Qin, 2023) studies the phenomenon of neural collapse and discovers that during training, ID samples tend to cluster around class weight vectors in the penultimate feature space, while OOD samples do not exhibit this clustering behavior.
+Therefore, NCI leverages the proximity between features and class weight vectors to effectively distinguish ID samples from OOD samples, while maintaining low computational cost.
 ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted.
-NAC  (Liu et al., 2024a) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data.
+NAC (Liu et al., 2024b) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data.
 By quantifying this statistical property, NAC aims to distinguish between ID and OOD data.
@@
 In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH.
-SCALE  (Xu et al., 2024) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data.
+SCALE  (Xu et al., 2024a) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data.
 Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.
@@
 Overview.
-Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize test data, either the complete test set or a series of unlabeled mini-batches, to enhance OOD detection performance through model adaptation.
-Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge is impossible.
+Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize unlabeled test data, either the complete test set or a series of mini-batches, to enhance OOD detection performance through model adaptation.
+Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge can be challenging and potentially limited.
 These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free.
@@
 Model-optimization-based. Model optimization-based methods enhance the trained model by leveraging unlabeled data during the post-training phase.
-A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection.
+A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024a) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection.
 The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them.
-Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024) explains how they help OOD detection from the lens of separability and learnability.
+Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024a) explains how they help OOD detection from the lens of separability and learnability.
 Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model.
@@
 The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification.
-In contrast, ATTA  (Gao et al., 2023b) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model.
+In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model.
 SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
@@
 Modifying the original trained model is infeasible in certain security-sensitive scenarios.
-Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” (MOF) techniques, are increasingly garnering interest.
+Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” techniques, are increasingly garnering interest.
 These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
-Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself.
+Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself.
 ETLT observes a linear correlation between the feature representation and the OOD score of a test input.
@@
 In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach.
-The core principle behind AdaOOD is the maintenance of a memory bank, similar to the approach taken by AUTO.
+Similarly, OODD (Yang et al., 2025) constructs a dynamic OOD dictionary, managed by a priority queue, which continuously updates with feature representations of test samples based on their OOD scores.
+The core principle behind AdaOOD and OODD is the maintenance of a memory bank, similar to the approach taken by AUTO.
 Online v.s.
@@
 Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment.
-In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023b) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
+In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023a) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
 More challenging scenario.
-In the context of test-time OOD detection scenarios, some scholars have proposed more challenging configurations that demand a higher level of capability from the models.
-MOL (Wu et al., 2023) introduces a more realistic problem scenario, namely Continuous Adaptive Out-of-Distribution (CAOOD) detection, aimed at addressing the challenge of constantly changing ID and OOD distributions in the real world.
+In the context of test-time OOD detection scenarios, some researchers have proposed more challenging configurations that demand a higher level of capability from the models.
+MOL (Wu et al., 2023a) introduces a more realistic problem scenario, namely Continuous Adaptive Out-of-Distribution (CAOOD) detection, aimed at addressing the challenge of constantly changing ID and OOD distributions in the real world.
 The meta-learning approach is employed to swiftly adapt models in response to the complexities encountered in various scenarios in MOL.
-Figure 9. Overview and trends in LPM-based OOD Detection Methods.
+Figure 5.
+Overview and trends in LPM-based OOD Detection Methods.
+(a) presents statistical data on influential research works in recent years, which involves papers indexed in prominent journals or conferences in the field.
+In the bar chart, the lower and upper parts represent the total number of papers in the first and second halves of each year, respectively.
+(b) shows a trend pie chart illustrating the proportion of keyword categories related to the research works in that year.
 5. Problem: Large Pre-trained Model-based OOD Detection
@@
 Recent research  (Vaze et al., 2022) highlights a correlation between higher ID classification accuracy and better OOD detection performance.
-Consequently, large pre-trained model-based OOD detection problem comes naturally.
-In recent years, large pre-trained models of various types, including single-modal (ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), Diffusion (Ho et al., 2020)), visual language models (VLMs) (CLIP (Radford et al., 2021), multi-modal Diffusion (Saharia et al., 2022) ALIGN (Jia et al., 2021), ), and large language models (LLMs) (GPT3 (Brown et al., 2020)), have been increasingly utilized for OOD detection tasks, as shown in Fig.
-9.
+Consequently, large pre-trained model-based (LPM-based) OOD detection problem comes naturally.
+In recent years, large pre-trained models of various types, including single-modal models (ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), Diffusion (Ho et al., 2020)), visual language models (VLMs) (CLIP (Radford et al., 2021), multi-modal Diffusion (Saharia et al., 2022), ALIGN (Jia et al., 2021), ), and large language models (LLMs) (GPT3 (Brown et al., 2020)), have been increasingly utilized for OOD detection tasks, as shown in Fig.
+5.
+Based on the trend chart, it can be anticipated that LPM-based OOD Detection will continue to be a key area of focus in the future.
 Leveraging the powerful representational capabilities of large pre-trained models has further relaxed the constraints of OOD detection tasks, leading to a focus on more challenging and realistic scenarios, which has emerged as a new hotspot.
-Given the number of ID shots exposed to the large pre-trained model, large pre-trained model-based OOD detection can be classified into Zero-shot, Few-shot, and Full-shot OOD detection, as shown in Fig.
-10.
-The performance evaluations of several relevant competitive methods are summarized in Table  2 to provide an understanding of the performance level of OOD detection in this area.
+Given the number of ID shots exposed to the large pre-trained model, OOD detection in this section can be classified into Zero-shot, Few-shot, and Full-shot OOD detection, as shown in Fig.
+6.
+The performance evaluations of several relevant competitive methods are summarized in Table  3 to provide an understanding of the performance level of OOD detection in this area.
 Figure 6.
@@
 Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data.
-Note that we do not need to have access to ID images and rely solely on textual category knowledge.
+Note that we rely solely on textual category knowledge, without the need for access to ID images.
 It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training.
-VLMs-based.
-Existing research on zero-shot OOD detection using VLMs can be categorized into two main approaches based on which VLMs are utilized as a basic backbone.
+Existing research on zero-shot OOD detection using vision-language models (VLMs-based) can be broadly categorized into two main approaches, depending on the underlying models employed: diffusion-based and CLIP-based methods.
 Diffusion-based.
-Traditional generative approaches, such as diffusion models  (Ho et al., 2020), capture the ID distribution effectively.
-Subsequently, these methods are used to identify OOD samples by assessing the likelihood that a given test input is derived from an OOD source.
-In a recent study in RONIN  (Nguyen et al., 2024), the diffusion model has been applied to achieve OOD object detection, which utilizes the diffusion model to generate ID inpainting, while the CLIP model is employed to calculate the similarity score.
+Recent generative approaches, such as diffusion models (Ho et al., 2020), effectively capture the ID data by modeling its underlying distribution.
+These models typically detect OOD samples by evaluating the likelihood of a test input belonging to the modeled ID distribution.
+For instance, RONIN (Nguyen et al., 2024) leverages a diffusion model for ID inpainting, combined with the CLIP model to compute similarity scores for OOD detection.
+However, most existing works focus more extensively on CLIP-based methods due to their superior zero-shot capabilities and flexibility.
 CLIP-based.
-Given the remarkable proficiency of CLIP in correlating texts and images, a substantial number of researchers have ventured into zero-shot OOD detection leveraging CLIP.
-Typically, in the zero-shot OOD detection setup, add-ons are added to the pre-trained model to better adapt it to the OOD detection task.
-ZOC  (Esmaeilpour et al., 2022) achieves zeto-shot OOD detection by training an image description generator on a large image captioning dataset  (Lin et al., 2014), enabling the model to generate candidate unseen labels.
-It’s essential to note that ZOC treats CLIP merely as a feature extractor and doesn’t impart OOD detection capabilities to CLIP itself.
-The subsequent work CLIPN  (Wang et al., 2023c) empowers CLIP with the ability to say “no” by a “no-prompt” encoder.
-Other approaches alleviate the need for additional training and focus on enhancing the performance of post-hoc zero-shot OOD detection.
-A simple baseline of clip-based zero-shot OOD detection is to use the normalized text-image similarity as an OOD score.
-Ming et al.
-(2022b) further substitute the score with the maximum concept matching (MCM) score and offer a comprehensive theoretical explanation of this modification.
-Moreover, NegLabel  (Anonymous, 2024), LAPT (Zhang et al., 2024) and CLIPScope (Fu et al., 2024) enriches the text information through a predetermined corpus.
-Concretely, NegLabel selects negative words (labels) from the corpus, which are dissimilar to ID labels, to enhance OOD detection performance.
-Futhermore, CLIPScope uses those labels to revise the original score by applying Bayesian rules.
-In contrast, LAPT leverages text-to-image generation models or image retrieval models to obtain images corresponding to the given ID and Negative ID labels, followed by prompt tuning.
-It is noteworthy that ZOC, CLIPN, NegLabel and CLIPScope each devise unique OOD scores for integrating into their respective models.
-Conversely, the MCM score stands out for its generalizability across different models.
-The calculation of MCM is as follows:
-(6) $S_{MCM}(x^{\prime};\mathcal{Y}_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}},$
-where $x^{\prime}$ is the test input image; $\mathcal{Y}_{in}=\{y_{1},y_{2},\ldots,y_{K}\}$ represents the label set; $y_{i}$ denotes the $i$-th label in the label set; $T(t_{i})$ corresponds to the concept vector for the text prompt $t_{i}$; $s_{i}(x^{\prime})$ is the cosine similarity score between the image feature $I(x^{\prime})$ and the concept vector $T(t_{i})$, computed as $s_{i}(x^{\prime})=\frac{I(x^{\prime})\cdot T(t_{i})}{\|I(x^{\prime})\|\|T(t_{i})\|}$; $e^{s_{i}(x^{\prime})/\tau}$ is the exponentiated similarity score normalized by the temperature parameter $\tau$; $K$ is the number of labels in the label set; $\tau$ is the temperature parameter of the softmax function; and $S_{MCM}$ is the maximum concept matching score.
-Some works have explored variations of zero-shot OOD detection or more challenging scenarios, such as detecting ID objects (Miyai et al., 2023c) or handling labels with noise (Ding and Pang, 2024), which will be discussed in the Section.7.
+In contrast to diffusion-based approaches, CLIP-based methods have received significant attention for zero-shot OOD detection.
+These methods exploit the powerful vision-language alignment capability of CLIP to perform OOD detection without requiring additional training on OOD data.
+In the following, we focus on the development and evolution of CLIP-based zero-shot OOD detection methods, discussing both approaches that require extra training and post-hoc methods, as well as recent advances leveraging textual information.
+A representative approach, ZOC (Esmaeilpour et al., 2022), introduces an image description generator trained on large-scale image captioning datasets (Lin et al., 2014), enabling the generation of candidate unseen labels for OOD detection.
+Notably, ZOC treats CLIP primarily as a feature extractor and does not inherently endow it with OOD discrimination capabilities.
+Building upon this, CLIPN (Wang et al., 2023c) enhances CLIP’s expressiveness by incorporating a “no-prompt” encoder, thereby empowering CLIP to explicitly reject unfamiliar inputs—a critical step toward robust OOD detection in open-world scenarios.
+Beyond methods that require further training, a parallel line of research investigates post-hoc zero-shot OOD detection, aiming to maximize CLIP’s potential without additional fine-tuning.
+The most straightforward baseline in this category computes the normalized text-image similarity as the OOD score.
+To improve upon this, MCM (Ming et al., 2022b) replaces the similarity score with a Maximum Concept Matching (MCM) score, providing a theoretically grounded and empirically robust alternative.
+The MCM score is defined as:
+(6) $S_{MCM}(x^{\prime};Y_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}}.$
+where $x’$ ’ denotes the test image, $Y_{in}$ is the set of in-distribution labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
+Recent works have further enriched the textual context used by CLIP to improve OOD detection.
+For example, NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024) all leverage curated corpora to augment the set of candidate labels.
+Specifically, NegLabel selects negative labels from a corpus that are semantically dissimilar to ID classes, thereby sharpening the model’s ability to distinguish OOD samples.
+CLIPScope (Fu et al., 2024) goes a step further by integrating these negative labels into a Bayesian scoring framework to revise the original OOD score.
+LAPT, in contrast, utilizes text-to-image generation or image retrieval models to obtain visual exemplars for both ID and negative labels, followed by prompt tuning to optimize OOD discrimination.
+It is worth emphasizing that methods such as ZOC (Esmaeilpour et al., 2022), CLIPN (Wang et al., 2023c), NegLabel (Anonymous, 2024), and CLIPScope (Fu et al., 2024) each propose distinct OOD scoring mechanisms tailored to their respective architectures.
+In contrast, the MCM score stands out for its generalizability and can be seamlessly integrated into diverse CLIP-based frameworks.
+Recent advances in zero-shot OOD detection have moved beyond static negative sampling by leveraging more adaptive strategies within the joint vision-language space.
+For example, AdaNeg (Zhang and Zhang, 2024) dynamically generates negative proxies at test time to better match the semantics of OOD samples, while AdaND (Cao et al., 2025) focuses on mitigating the impact of noisy samples by decoupling noise detection from classification using CLIP-based pseudo-labels.
+In contrast, SimLabel (Zou et al., 2025) distinguishes OOD inputs by analyzing the consistency of image-label similarities among semantically related classes, automatically constructing similar class groups via hierarchical structures and large language models.
+These approaches thus differ in their strategies for negative sampling, noise handling, and semantic grouping, reflecting the diverse directions in recent OOD detection research.
+Complementing these proxy- and consistency-based advances, several works have proposed to enhance OOD detection by expanding the diversity and expressiveness of negative or outlier labels.
+OLE (Ding and Pang, 2024) exposes CLIP to a large set of outlier labels, selected from external corpora and clustered into prototypes, to better model the outlier space.
+Furthermore, OLE augments these prototypes through interpolation with ID representations, effectively filling the semantic gap between ID and OOD regions in the embedding space.
+Similarly, recent work on prompt learning for near-OOD detection (Jung et al., 2024) introduces novel scoring functions, such as the Empty-Class Score, which measures the alignment between an image and the context vector excluding any class name, and relative scores that subtract this alignment from standard logit-based OOD scores.
+This simple post-hoc adjustment provides substantial gains, especially for challenging near-OOD cases, without requiring retraining or architectural changes.
+Beyond label- and score-centric approaches, matrix approximation techniques have also been explored.
+SeTAR (Li et al., 2024b) applies selective low-rank approximation to CLIP’s weight matrices, aiming to retain the most informative components for OOD discrimination while filtering out spurious correlations.
+In a different vein, OT-DETECTOR (Liu et al., 2025a) formulates OOD detection as an optimal transport problem, quantifying the semantic and distributional discrepancy between test samples and ID labels, and further refining detection by semantic-aware content refinement.
+Collectively, these developments illustrate a clear methodological evolution: from static post-hoc scoring and negative label augmentation, to adaptive proxy generation, label consistency modeling, prototype-based outlier exposure, and low-rank or optimal transport-based representations.
+This progression not only improves OOD detection accuracy but also enhances robustness in multi-label, semantically rich, and highly dynamic deployment scenarios.
+Furthermore, the interplay between proxy-based, consistency-based, and low-rank approximation methods suggests promising directions for unified frameworks that leverage the strengths of each paradigm.
 LLM-based.
-With the flourishing development of LLMs, new opportunities have emerged in the field of OOD detection with the application of LLMs.
-The strength of LLMs resides in their extensive world knowledge, enabling them to furnish comprehensive details about ID labels.
-Nevertheless, the issue of generating false or misleading information, known as hallucinations, poses a significant challenge when employing LLMs for OOD detection tasks.
-Existing methods (Huang et al., 2024; Dai et al., 2023) are based on an observation: two categories of images may have very similar visual features but different properties in the semantic space.
+With the rapid advancement of LLMs, new opportunities have emerged in the field of OOD detection by leveraging the extensive world knowledge and powerful semantic reasoning capabilities of LLMs.
+The primary strength of LLMs lies in their ability to provide comprehensive and context-rich descriptions of ID labels, which can be crucial for distinguishing between ID and OOD samples.
+However, a significant challenge in employing LLMs for OOD detection is their tendency to generate hallucinations—false or misleading information—which can undermine the reliability of OOD identification.
+Recent works have sought to address these challenges by exploring the nuanced relationship between visual and semantic spaces.
+For instance, existing methods (Huang et al., 2024; Dai et al., 2023) observe that visually similar image categories may possess distinct properties in the semantic space, which can be exploited for more effective OOD detection.
 Dai et al.
-(2023) introduce a consistency-based calibration method with an object detector to mitigate hallucinations and utilize LLMs’ world knowledge for describing ID classes.
-However,  Huang et al.
-(2024) propose ODPC to generate peer-class based on ID labels and OOD samples within text and image modality.
-It should be noted that ODPC requires ID images for training MLP, indicating that it is not a zero-shot method.
-Regarding the OOD score,  Dai et al.
-(2023) adopt MSP, however, ODPC uses a KNN-based score.
-Later, a study  (Liu et al., 2024b) on LLMs was proposed to study three problems with LLMs doing OOD detection: the propensity of LLMs for near-OOD and far-OOD, the effect of different fine-tuning methods, and the choice of OOD score.
-Liu et al.
-(2024b) find that LLMs are natural far-OOD detectors, generative fine-tuning is better than discriminative fine-tuning, and cosine distance is sufficient as OOD score because the embedding space of LLMs exhibits anisotropy.
+(2023) introduce a consistency-based calibration method that incorporates an object detector to mitigate hallucinations, thereby leveraging the world knowledge of LLMs for more accurate ID class descriptions.
+In contrast, Huang et al.
+(2024) propose ODPC, which generates peer classes based on both ID labels and OOD samples across textual and visual modalities.
+Notably, ODPC requires ID images for training an MLP, indicating that it is not a zero-shot approach.
+In terms of OOD scoring, Dai et al.
+(2023) adopt the MSP, while ODPC utilizes a KNN-based scoring strategy.
+Building upon these foundations, subsequent studies have systematically investigated the behavior of LLMs in OOD detection.
+For example, Liu et al.
+(2024a) analyze the propensity of LLMs to detect near-OOD versus far-OOD samples, the impact of different fine-tuning strategies, and the suitability of various OOD scoring functions.
+Their findings reveal that LLMs are naturally adept at detecting far-OOD samples, that generative fine-tuning outperforms discriminative fine-tuning, and that cosine distance is an effective OOD score due to the anisotropic nature of LLM embedding spaces.
+More recently, LLM-based approaches such as CMA (Lee et al., 2025), COOD (Liu et al., 2024c), and EOE (Cao et al., 2024) have extended OOD detection into multi-label and concept-rich regimes.
+These methods utilize LLMs not only as knowledge generators, but also as proxy creators and synthetic data generators (Abbas et al., 2025), thereby enriching the semantic context available for OOD detection.
+The synergy between VLMs for visual grounding and LLMs for semantic reasoning has been further highlighted by works such as ReGuide (Kim et al., 2025), which propose self-guided, image-adaptive concept generation to enhance robustness.
+Collectively, these advancements demonstrate the growing potential of LLM-based techniques to address the complex challenges of OOD detection by effectively integrating visual and semantic information.
 Remark.
@@
 However,  Fort et al.
-(2021) argue that in deployment scenarios, labels for OOD are either readily available.
+(2021); Ding and Pang (2024) argue that in deployment, labels for OOD are either readily available.
 They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance.
@@
 Table 3.
-Performance evaluation of some competitive methods using the ImageNet-1K dataset as the ID dataset and iNaturalist, SUN, Places, and Textures as OOD datasets.
-The best result is emphasized in bold.
+Performance evaluation of some competitive methods based on CLIP-ViT-B/16 using the ImageNet-1K dataset as the ID dataset and iNaturalist, SUN, Places, and Textures as OOD datasets.
+The results are cited from (Ming et al., 2022b; Chen et al., 2024b; Zhang and Zhang, 2024; Li et al., 2024c) and the best result is emphasized in bold.
 Scenario Method iNaturalist SUN Places Textures Average
@@
 NegLabel  (Anonymous, 2024) 99.49 1.91 95.49 20.53 91.64 35.59 90.22 43.56 94.21 25.40
+LAPT  (Zhang et al., 2024a) 99.63 1.16 96.01 19.12 92.01 33.01 91.06 40.32 94.68 23.40
+CSP  (Chen et al., 2024b) 99.60 1.54 96.66 13.66 92.90 29.32 93.86 25.52 95.76 17.51
+AdaNeg  (Zhang and Zhang, 2024) 99.71 0.59 97.44 9.50 94.55 34.34 94.93 31.27 96.66 18.92
 Few-shot CoOp  (Zhou et al., 2022a) 93.77 29.81 93.29 40.83 90.58 40.11 89.47 45.00 91.78 51.68
 LoCoOp  (Miyai et al., 2023a) 96.86 16.05 95.07 23.44 91.98 32.87 90.19 42.28 93.52 28.66
-NegPrompt  (Li et al., 2024b) 98.73 6.32 95.55 22.89 93.34 27.60 91.60 35.21 94.81 23.01
+NegPrompt  (Li et al., 2024c) 98.73 6.32 95.55 22.89 93.34 27.60 91.60 35.21 94.81 23.01
 5.2. Few-shot OOD Detection Approaches
@@
 Similarly,  Dong et al.
-(2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  ming2023does.
+(2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  Ming and Li (2024).
 Following this,  Kim et al.
@@
 (2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
-Fine-tuning based.
-A prevalent method for fine-tuning CLIP using limited ID data involves prompt learning (Zhou et al., 2022b, 2023), where the fine-tuning focuses on the prompt’s contextual words, maintaining the pre-trained parameters unchanged.
-To distinguish between ID and OOD data, both LoCoOp  (Miyai et al., 2023a) and ID-like Prompting Learning  (Bai et al., 2024) are designed to enhance the learning of context vectors from the near-ID feature perspective.
-LoCoOp employs entropy maximization to distance ID-irrelevant local features (such as the backgrounds in ID images) from the textual embeddings of ID classes.
-Similarly, ID-like Prompting Learning generates ID-like data—outliers within the vicinity of ID samples—to refine the context vector.
-Additionally, it incorporates a diversity loss to enhance the variance among the sampled OOD data.
-However, recent work EOK-Prompt (Zeng et al., 2024) and GalLop (Lafon et al., 2024) argues that optimizing only a prompt using numerous features from an image wastes valuable information.
-Instead, EOK proposes an orthogonal method to optimize the novel local prompt, making better use of the limited image data, while GalLop optimizes global prompts and local prompts by leveraging global and local features, respectively.
-Furthermore, based on the mentioned studies, DSFG  (Dong et al., 2023) posits that traditional fine-tuning approaches applied to CLIP may inadvertently lead to a loss of critical OOD knowledge.
-To address this issue, DSFG adopts a strategy to retain broad OOD knowledge by merging the original features with those modified through fine-tuning, followed by the training of a classifier.
-DSFG’s plug-and-play capability makes it seamlessly compatible with all fine-tuning-based approaches, enhancing its practicality and value.
-The aforementioned few-shot tuning methods inevitably suffer from overfitting on limited-shot instances.
-Chen et al.
-(2024b) proposed a training-free few-shot OOD detection method, Dual-Adapter(Chen et al., 2024b), which constructs two types of adapters by extracting positive and negative features to aid in OOD detection.
-The currently mentioned few-shot methods lack OOD supervision, so CLIP-OS (Sun et al., 2024) tries to find such supervision and also achieves stunning results
+Few-shot OOD Detection.
+Few-shot OOD detection aims to leverage a limited amount of ID data to adapt LPMs, primarily through prompt learning or parameter-efficient fine-tuning.
+Early approaches such as CoOp (Zhou et al., 2022b) and ZegCLIP (Zhou et al., 2023) focus on optimizing the contextual words in prompts while keeping the backbone fixed, thus preserving the generalization ability of the pre-trained models.
+Building upon this, LoCoOp (Miyai et al., 2023a) and IDPL (Bai et al., 2024a) further enhance context vector learning from a near-ID perspective: LoCoOp maximizes entropy to separate ID-irrelevant local features (e.g., backgrounds) from textual embeddings of ID classes, while IDPL synthesizes ID-like outliers near the ID boundary and employs a diversity loss to encourage variance among sampled OOD candidates.
+In addition, MMFT (Kim and Hwang, 2025) projects image and text features into a shared hyperspherical space and introduces a cross-modal alignment loss function to promote the alignment of image and text representations in the hyperspherical space.
+During training, the loss function brings the image and text embeddings closer together while effectively distinguishing OOD texts from ID texts.
+However, these prompt-only methods may underutilize the rich information contained in image features.
+Recent advances such as GalLoP (Lafon et al., 2024) and Local-Prompt (Zeng et al., 2025) address this limitation by jointly optimizing global and local prompts.
+GalLoP leverages both global and region-level features via prompt dropout and region-level regularization to improve generalization and robustness, whereas Local-Prompt introduces orthogonal local prompt optimization to better exploit limited image data.
+Beyond prompt learning, DSFG (Dong et al., 2023) highlights that conventional fine-tuning can erode critical OOD knowledge.
+To mitigate this, DSFG merges original and fine-tuned features before classifier training, thereby retaining broad OOD awareness and offering plug-and-play compatibility with other fine-tuning-based methods.
+Recently, SCT (Yu et al., 2024a) proposes a self-calibrated tuning framework that dynamically balances ID classification and OOD regularization by introducing modulation factors based on sample uncertainty, enabling more reliable adaptation with only a few ID samples.
+Complementary to this, GaCoOp (Tong et al., 2025) addresses the gradient conflict between ID classification and OOD regularization in prompt optimization.
+By decomposing and projecting gradients, GaCoOp aligns the optimization directions, thereby mitigating the adverse effects of background bias and improving few-shot OOD detection performance, especially when data is scarce.
+Despite these advancements, overfitting remains a challenge due to the scarcity of labeled data.
+To tackle this, Dual-Adapter (Chen et al., 2024c) proposes a training-free few-shot method that constructs positive and negative adapters in a plug-and-play manner, requiring only a handful of ID samples.
+Meanwhile, to address the intrinsic modality gap in VLMs, SUPREME (Wang et al., 2025a) introduces multi-modal prototypes and alignment modules, while CLIP-OS (Sun et al., 2024) regularizes the ID/OOD boundary by synthesizing OOD data through mixing ID features, achieving strong performance without external OOD supervision.
+Additionally, NegPrompt (Li et al., 2024c) further refines the boundary by optimizing negative prompts, demonstrating the continued evolution and diversification of few-shot OOD detection strategies.
 Meta-learning based.
@@
 Overview.
-This setup is generally less realistic than the first two (zero-shot and few-shot).
-However, we list them separately to ensure a comprehensive review of existing methods across the spectrum.
+While this setup is generally less realistic than the first two (zero-shot and few-shot), we list them separately to ensure a comprehensive review of existing methods across the spectrum.
 Given the full set of ID data and corresponding labels, VLMs can enhance OOD detection significantly by fine-tuning.
@@
 6. Evaluation and Application
+Table 4.
+Summary of Datasets.
+The CARLA System is a simulation platform designed for evaluating OOD Detection in the field of autonomous driving, hence its entire row is filled with dashes.
+“*”indicates that this dataset is divided by category, with some data used as ID and other data as OOD.
+“/” symbol in the Usage column indicates that this dataset can be used as both ID and OOD data.
+TASK Dataset Name Data Type # Classes # Samples Reference Usage
+Image Classification CIFAR-10 Images 10 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022) ID/OOD
+CIFAR-100 Images 100 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022) ID/OOD
+MNIST Images 10 70,000 (Hendrycks and Gimpel, 2017; Zhao et al., 2024) ID/OOD
+ImageNet-1K Images 1,000 1,431,167 (Ming et al., 2022b; Wang et al., 2023c) ID
+iNaturalist Images 5,089 675,170 (Ming et al., 2022b; Wang et al., 2023c) OOD
+SUN Images 397 108,754 (Ming et al., 2022b; Wang et al., 2023c) OOD
+Places Images >205 >2,500,000 (Ming et al., 2022b; Wang et al., 2023c) OOD
+Textures Images 47 5,640 (Ming et al., 2022b; Wang et al., 2023c) OOD
+Semantic Segmentation Cityscapes Images 30 25,000 (Gao et al., 2023a) ID
+Road Anomaly Dataset Images > 5 60 (Gao et al., 2023a) OOD
+Object Detection PASCAL VOC Images 20 2,913 (Nguyen et al., 2024; Du et al., 2022b) ID
+BDD100K Images 10 100,000 (Nguyen et al., 2024; Du et al., 2022b) ID
+MS-COCO Images 80 > 930 (Nguyen et al., 2024; Du et al., 2022b) OOD
+OpenImages Images 601 1761 (Nguyen et al., 2024; Du et al., 2022b) OOD
+Autonomous Driving CARLA System - - - (Mao et al., 2024) -
+Medical Image Analysis Kvasir-Capsule* Images 14 4,741,621 (Tan et al., 2024) ID/OOD
+Text Category News Category* Text 41 210,000 (Arora et al., 2021) ID/OOD
+SST-2 Text 2 215,154 (Arora et al., 2021) ID/OOD
+Intent Detection CLINC150* Text 150 22,500 (Zhan et al., 2022) ID/OOD
+Banking77* Text 77 13,083 (Zhan et al., 2022) ID/OOD
+StackOverflow* Text 20 20,000 (Zhan et al., 2022) ID/OOD
+Audio MSCW Audio 31 >23,400,000 (Bukhsh and Saeed, 2023) ID/OOD
+Vocalsound Audio 6 21,024 (Bukhsh and Saeed, 2023) ID/OOD
+Graph data TU/OGB* Graph data 10 19,766 (Wang et al., 2024a) ID/OOD
 6.1. Evaluation metrics
@@
 This categorization corresponds to dividing OOD detection tasks into near-OOD and far-OOD detection.
-It is evident the near-OOD detection task is more challenging, however, numerous methods  (Fort et al., 2021; Ming et al., 2022b; Bai et al., 2024) have demonstrated excellent performance in this area.
+It is evident the near-OOD detection task is more challenging, however, numerous methods  (Fort et al., 2021; Ming et al., 2022b; Bai et al., 2024a) have demonstrated excellent performance in this area.
 Recently,  Yang et al.
 (2023b); Bai et al.
-(2023) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously.
+(2023); Yu et al.
+(2024b) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously.
 This is crucial to prevent the loss of model generalization.
@@
 6.3. Application
-Table 3.
-Summary of Datasets.
-The CARLA System is a simulation platform designed for evaluating OOD Detection in the field of autonomous driving, hence its entire row is filled with dashes.
-Other “-” symbols represent numbers that vary depending on usage.
-More detailed descriptions can be found in the code repo.
-TASK Dataset Name Data Type # Classes # Samples Papers
-Image Classification CIFAR-10 Images 10 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
-CIFAR-100 Images 100 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
-MNIST Images 10 70,000 (Hendrycks and Gimpel, 2017; Zhao et al., 2024)
-ImageNet-1K Images 1,000 1,431,167 (Ming et al., 2022b; Wang et al., 2023c)
-iNaturalist Images 5,089 675,170 (Ming et al., 2022b; Wang et al., 2023c)
-SUN Images 397 108,754 (Ming et al., 2022b; Wang et al., 2023c)
-Places Images >205 >2,500,000 (Ming et al., 2022b; Wang et al., 2023c)
-Textures Images 47 5,640 (Ming et al., 2022b; Wang et al., 2023c)
-Semantic Segmentation Cityscapes Images - 25,000 (Gao et al., 2023b)
-Road Anomaly Dataset Images - 100 (Gao et al., 2023b)
-Object Detection PASCAL VOC Images 20 2,913 (Nguyen et al., 2024; Du et al., 2022b)
-BBD100K Video->Image Variable 100,000 (Nguyen et al., 2024; Du et al., 2022b)
-Autonomous Driving CARLA System - - - (Mao et al., 2024)
-Medical Image Analysis Kvasir-Capsul Images 4,741,621 (Tan et al., 2024)
-Text Category News Category Text - 210,000 (Arora et al., 2021)
-SST-2 Text - 215,154 (Arora et al., 2021)
-Intent Detection CLINC150 Text 150 22,500 (Zhan et al., 2022)
-Banking Text 77 13,083 (Zhan et al., 2022)
-StackOverflow Text 20 20,000 (Zhan et al., 2022)
-Audio MSCW Audio - >23400000 (Bukhsh and Saeed, 2023)
-Vocalsound Audio - 21,024 (Bukhsh and Saeed, 2023)
-Graph data TU Graph data - Variable (Wang et al., 2024)
-OGB Graph data - Variable (Wang et al., 2024)
+This section presents a comprehensive review of both academic and real-world applications of OOD detection.
+To promote further research and enable standardized evaluation, commonly used benchmark datasets are systematically summarized in Table 4.
+To provide a comprehensive overview of recent advances in OOD detection, we first organize our discussion based on fundamental technical domains (Computer Vision, Natural Language Processing, and Beyond) as these domains have distinct methodological characteristics and theoretical foundations that shape how OOD detection is approached.
+This domain-based categorization allows us to systematically analyze the core technical innovations and challenges specific to each field.
+We then extend our analysis to an application-oriented perspective, as shown in Table 6.3, which bridges the gap between theoretical advances and practical implementations, illustrating how these domain-specific techniques are adapted and deployed in real-world scenarios.
+Table 5. Representative applications of OOD detection across five major domains, highlighting key use cases in industrial, medical, safety, scientific, and general-purpose settings.
+Domain Task Application Description Reference
+Industrial Applications Autonomous Driving Detecting unseen obstacles or traffic scenarios to ensure driving safety in autonomous vehicles. (tes, )
+3D Object Detection Identifying out-of-distribution objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
+Object Detection Finding anomalous or unknown items in industrial production lines. (Syed et al., 2024; Chen et al., 2024d; Zhang et al., 2025b)
+Earth Observation Imagery Detecting novel land-cover types or changes in satellite and remote sensing images. (Le Bellier and Audebert, 2024)
+Edge Computing Real-time detection of abnormal data on edge devices to enhance responsiveness and security. (Li et al., 2025)
+Time Series Analysis Monitoring equipment by identifying abnormal patterns for predictive maintenance. (web, )
+Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Oh et al., 2024; Pokhrel et al., 2024; Smedsrud et al., 2021)
+Biological Classification Identifying previously unseen species and enhancing biodiversity monitoring. (Arslan et al., 2024)
+Security and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
+Facial Expression Recognition Identifying unknown or abnormal facial expressions in surveillance. (Zhang et al., 2024b)
+Human Action Recognition Detecting anomalous or dangerous actions in monitored environments. (Sim et al., 2023; Mandal et al., 2019; Xu et al., 2024b)
+Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021; Gong et al., 2022)
+Scientific Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
+Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Giger and Csillaghy, 2024; Das et al., 2024)
+Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b; Jiang et al., 2025)
+Mathematical Reasoning Identifying out-of-distribution problem types in automated reasoning systems. (Wang et al., 2024b)
+Spiking Neural Networks Recognizing abnormal neural signals in bio-inspired computing models. (Terres-Escudero et al., ; Orchard et al., 2015)
+General-Purpose Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
+Semantic Segmentation Identifying unknown regions or objects in pixel-level segmentation tasks. (Gao et al., 2023a; Lis et al., 2019)
+Text Classification Detecting unseen topics or intents in text data. (Zhan et al., 2024)
+Intent Detection Recognizing new user intents in intelligent assistants and chatbots. (Wang et al., 2024c; Zhan et al., 2022)
+Question-Answering Systems Detecting questions that are out-of-scope or unanswerable by the system. (Zhang et al., 2024c)
+Document Classification Detecting anomalies in documents containing multiple data modalities. (Constantinou et al., 2024)
+Reinforcement Learning Detecting abnormal behavior or novel states in agent-environment interactions. (Chen et al., 2025b; Sedlmeier et al., 2019)
 6.3.1. Computer Vision
@@
   • Image Classification.
-The majority of tasks discussed in this paper focus on OOD detection within the realm of image classification.
-In such task scenarios, commonly utilized ID datasets include MNIST  (Xiao et al., 2017), CIFAR-10  (Krizhevsky et al., 2009), CIFAR-100  (Krizhevsky et al., 2009), and ImageNet-1K  (Deng et al., 2009).
-Various OOD datasets are constructed accordingly for evaluating different methods  (Yang et al., 2022b), with the most frequently employed datasets being iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017) and Textures (Cimpoi et al., 2014).
-Additionally, full-spectrum OOD is typically evaluated on three benchmarks: DIGITS, OBJECTS, and COVID, as proposed by  Yang et al.
-(2023b).
+The majority of OOD detection research targets image classification, utilizing ID datasets such as MNIST (Xiao et al., 2017), CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009).
+Common OOD datasets include iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017), and Textures (Cimpoi et al., 2014).
+Full-spectrum OOD evaluation often employs the DIGITS, OBJECTS, and COVID benchmarks (Yang et al., 2023b).
+In addition to the aforementioned methods, most current research classifies OOD datasets based on the difficulty of distinguishing them from ID datasets.
+The most popular approach is to use ImageNet-1k as the ID dataset, with SSB-hard (Vaze et al., 2022) and NINCO (Bitterwolf et al., 2023) categorized as near-OOD (representing categories that are difficult to distinguish), and iNaturalist (Van Horn et al., 2018), Textures (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022c) categorized as far-OOD (representing categories that are easier to distinguish
   • Semantic Segmentation.
-Recent works  (Gao et al., 2023b) have started delving into the dense OOD detection task, also known as anomaly segmentation.
+Recent works  (Gao et al., 2023a) have started delving into the dense OOD detection task, also known as anomaly segmentation.
 The datasets used for evaluation include the Cityscapes dataset  (Cordts et al., 2016), the Road Anomaly dataset  (Lis et al., 2019), and the recently developed SOOD-ImageNet  (Bacchin et al., 2024).
   • Object Detection.
-The application of methods related to OOD detection in the field of object detection is relatively nascent, with only a few studies exploring this area  (Du et al., 2022b).
-Evaluations are commonly performed using datasets like PASCAL-VOC  (Everingham et al., 2015) and Berkeley DeepDrive-100K  (Yu et al., 2018).
+OOD detection in object detection is relatively new (Du et al., 2022b), with evaluations on datasets like PASCAL-VOC (Everingham et al., 2015) and Berkeley DeepDrive-100K (Yu et al., 2018).
+Novel approaches include Situation Monitor, which uses diversity-driven budding ensemble architectures for zero-shot OOD detection in Transformer-based models (Syed et al., 2024), Proto-OOD, which enhances OOD object detection with prototype feature similarity (Chen et al., 2024d), and VisTa, which employs visual-contextual and text-augmented methods for zero-shot object-level OOD detection (Zhang et al., 2025b).
+  • 3D Object Detection.
+OOD detection in LiDAR-based 3D object detection generates synthetic OOD data and trains MLPs to distinguish ID and OOD targets, with new evaluation protocols for realistic scenarios (Kösel et al., 2024).
   • Autonomous Driving.
-Autonomous driving has long been a crucial practical application of OOD Detection.
-Recently, Mao et al.
-(2024) utilized the CARLA (Dosovitskiy et al., 2017) system to simulate and evaluate the performance of OOD Detection in autonomous driving scenarios.
+Autonomous driving is a critical application of OOD detection.
+Recent work uses the CARLA simulator (Dosovitskiy et al., 2017) to evaluate OOD detection performance in driving scenarios (Mao et al., 2024).
+Mao et al.
+(2024) propose a language-augmented latent representation method, leveraging the image-text cosine similarity from the CLIP model to improve transparency and controllability in detection.
+Another recent work focus on real-time OOD perception in trajectory prediction, further enhancing vehicle safety (Banerjee et al., 2024).
+In real-world applications, Tesla’s “Shadow Mode” (tes, ) identifies cases where AI predictions differ from human drivers in real traffic—these OOD samples are collected and used to retrain and improve autonomous driving models.
   • Medical Image Analysis.
-In the field of medical image analysis, OOD detection is crucial.
-Depending on the specific category of medical image, OOD detection employs various datasets including CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021).
-OOD detection has significant applications across various fields such as human action recognition (Sim et al., 2023) and solar image analysis (Giger and Csillaghy, 2024; Das et al., 2024).
-For further details, please refer to the accompanying code repository due to space limitations.
-The same applies to the subsequent paragraphs.
+OOD detection is vital in medical imaging, using datasets like CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021) depending on the image category (Tan et al., 2024).
+Recent studies focus on reliable OOD detection methods for digital pathology (Oh et al., 2024) and test-time augmentation for improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024).
+In practice, IDx-DR (Dis, ), the first FDA-approved autonomous AI diagnostic system, implements OOD detection by rejecting images that fail to meet quality standards or contain anomalies, deferring these cases to human experts for assessment.
+  • Human Action Recognition.
+OOD detection plays a crucial role in ensuring the robustness of Human Action Recognition (HAR) models, particularly when faced with previously unseen actions.
+By leveraging advanced techniques such as attention-based debiasing (Sim et al., 2023), generative feature synthesis (Mandal et al., 2019) and energy-based skeleton modeling (Xu et al., 2024b), recent research has significantly improved OOD detection performance.
+These methods are evaluated on benchmark datasets like UCF-101 (Soomro et al., 2012), HMDB-51 (Kuehne et al., 2011) and Kinetics-400 (Kay et al., 2017).
+  • Solar Image Analysis.
+In space weather forecasting, OOD detection serves as a crucial tool for identifying solar anomalies.
+Recent unsupervised  (Giger and Csillaghy, 2024; Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.
+  • Facial Expression Recognition.
+In facial expression analysis, the challenge of distinguishing unknown expressions from known categories has led to advanced OOD detection methods.
+Recent approach (Zhang et al., 2024b) focuses on leveraging the unique characteristics of facial expression features, incorporating attention map consistency and cycle training mechanisms to effectively identify OOD samples.
+  • Facial Authentication.
+In facial authentication systems, OOD detection is increasingly important for ensuring robust and secure identity verification.
+The FOOD framework (Kahya et al., 2024) exemplifies this trend by employing 60 GHz FMCW radar and a convolutional encoder-decoder architecture to achieve both accurate classification and reliable anomaly detection.
+  • Earth Observation Imagery.
+For earth observation imagery, OOD detection enhances the identification of rare or unexpected events, such as natural disasters.
+Diffusion models like ODEED (Le Bellier and Audebert, 2024) have demonstrated strong performance in detecting OOD samples, with evaluations conducted on benchmark datasets such as SpaceNet 8 to validate their effectiveness.
 6.3.2. Natural Language Processing
-OOD detection is also explored in various tasks across numerous Natural Language Processing applications. The two most common applications are as follows:
+OOD detection has become a crucial research topic in NLP, especially with the advent of LLMs and their deployment in real-world applications (Liu et al., 2025c; Lang et al., 2023).
+Below, we summarize key tasks and representative methods:
   • Intent Detection.
-Intent Detection is a significant application of OOD detection in NLP.
-The datasets used for evaluation include CLINC150  (Larson et al., 2019), Banking  (Casanueva et al., 2020), StackOverflow  (Xu et al., 2015), among others.
-  • Text Catagory.
-In text category OOD detection applications, datasets like News Category  (Misra, 2022) and SST-2 (Socher et al., 2013) are commonly utilized to form ID/OOD pairs and assess the models’ detection capabilities.
+Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022; Akbari et al., 2025).
+Common datasets for evaluation include CLINC150 (Larson et al., 2019), Banking77 (Casanueva et al., 2020), and HWU64 (Liu et al., 2020b).
+These datasets are specifically designed to evaluate OOD detection in intent classification tasks.
+  • Text Classification.
+In text classification tasks, OOD detection helps identify texts that do not belong to any of the known categories.
+Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for in-distribution data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset.
+Methods like VI-OOD (Zhan et al., 2024) enhance detection by optimizing joint distributions.
+  • Question-Answering Systems.
+Question-answering systems require robust mechanisms to handle diverse user queries while maintaining response accuracy.
+The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches.
+Using standard in-distribution datasets (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), PubMedQA (Jin et al., 2019)) and OOD datasets (SQUAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2019)), recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
+  • Mathematical Reasoning.
+Mathematical problem-solving systems face challenges when encountering problem types not seen during training.
+Using MultiArith (Roy et al., 2016) as the primary in-distribution dataset, researchers evaluate OOD detection against diverse benchmarks including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), SVAMP (Patil et al., 2021), AddSub (Hill et al., 2015), SingleEq (Wang et al., 2024b), and SingleOp (Wang et al., 2024b).
+Recent advances, such as TV-Score (Wang et al., 2024b), analyze embedding trajectory variations to enhance OOD detection performance.
+  • Document Classification.
+In (multimodal) document classification, where documents integrate both text and images, OOD detection is essential for identifying unseen document types.
+Datasets such as Tobacco3482 (Aggarwal et al., 2017) and FinanceDocs (Constantinou et al., 2024) are commonly used for evaluation.
+The AHM (Constantinou et al., 2024) approach enhances OOD detection by improving the representation of multimodal features.
 6.3.3. Beyond Computer Vision and Natural Language Processing
 In addition to the two data modalities mentioned above, OOD detection still has many important applications across various types of data.
-  • Audio data. In audio OOD detection, MSCW (Micro-EN) (Mazumder et al., 2021) and Vocalsound  (Gong et al., 2022) are usually used as ID datasets, and they also act as OOD for each other.
-  • Graph data.
-Recent studies have proposed various approaches for OOD detection in graph data.
-The existing graph-level OOD detection benchmark comprises a total of 10 pairs of datasets from TU  (Morris et al., 2020) and OGB datasets (Hu et al., 2020), which has been widely utilized.
+  • Audio Recognition.
+In the field of audio recognition, OOD detection is essential for identifying inputs such as synthesized speech and dialects that diverge from the training distribution.
+Traditional probabilistic scoring methods often underperform on atypical audio samples (Lane et al., 2006), motivating the adoption of alternative approaches including confidence-based topic classification (Lane et al., 2006), neural sentence embeddings (Ryu et al., 2017), and deep nearest neighbor analysis (Bukhsh and Saeed, 2023).
+Autoencoder architectures built upon WavLM (Du et al., 2024b) have demonstrated improved performance, particularly in detecting synthetic speech.
+Standard datasets for evaluating OOD detection in this domain include MSCW (Mazumder et al., 2021) and Vocalsound (Gong et al., 2022), which provide both in-distribution and out-of-distribution samples for comprehensive benchmarking.
+  • Graph Neural Networks.
+Within graph neural networks (GNNs), OOD detection encompasses both node-level and graph-level tasks.
+Commonly explored techniques include uncertainty-based graph classification (Bazhenov et al., 2022), energy-based detection (Wu et al., 2023b; Jiang et al., 2025), and semi-supervised learning (Huang et al., 2022b).
+The HGOE framework (Junwei et al., 2024) enhances graph-level detection by incorporating anomalous samples during training, while PGR-MOOD (Shen et al., 2024) utilizes diffusion models to identify out-of-distribution molecular structures.
+GOLD (Wang et al., 2025b) generates pseudo-OOD data with significant differences from the ID embeddings by using a latent generative model to mimic the embeddings of training data and training it jointly with a GNN encoder and an OOD detector.Benchmarking typically relies on paired datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020), which facilitate systematic evaluation of OOD detection methods in GNNs.
   • Reinforcement Learning.
-Currently, there is a rising trend of integrating OOD detection with reinforcement learning  (Leonardos et al., 2023) to bolster model robustness.
-Mohammed and Valdenegro-Toro (2021) provide a benchmark and explore methods for crafting tailored reinforcement learning environments that can generate OOD data.
+OOD detection plays a crucial role in reinforcement learning (RL) by enhancing system robustness when encountering unprecedented states or actions (Leonardos et al., 2023).
+Common approaches include uncertainty quantification through dropout techniques and bootstrapping methods (Sedlmeier et al., 2019), as well as systematic environment modifications (Mohammed and Valdenegro-Toro, 2021).
+In the context of offline RL, advanced methods such as ADAC (Chen et al., 2025b) systematically assess OOD actions to effectively address distribution shifts.
+Comprehensive evaluation frameworks for OOD detection in RL primarily concentrate on various forms of environmental perturbations (Mohammed and Valdenegro-Toro, 2021), providing standardized testing scenarios for robustness assessment.
+  • Time Series Analysis.
+In the domain of time series analysis, the inherent temporal dependencies present significant challenges for conventional OOD detection approaches.
+SRS (Belkhouja et al., 2023) addresses this by leveraging time series decomposition in conjunction with deep generative models to estimate conditional likelihoods, while CODiT (Kaur et al., 2023) introduces temporal equivariance to better capture sequential patterns.
+Recent studies have further evaluated modality-agnostic algorithms on multivariate time series datasets, outlining persistent challenges and future research directions (Gungor et al., 2025).
+In practice, OOD detection has been widely adopted in fraud detection services, such as the AWS Fraud Detector (web, ), which utilizes machine learning to identify transaction patterns that deviate from historical norms, effectively flagging potentially fraudulent activities in real time.
+  • Biological Classification.
+In biological classification, OOD detection plays a vital role in identifying previously unseen species or anomalies that extend beyond the categories present in training data.
+The integration of DNA barcodes has been shown to enhance image-based classification models through re-ranking strategies, thereby improving robustness and detection accuracy (Arslan et al., 2024).
+  • Spiking Neural Networks.
+Spiking Neural Networks (SNNs), recognized for their event-driven efficiency, present distinctive challenges for OOD detection due to their sparse latent representations.
+Recent methods have proposed combining feedforward learning with distance-based scoring and gradient-free attribution to address these challenges (Terres-Escudero et al., ).
+The N-MNIST dataset (Orchard et al., 2015) is commonly used for evaluation, with OOD samples generated through diverse event patterns.
+  • Astronomical Imaging.
+In astronomical imaging, OOD detection is essential for distinguishing between simulated and real galaxy images, where conventional statistical techniques often prove inadequate.
+Approaches such as Bayesian model comparison and sparse variational autoencoders (Zhou et al., 2024) have been utilized to uncover subtle discrepancies.
+The SDSS dataset is widely adopted for benchmarking, with simulation-induced biases serving as OOD samples.
+  • Edge Computing.
+Within edge computing, efficient and privacy-preserving OOD detection is critical for devices with limited computational resources.
+SecDOOD (Li et al., 2025) addresses these constraints by employing a collaborative cloud-device architecture and HyperNetwork-based parameter generation, thereby eliminating the need for device-side backpropagation.
+CIFAR-10 and Tiny ImageNet are standard datasets for in-distribution and OOD evaluation in this context.
 7. Emerging Trends And Open Challenges
@@
 Meta-Learning Adaptation.
-Faced with the quick adaptation challenge in test-time OOD detection, mete-learning algorithms, which provide a learning-to-learn paradigm to efficiently adapt the model to new test data, may be the solution.
-Additionally, in addressing the dauntingly vast sample space of potential outliers inherent in training-driven OOD detection, improved sampling methods could offer a pathway to the efficient utilization of outliers  (Ming et al., 2022a).
-Theoretically-Driven Score Designing.
-In the traditional single-modal regime, numerous post-hoc scores have been carefully designed for OOD detection, effectively reducing training costs.
-However, as the field has progressed into the multi-modal domain, there is an increasing demand for theoretically-driven score designs.
-MCM  (Ming et al., 2022b) is a notable example, but it is merely an extension of Softmax and does not deeply explore the relationships between text and image.
-Therefore, more advanced scores are needed.
+Rapid adaptation to novel data distributions at test time remains a significant challenge in OOD detection.
+Meta-learning algorithms, which follow a learning-to-learn paradigm, have emerged as promising solutions for enabling efficient adaptation to unseen test data.
+Additionally, the vast sample space of potential outliers in training-driven OOD detection motivates the development of improved sampling strategies for more effective utilization of outlier information (Ming et al., 2022a).
+Model Selection.
+Selecting the most suitable OOD detection approach for a given task is non-trivial, as detection performance often varies across data distributions.
+Automated model selection frameworks, such as MetaOOD (Qin et al., 2025), leverage meta-learning to facilitate zero-shot model selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) utilize proportion estimation and likelihood ratio tests to enhance ensemble and score fusion strategies.
+Interestingly, recent work has shown that partially trained, immature models can exhibit effective OOD detection capabilities (Montazeran and Köthe, 2025), suggesting new directions for leveraging model dynamics in detection.
+Active Learning.
+Active learning strategies, which prioritize labeling the most informative data points, offer efficiency gains in data-scarce settings.
+In OOD detection, active learning can focus annotation efforts on ambiguous or high-uncertainty samples, thereby improving the model’s discrimination between in-distribution and out-of-distribution data.
+SISOM (Schmidt et al., 2024) exemplifies a unified approach that combines active learning and OOD detection by selecting samples based on feature space distances.
+Continual and Incremental Learning.
+As real-world data is often encountered in a streaming and evolving manner, continual and incremental learning paradigms are increasingly relevant for OOD detection.
+These approaches enable models to adapt to new tasks and data distributions over time while retaining knowledge of previous ones.
+Recent efforts include continual unsupervised OOD detection (Doorenbos et al., 2024), systematic evaluation benchmarks like OpenCIL (Miao et al., 2024b), and hierarchical two-sample tests for continual OOD detection (Liu et al., 2025d).
+Theoretically-Driven Score Design.
+While many post-hoc scoring methods have been proposed for OOD detection in single-modal regimes, the transition to multi-modal domains necessitates more principled, theoretically motivated score designs.
+For example, MCM (Ming et al., 2022b) extends softmax-based approaches, but further innovation is needed to better capture the complex relationships in multi-modal data.
 7.2. More practical scenarios of OOD detection
@@
 Quick Test-Time Adaptation.
-The advent of test-time scenarios like CAOOD  (Wu et al., 2023) significantly advances the application of OOD detection in real-world contexts, promising improved reliability and adaptability.
+The introduction of test-time adaptation scenarios, such as CAOOD (Wu et al., 2023a), marks a step forward in making OOD detection more reliable and adaptable in dynamic real-world environments.
+Federated Environments.
+In practical scenarios, data is often distributed across multiple clients with heterogeneous distributions.
+Federated learning (FL) enables collaborative model training while preserving privacy, but also brings unique OOD detection challenges due to covariate and semantic shifts among clients.
+Addressing these challenges is crucial for robust federated models.
+For instance, FOOGD (Liao et al., 2024) combines score matching and feature invariance to handle simultaneous OOD shifts in federated settings.
 Multi-Modal Detection.
-Exploring multi-modal OOD detection enhances our grasp of data dynamics and boosts model efficacy across varied sensory inputs.
-Additionally, the forthcoming integration of LLMs into multi-modal OOD detection stands to transform the field, marrying advanced linguistic analysis with other data types to forge more effective and versatile detection systems for complex applications.
+The proliferation of multi-modal data—ranging from images and text to audio and video—demands OOD detection methods capable of handling diverse inputs.
+Benchmarks such as MultiOOD (Dong et al., 2024) provide comprehensive testbeds for multi-modal OOD detection, introducing algorithms like A2D and NP-Mix for challenging scenarios.
+Further advancements, including DIAEF (Gao et al., 2024) and DPU (Li et al., 2024d), focus on cross-modal alignment and dynamic prototype updating, respectively.
+The integration of large language models (LLMs) into multi-modal OOD detection is poised to further enhance the field by combining advanced linguistic reasoning with other modalities.
+Multi-Label OOD Detection.
+Real-world applications often involve multi-label data, where each instance may belong to multiple categories.
+Detecting OOD samples in these scenarios is challenging due to increased dimensionality and label correlations.
+Recent approaches (Mei et al., 2024; Aguilar et al., 2025) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
 Open-Vocabulary Scenario.
-Existing LPM-based methods assume that ID categories are accessible.
-However, this assumption does not hold in open-vocabulary scenarios.
-Li et al.
-(2024b) addresses this by learning migratable negative prompts.
-It uses a subset of ID labels for prompt learning, and these learned prompts can be applied to other unseen labels.
-OOD detection in open-vocabulary settings further relaxes the requirement for prior access to ID knowledge, paving the way for the development of more generalized models.
-Noise Settings.
-While previous research has largely focused on standard clean label settings, few studies have explored the other side of the coin: noisy label scenarios.
-Humblot-Renaux et al.
-(2024) investigate practical scenarios and examine the impact of varying levels of noise on OOD detection.
-Then they provide several key takeaways for improving OOD detection in noisy label environments.
+Traditional approaches often assume access to all in-distribution categories, an assumption that does not hold in open-vocabulary contexts.
+Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of in-distribution labels is available, thus paving the way for more generalizable models.
+Noisy Label Environments.
+Most prior studies focus on clean label settings, but real-world data often contains label noise.
+Investigations into noisy label scenarios (Humblot-Renaux et al., 2024) offer insights and practical recommendations for improving OOD detection robustness under such conditions.
+Long-Sequence Processing.
+In NLP, the detection of OOD inputs becomes increasingly complex when language models process sequences that exceed their pre-trained context window.
+In such scenarios, models frequently encounter unfamiliar dependency structures, increased attention demands, and altered positional encoding behaviors, all of which deviate from patterns seen during training.
+These distributional shifts impede the model’s generalization capabilities, highlighting the importance of thoroughly understanding their effects.
+Advancing OOD detection in this context is essential for ensuring reliable model performance on sequences that extend beyond standard context boundaries (Liu et al., 2025c).
 7.3. New applications of OOD detection
 Additional Modalities.
-While OOD detection has made strides across diverse sectors, its potential in speech and physiological signal analysis remains largely untapped.
-In particular, OOD issues are prevalent in emotion-related physiological signals due to the variability in subjects’ emotional states.
-Therefore, this represents a promising area for further exploration.
-Some of the current methods  (Dong et al., 2024) also provide benchmarks for reference.
-Human-in-the-loop Application.
-A pivotal future direction, proposed by  Vishwakarma et al.
-(2024), is the integration of human insight into the detection process.
-This human-in-the-loop approach, particularly in high-stakes decision-making, will be vital for enhancing the accuracy and responsiveness of OOD detection systems, merging human intuition with algorithmic precision.
+While significant advances have been made in vision and language domains, OOD detection in modalities such as speech and physiological signals remains underexplored.
+These areas, particularly emotion-related physiological signals with high inter-subject variability, present promising directions for future research (Dong et al., 2024).
+Human-in-the-Loop Systems.
+The integration of human expertise into OOD detection is emerging as a critical direction, especially in high-stakes or safety-critical applications.
+Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) incorporate human feedback to refine detection thresholds and guide annotation, balancing algorithmic precision with human intuition.
+The AHA framework, for instance, strategically targets ambiguous regions for selective annotation, improving generalization and error correction under limited supervision.
 Web Image Scraping.
-To automate the process of scraping images from the Internet,  Miyai et al.
-(2023c) propose a zero-shot ID detection task, a concept derived from zero-shot OOD detection.
-An image will be classified as an ID image if it contains ID objects; it will only be considered an OOD image if it lacks any ID objects.
-It presents a novel perspective on the application of OOD detection and merits further exploration
+OOD detection is also finding novel applications in automating web image scraping.
+For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as in-distribution if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
 8. Conclusion
-OOD detection is critical for trustworthy machine learning.
+OOD detection is a critical component for trustworthy machine learning.
 In this paper, we provide a comprehensive review of recent advances in OOD detection, focusing for the first time on the problem scenario perspective: training-driven, training-agnostic, and large pre-trained model-based OOD detection.
 We also summarize extensively used evaluation metrics, experimental protocols, and diverse applications.
-We believe that our novel taxonomy of existing papers and extensive discussion of emerging trends will contribute to a better understanding of the current state of research, assist practitioners in selecting suitable approaches, and inspire new research hotspots.
+We believe that our taxonomy of existing papers and extensive discussion of emerging trends will contribute to a better understanding of the current state of research, assist practitioners in selecting suitable approaches, and inspire new research hotspots.
+9. Acknowledgements
+We sincerely appreciate the constructive comments and suggestions provided by all the reviewers and the editor, which have greatly improved the quality of our work.
+We are truly grateful for their valuable time and insightful feedback.

### v4

@@
 Salehi et al.
-(2022) offer a comprehensive review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.
+(2022) offer a review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.
 Yang et al.
@@
 CADRef  (Ling et al., 2025), ITP  (Xu and Yang, 2025), NCI  (Liu and Qin, 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Density-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] [ Test-time Adaptive Approaches
-(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-update needed,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-update free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
+(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-optimization-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-optimization-free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
 model-based OOD detection
@@
 Anomaly Detection.
-Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009; Xing et al., 2025; Wang and Nalisnick, 2025).
-These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003) or system failures  (Du et al., 2017).
-The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations.
-Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency.
-While both AD and OOD detection aim to identify unusual or unexpected data points, compared to OOD detection, AD differs in three key aspects.
-First, regarding supervision, AD methods are typically unsupervised, training only on normal samples without detailed labels and detecting anomalies during testing, whereas OOD detection involves labeled category discrimination.
-Second, AD focuses on perceptual anomalies, where images belong to known categories but contain anomalous regions, while OOD detection targets semantic differences, with test samples from unseen categories.
-Additionally, although both methods commonly use AUROC (Wehler et al., 2001) for evaluation, AD specifically employs AUPRO (Bertoldo et al., 2024) to assess pixel-level detection performance.
+Anomaly detection (AD) identifies data points that deviate from normal patterns (Chandola et al., 2009), with wide applications in fraud detection (Pourhabibi et al., 2020), network security (Lazarevic et al., 2003), and system monitoring (Du et al., 2017).
+While both AD and OOD detection identify unusual samples, AD differs in three main aspects.
+First, AD is typically unsupervised, using only normal samples for training, whereas OOD detection involves labeled categories.
+Second, AD focuses on perceptual anomalies, while OOD detects semantic shifts from unseen categories.
+Lastly, although both use AUROC (Wehler et al., 2001) for evaluation, AD employs AUPRO (Bertoldo et al., 2024) for pixel-level performance.
+Beyond these standard metrics, practical evaluation also assesses human-centric and operational criteria, such as detection latency (Lavin and Ahmad, 2015; Sinha et al., 2024), the economic cost of errors (Xu et al., 2013; Kim et al., 2022), model interpretability (Jiang et al., 2023c; Doshi and Yilmaz, 2023), and computational footprint (Ibidunmoye et al., 2015; Huertas-Garc´ıa et al., 2023).
 Novelty Detection.
-Novelty detection (ND) focuses on identifying new or unknown data points that a model has not seen during training  (Pimentel et al., 2014).
-This process is essential in situations where the system needs to adapt to evolving conditions or when it’s crucial to flag new, unseen scenarios.
-Unlike AD, which seeks to find patterns that deviate from the norm, novelty detection aims to discover entirely new patterns.
-Applications include identifying new trends in social media  (Li et al., 2017), discovering new species in biological data  (Pimentel et al., 2014), or detecting new topics in document streams  (Aksoy et al., 2012).
-Essentially, ND deals with surprises in familiar contexts, while OOD detection deals with data from unfamiliar contexts.
+Novelty detection (ND) aims to identify previously unseen data points (Pimentel et al., 2014), enabling systems to adapt to new conditions or scenarios.
+Unlike AD, which finds deviations within known patterns, ND discovers new patterns, such as emerging social media trends (Li et al., 2017), novel species (Pimentel et al., 2014), or new document topics (Aksoy et al., 2012).
+In essence, ND detects surprises within familiar contexts, while OOD detection addresses data from entirely unfamiliar contexts.
 Open Set Recognition.
-Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016).
-This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that are not present during training.
-Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully.
-OSR is concerned with identifying new classes that are not present during training, which is generally achieved by dividing the categories of the same dataset into base classes and new classes, which means that new classes and base classes usually come from the same domain.
-In contrast, OOD detection focuses on identifying any data that is different from the training distribution, regardless of whether it belongs to a new class or a different domain altogether.
-Essentially, OSR is a subset of OOD detection, specifically aimed at classifying new class types within the same domain.
+Open Set Recognition (OSR) extends traditional classification by identifying both known classes and instances from unknown categories (Bendale and Boult, 2016).
+Unlike OOD detection which identifies any OOD data, OSR specifically focuses on recognizing novel classes within the same domain.
+This makes it particularly relevant for applications like robotics (Marsland et al., 2005) and autonomous systems (Amini et al., 2018).
 Outlier Detection.
-Outlier detection (OD) is the process of identifying data points that are significantly different from the majority of the data  (Singh and Upadhyaya, 2012).
-It is similar to AD but focuses more on the identification of individual data points rather than patterns.
-Outliers can arise due to variability in measurement, experimental errors, or genuinely novel variations.
-Applications include fraud detection, fault detection, and removing anomalous data from datasets to improve model accuracy  (Pawar et al., 2014).
-Compared to OOD detection, OD is more of a transductive scenario, as it inherently has access to outliers.
-However, classical OOD detection only encounters outliers at the time of deployment.
+Outlier detection (OD) identifies individual data points that significantly deviate from the majority (Singh and Upadhyaya, 2012).
+Unlike OOD detection which encounters outliers only during deployment, OD is transductive with inherent access to outliers.
+It finds applications in fraud detection and data cleaning (Pawar et al., 2014).
 Zero-shot learning.
-Zero-shot learning is a paradigm in machine learning where the goal is to recognize objects without having seen examples of these objects during the training phase  (Romera-Paredes and Torr, 2015; Wang et al., 2019; Pourpanah et al., 2022).
-The fundamental challenge in zero-shot learning is to effectively transfer knowledge from seen to unseen classes  (Romera-Paredes and Torr, 2015).
-It has been primarily tackled by learning semantic relationships between classes through attributes or by embedding both seen and unseen classes into a shared semantic space.
-While both deal with unknowns during inference, zero-shot learning tries to classify new categories  (Xian et al., 2018), whereas OOD detection flags data that is anomalous or unfamiliar to the training data distribution  (Lee et al., 2018b).
+Zero-shot learning aims to recognize objects without seeing their examples during training by transferring knowledge from seen to unseen classes (Romera-Paredes and Torr, 2015; Wang et al., 2019).
+Unlike OOD detection which flags anomalous data (Lee et al., 2018b), zero-shot learning actively classifies new categories through semantic relationships (Xian et al., 2018).
 Selective Classification.
-Selective classification, also known as reject option classification, provides a mechanism for models to abstain from making a decision when they are not sufficiently confident in their predictions  (Geifman and El-Yaniv, 2017; Chang et al., 2024).
-Selective classification involves a model deciding when to make a prediction based on its confidence level, effectively choosing not to predict when uncertain  (Wiener and El-Yaniv, 2011).
-On the other hand, OOD detection identifies data points that differ from the training distribution, aiming to flag them as unfamiliar to the model.
-While both methods manage uncertainty, selective classification deals with confidence in making predictions on ID data (El-Yaniv et al., 2010; Chang et al., 2024), and OOD detection focuses on recognizing and handling data that is not represented in the training set.
+Selective classification (or reject option classification) allows models to abstain from predictions when confidence is insufficient  (Geifman and El-Yaniv, 2017; Chang et al., 2024).
+While OOD detection identifies samples from different distributions, selective classification focuses on managing prediction uncertainty within the training distribution  (El-Yaniv et al., 2010).
 Misclassification (Failure) Detection.
-Misclassification detection aims to identify incorrect predictions on in-distribution data, exposing a model’s unreliability in its intended domain (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d).
-This is especially important in safety-critical areas such as autonomous driving and healthcare.
-Failure detection generalizes this task to capture a broader range of model breakdowns, including uncertainty and data corruption.
-While both misclassification and OOD detection seek to improve model reliability, the former focuses on errors within the training distribution (Jaeger et al., 2023), whereas the latter targets inputs outside it.
+Misclassification detection identifies incorrect predictions on ID data (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d), critical for domains like autonomous driving.
+It differs from OOD detection by focusing on errors within the training distribution (Jaeger et al., 2023), rather than external inputs.
 3. Problem: Training-driven OOD Detection
@@
 Reconstruction-based.
-The reconstruction-based approach detects OOD samples by measuring the difference between representations of data before and after reconstruction, relying on the model’s ability to accurately recover input semantics.
-The core idea is that OOD data contain semantic features that differ from those of ID data.
-By quantifying these differences, the model can effectively distinguish OOD samples.
-An overview is illustrated in Fig.
+Reconstruction-based OOD detection methods focus on measuring representational differences between original and reconstructed data.
+These approaches leverage the model’s capability to recover input semantics, operating on the premise that OOD samples contain distinct semantic features from ID data, as shown in Fig.
 3 (a).
-MoodCat  (Yang et al., 2022) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process.
-Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data.
-Following this, MOOD  (Li et al., 2023a) and MOODv2  (Li et al., 2024a) leverage pretraining tasks based on masked image modeling, exhibiting significant advantages in learning the internal distribution of data.
-PRE  (Osada et al., 2023) introduces the concept of normalized flow, combined with a penalty based on typicality to constrain reconstruction errors, which discerns differences between OOD and ID data well.
+Several notable works have advanced this direction.
+MoodCat  (Yang et al., 2022) employs masked image modeling with constrained synthesis based on classification outputs.
+Zhou (2022) enhances detection by incorporating feature activation analysis through an auxiliary module.
+Building upon masked image modeling, MOOD  (Li et al., 2023a) and its successor MOODv2  (Li et al., 2024a) demonstrate improved capabilities in learning data distributions.
+PRE  (Osada et al., 2023) further develops this approach by combining normalized flow with typicality-based penalties to effectively distinguish between OOD and ID samples.
 Figure 3.
-Illustration of some subclasses within the Training-driven approaches: (a) Reconstruction-based : After extracting the feature representation $F(x)$ of the image, it is fed into a reconstruction network (e.g., VAE or DDPM) to obtain the reconstructed feature representation $F^{\prime}(x)$.
-By comparing the dissimilarity between the original and reconstructed images, OOD data can be identified, as OOD samples typically exhibit higher dissimilarity before and after reconstruction.
-(b) Prototype-based : The model learns more optimal ID prototypes during the training process to distinguish OOD data.The symbol $h$ typically denotes the features of the penultimate layer in the network.
-(c) OOD Synthesis : During training, real OOD data is not available; instead, OOD data is synthesized by extrapolating information from the ID data.
-(d) Outlier Mining : During training, real OOD data is available, but typical OOD samples are generated using this real OOD data.
-(e) Boundary Regularization : The model’s decision boundary is optimized by using real OOD data, typically without involving the synthesis of OOD data.
-Recently, diffusion models (Ho et al., 2020) have made remarkable strides in both training stability and the quality of generated images.
-Leveraging diffusion models to detect OOD data has become a new research direction.
+Illustration of main Training-driven OOD detection approaches: (a) Reconstruction-based: Identifies OOD samples by comparing the dissimilarity between original feature representation F(x) and its reconstruction F’(x) through networks like VAE or DDPM.
+(b) Prototype-based: Learns optimal ID prototypes during training for OOD distinction, where h represents penultimate layer features.
+(c) OOD Synthesis: Generates synthetic OOD samples by extrapolating from ID data when real OOD data is unavailable.
+(d) Outlier Mining: Leverages available real OOD data to generate typical OOD samples during training.
+(e) Boundary Regularization: Optimizes model decision boundary using real OOD data without synthesis.
+Recent advances in diffusion models (Ho et al., 2020) have led to improved training stability and image quality, prompting their application in OOD detection.
 Graham et al.
-(2023) address this issue by introducing DDPM, utilizing it for the reconstruction of noise-disturbed images.
-The model employs multidimensional reconstruction error for the identification of OOD data.
-The information bottleneck of this model can be externally regulated.
-Similarly, LMD  (Liu et al., 2023a) utilizes diffusion models for OOD validation.
-LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold.
-DiffGuard  (Gao et al., 2023b) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
+(2023) introduce DDPM for reconstructing noise-corrupted images, using multidimensional reconstruction error for OOD identification and allowing external regulation of the information bottleneck.
+LMD (Liu et al., 2023a) also applies diffusion models, disrupting and reconstructing data to separate OOD samples from the original manifold.
+DiffGuard (Gao et al., 2023b) leverages pre-trained diffusion models to enhance semantic discrepancies between reconstructed OOD and original images.
 Probability-based.
@@
 Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models.
-It measures whether data should be classified as in-distribution by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
+It measures whether data should be classified as ID by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
 Logits-based.
@@
 LogitNorm  (Wei et al., 2022a) uses logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence.
-Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores.
-It adjusts the learning intensity between samples during training, rendering the learning process robust.
-DML  (Zhang and Xiang, 2023) challenges the potential hindrance to OOD detection performance.
-Drawing on empirical insights, DML enhances OOD detection performance by decoupling logits and balancing individual components, thereby mitigating the impact of each attribute on the results.
-It enhances OOD detection performance by decoupling logits and scenario MaxNorm as a parameter, thereby balancing the influence of each attribute on the outcome.
+UE-NL (Huang and Qiao, 2023) leverages Bayesian principles to jointly learn embeddings and uncertainty scores, normalizing logits to improve robustness in OOD detection.
+DML (Zhang and Xiang, 2023) further advances this area by decoupling logits and balancing key components, thereby mitigating attribute interference and enhancing detection performance.
 OOD Synthesis.
@@
 Lee et al.
-(2018a) add additional regularization terms by incorporating two supplementary terms into the original loss function.
-The first term compels the classifier to increase uncertainty for OOD samples, while the second term encourages the GAN to generate appropriate OOD samples.
-Specifically, a joint training scheme is designed to select suitable boundary samples in the low-density regions of ID samples.
-Similarly, VOS  (Du et al., 2022a) employs adaptive synthesis of anomaly values.
-It samples anomaly values from the low-likelihood regions of class-conditional distributions in the feature space, eliminating the need for reliance on external data.
-This enables the model to synthesize OOD data for predictions.
-It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention.
-Noticing the additional distributional assumptions posed by previous methods for OOD data, some researchers approach the problem from another perspective.
-For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality.
-It selects boundary points through non-parametric density estimation based on nearest neighbors, with these boundary points situated between ID and OOD data.
-Additionally, SSOD  (Pei, 2024) introduces self-supervised sampling to implicitly generate OOD data.
-It directly samples natural OOD signals from the backgrounds of ID data images, thus overcoming limitations introduced by biases during the OOD synthesis stage.
-In contrast to the implicit generation methods of OOD samples mentioned above, some approaches take a different route by directly leveraging explicit construction of OOD samples from ID samples for learning.
-To acquire OOD samples, CODEs  (Tang et al., 2021) initially generate seed OOD examples by slicing and stitching samples from internal distributions of different categories.
-These examples are then inputted into Chamfer GAN for distribution transformation, yielding high-quality OOD samples.
-CMG  (Wang et al., 2022a) then generates pseudo OOD data by providing class embeddings mixed as conditions to a Conditional Variational Autoencoder (CVAE), and subsequently utilizes this data to fine-tune the classifier for OOD detection.
-SEM  (Yang et al., 2023b) introduces a problem setting for full-spectrum OOD detection, where negative samples are generated using Mixup operations during training.
-It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data.
-Forte (Ganguly et al., 2025) utilizes per-point statistics (such as precision, recall, and density) to capture the local geometric properties of the data manifold in the feature space and performs typicality estimation using density-based non-parametric methods (such as Gaussian mixture models).
-This approach helps improve the accuracy of OOD synthetic data detection.
-SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples.
-It achieves this by employing CLIP to eliminate ID object regions within the training samples.
-The latent diffusion model is then utilized to substitute these regions with authentic features, all the while taking into account the contextual background.
-This methodology thereby establishes the model’s capability for rejection.
-However, ensuring the quality of generated data is often challenging.
-Furthermore, the quality of the generated data itself may have inherent flaws.
-ATOL  (Zheng et al., 2023) introduces an auxiliary task, wherein both auxiliary ID and auxiliary OOD data coexist.
-In a low-dimensional latent space, distinct regions are manually sought for both auxiliary ID and auxiliary OOD data, ensuring non-overlapping characteristics in the input space.
-Subsequently, data generation by the generator guarantees the non-overlapping properties of the input space.
-Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances.
-In recent studies, HamOS (Li and Zhang, 2025) employs Hamiltonian Monte Carlo (Duane et al., 1987) sampling, treating the synthesis process as sampling from a Markov chain.
-In the absence of OOD data, it synthesizes high-quality OOD virtual anomaly points by constructing a Markov chain (Neal, 1993).
-It projects the features onto a hyperspherical feature space and uses K-nearest neighbor distances to guide the generation of anomaly points.
-POP (Gong et al., 2025) does not require the synthesis of specific outliers, but instead directly reshapes the decision boundaries through virtual OOD prototypes, improving the model’s OOD detection capability by penalizing misclassifications.
+(2018a) and VOS (Du et al., 2022a) both synthesize OOD samples from low-density or low-likelihood regions of the ID feature space, but VOS enables adaptive outlier generation without external data or manual tuning.
+In contrast, NPOS (Tao et al., 2023) avoids distributional assumptions by using non-parametric density estimation to select boundary samples, enhancing flexibility and generality.
+SSOD (Pei, 2024) further differs by leveraging self-supervised sampling to extract OOD signals directly from ID backgrounds, addressing synthesis bias.
+OOD detection methods often generate synthetic outliers from in-distribution data.
+Techniques include recombining features with GANs (CODEs (Tang et al., 2021)), using conditional VAEs (CMG (Wang et al., 2022a)), or applying Mixup (SEM (Yang et al., 2023b)).
+Alternatively, other approaches estimate data typicality by analyzing the local geometric and density statistics of the feature manifold (FORTE (Ganguly et al., 2025)).
+SHIFT (Kwon et al., 2023) leverages CLIP and latent diffusion models to generate OOD images by replacing ID object regions with contextually consistent features, though maintaining high sample quality remains a challenge.
+ATOL (Zheng et al., 2023) introduces an auxiliary task that constructs non-overlapping regions for ID and OOD data in latent space, using generated samples to reinforce this separation and aligning genuine ID data to enhance reliability.
+Recent studies show that HamOS (Li and Zhang, 2025) uses Hamiltonian Monte Carlo (Duane et al., 1987) to generate virtual OOD anomalies by sampling from a Markov chain (Neal, 1993) when real OOD data is unavailable.
+It projects features onto a hypersphere and leverages K-nearest neighbor distances to guide anomaly generation.
+POP (Gong et al., 2025) instead refines decision boundaries using virtual OOD prototypes and penalizes misclassifications, enhancing OOD detection without explicit outlier synthesis.
 Prototype-based.
@@
 The optimization process during model training pertains to the prototypes of classes.
-AROS (Mirzaei and Mathis, 2025) focuses more on the feature separation between ID and OOD data.
-It proposes using the Lyapunov stability theorem to constrain the stability of the neural network’s embedding space, ensuring that the embeddings of ID and OOD data converge to distinct stable equilibrium points.
-To obtain OOD data, it also employs a technique of sampling from the low-probability regions of the ID data to generate pseudo-OOD.
-CNC (Harun et al., 2025) proposed controlling the degree of neural collapse (NC) at different layers of the DNN network.
-In the Encoder, entropy regularization is introduced to mitigate NC, while in the Projector, the projection head is used to enhance the NC phenomenon, forming more compact prototype and thereby improving OOD detection performance.
+AROS (Mirzaei and Mathis, 2025) enhances feature separation between ID and OOD data by applying the Lyapunov stability theorem, guiding embeddings of ID and OOD samples toward distinct stable points.
+Pseudo-OOD samples are generated by sampling from low-probability regions of the ID distribution.
+CNC (Harun et al., 2025) controls neural collapse (NC) at different network stages: entropy regularization in the encoder alleviates NC, while the projector head promotes NC, resulting in more compact prototypes and improved OOD detection.
 ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores.
@@
 Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning.
-However, PALM (Lu et al., 2024) observes that modeling each class with a single prototype may fail to capture the diversity of internal information within the data.
-PALM introduces a mixed-prototype strategy to optimize the modeling process, utilizing a strategy of multiple prototypes to learn informative representations for each class.
-By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function.
-PFS (Wu et al., 2025) also introduces a clustering loss that encourages features of the same ID class to cluster more closely, thereby improving the consistency of ID data representation.
+However, PALM (Lu et al., 2024) notes that representing each class with a single prototype often fails to capture the internal diversity of data.
+To address this, PALM adopts a mixed-prototype strategy, assigning multiple prototypes per class to better model informative representations.
+By jointly learning class-level prototypes and contrasting them across classes, PALM’s loss function encourages intra-class compactness and inter-class separability at prototype level.
+Similarly, PFS (Wu et al., 2025) introduces a clustering loss that further promotes tight feature clustering within each ID class, enhancing the consistency of ID representations.
 A special case: Long-tail ID data.
@@
 Many approaches are proposed to address the challenge of ID imbalance and enhance OOD detection capabilities.
-OLTR  (Liu et al., 2019) addresses the robustness of tail recognition by associating visual concepts between the head and tail embeddings.
-It accomplishes this by dynamically adjusting embedding norms through the utilization of visual memory.
-The approach involves two components to enhance the robustness of long-tailed data: one utilizes direct features from standard embeddings, while the other employs memory features that store discriminative centroids derived from direct features, thereby augmenting the direct features of the images.
-Mehta et al.
-(2022) propose a tailored complex subset mixing strategy designed specifically for handling medium-to-tail classes.
-This method creates mixed samples by pairing two independent samples from different tail categories and calculates the distance between the mixed sample and prototypes of specific classes.
-The mixing strategy is ultimately combined with prototype learning to effectively address the challenges posed by long-tail scenarios.
-AREO  (Sapkota and Yu, 2023) introduces a method to quantify sample uncertainty through evidential learning, akin to confidence.
-The entire training process is governed by an innovative multi-scheduler learning mechanism, dynamically adjusting training parameters based on the importance of different classes to ensure the model focuses on the features of minority classes, thus balancing the disparities between majority and minority classes.
-Overall, AREO identifies higher uncertainty in certain samples during training and dynamically adjusts parameters to mitigate overfitting to these samples.
-Existing OOD detection methods predominantly assume a uniform distribution of probabilities from OOD to ID.
-In the presence of class imbalances,  Jiang et al.
-(2023a) propose an alternative strategy to enhance current OOD detection methodologies.
-This involves recalibrating the OOD score based on the class prior distribution of the ID and the KL divergence from the output of a pre-trained model.
-This approach signifies the potential to augment the robustness of OOD detection beyond the achievements of previously optimized methodologies, further fortifying the overall performance.
-Additionally, Open-Sampling  (Wei et al., 2022b) leverages noisy labels from the OOD dataset to rebalance the class priors of the ID training dataset.
-These labels are sampled from a predefined distribution that complements the original class prior distribution.
-COOD  (Hogeweg et al., 2024) uses the supervised model to combine individual OOD measures into a single ensemble, similar to the concept of random forests.
-This approach addresses the limitations of individual OOD methods and can also overcome issues related to data imbalance.
+OLTR (Liu et al., 2019), POP (Mehta et al., 2022), and AREO (Sapkota and Yu, 2023) each address long-tailed recognition from different perspectives.
+OLTR enhances tail class robustness by leveraging visual memory to bridge head and tail embeddings.
+POP focuses on medium-to-tail classes through a prototype-based mixing strategy, generating mixed samples to improve class separation.
+In contrast, AREO models sample uncertainty via evidential learning and dynamically adjusts training using a multi-scheduler mechanism to better balance majority and minority classes.
+Most existing OOD detection methods assume a uniform probability distribution between OOD and ID samples.
+To address class imbalance, Jiang et al.
+(2023a) recalibrate OOD scores by incorporating class priors and KL divergence, improving robustness under skewed distributions.
+Open-Sampling (Wei et al., 2022b) rebalances ID class priors using noisy OOD labels sampled from a complementary distribution.
+In contrast, COOD (Hogeweg et al., 2024) ensembles multiple OOD measures via a supervised model, effectively mitigating individual method limitations and handling data imbalance.
+Figure 4.
+Illustration of training-agnostic OOD detection approaches.
+Both methods require access to a pre-trained model.
+Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing.
+“A’/B’ ” means that the original features are deformed; “Bank” means that some samples are stored; “Clustering” means that clustering is performed for ID images.
 3.2. OOD Detection Approaches with Both ID and OOD Data
@@
 3 (d).
-POEM  (Ming et al., 2022a) raises methodology leans towards excavating anomalies that carry more representative characteristics.
-Li and Vasconcelos (2020) propose a method of data resampling to acquire representative outlier data for training, assigning higher priority score reweighting to hard negative instances.
-Iterative optimization, leveraging adversarial principles, selects target data.
-POEM employs posterior sampling to unearth anomalies with elevated boundary scores from an extensive auxiliary dataset, facilitating a nuanced comprehension of the intricacies within OOD samples.
-DAOL  (Wang et al., 2023a) takes the disparity between authentic data and auxiliary OOD data as a starting point to formulate an OOD dataset.
-Utilizing the Wasserstein ball, it models the distribution of all OOD data, selecting the most challenging OOD data within the ball for training.
+Different approaches have been proposed for selecting representative OOD samples: POEM  (Ming et al., 2022a) utilizes posterior sampling to identify high boundary score anomalies, while  Li and Vasconcelos (2020) employs data resampling with priority score reweighting for hard negative instances.
+In contrast, DAOL  (Wang et al., 2023a) models OOD distribution using Wasserstein balls to select challenging OOD samples based on the disparity between real and auxiliary data.
 Beyond solely relying on raw data, another direction in addressing this issue involves synthesizing representative outlier data by utilizing authentic OOD data through information extrapolation.
@@
 Long-tail ID data.
-Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information.
-The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two.
-To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024a) incorporates a learnable tail class prototype during the training process.
-This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.
+Recent studies have explored OOD detection in scenarios with imbalanced ID training data.
+Different approaches have been proposed to address this challenge: PASCAL  (Wang et al., 2022b) uses contrastive loss to separate tail-class from OOD data, while COCL  (Miao et al., 2024a) introduces a learnable tail class prototype to better distinguish between them.
 Choi et al.
-(2023) take a different approach, suggesting that factors affecting the performance of OOD detection may be related to the imbalance in the cross-class distribution of auxiliary OOD data.
-Consequently, it proposes an energy regularization loss, specifically regularizing auxiliary samples from the majority classes to address the issue of class imbalance in OOD data.
-EAT  (Wei et al., 2024) utilizes dynamically assigned virtual labels during the model training process to train OOD data, thereby expanding the classification space.
+(2023) focuses on balancing cross-class distribution of auxiliary OOD data through energy regularization, and EAT  (Wei et al., 2024) expands the classification space by dynamically assigning virtual labels to OOD data during training.
 Despite the success and considerable attention received by methods like Outlier Exposure in the research community, there are voices questioning the essence of allowing access to OOD data during training.
@@
 In contrast, test-time adaptive methods leverage information shared among test samples to enhance OOD detection.
-Figure 4.
-Illustration of training-agnostic OOD detection approaches.
-Both methods require access to a pre-trained model.
-Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing.
-“A’/B’ ” means that the original features are deformed; “Bank” means that some samples are stored; “Clustering” means that clustering is performed for ID images.
 4.1. Post-hoc OOD Detection Approaches
@@
 For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category.
-Departing from the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by utilizing the maximum value of the logits vector generated by the neural network.
-These logits constitute the raw, pre-Softmax layer outputs of the neural network, representing the model’s confidence in each category without undergoing Softmax transformation.
-Alternatively, the Energy  (Liu et al., 2020a) method defines an energy score conceptually derived from Helmholtz Free Energy.
-This score can theoretically reflect the input probability density.
-By assigning lower energy scores to ID data, this method distinguishes them from anomalous data.
-Additionally, it demonstrates reduced susceptibility to model overconfidence issues.
+Departing from the MSP method, MaxLogits (Hendrycks et al., 2022) detects OOD samples using the maximum logit value, while Energy (Liu et al., 2020a) employs an energy-based score related to input probability density.
+Compared to MaxLogits, Energy is less sensitive to model overconfidence and provides a more theoretically grounded metric.
 GEN  (Liu et al., 2023b) introduces the concept of generalized entropy and directly utilizes Bregman divergence to compute the statistical distance between the model’s probability output and uniform distribution, aiming to identify OOD data.
@@
 Gradient-based.
-Grad  (Lee and AlRegib, 2020) research indicates that gradient-based methods also contribute to OOD detection by quantifying model uncertainty through the gradients propagated in reverse during backpropagation.
-If the input sample is an ID sample, the gradients of the model with respect to these samples tend to be relatively small and stable.
-Conversely, the gradients are typically larger or more irregular for OOD samples.
-GradNorm  (Huang et al., 2021) posits that the gradient magnitude for ID data surpasses that of OOD data.
-Leveraging this observation, it employs gradient vector norms, computing KL divergence through softmax output and uniform distribution backpropagation to detect OOD data.
-In contrast, GradOrth  (Behpour et al., 2023) adopts an alternative perspective, recognizing that crucial features of OOD data reside in a low-rank subspace.
-Consequently, it shifts focus to calculating gradient projection norms in this subspace to identify OOD data.
-GAIA  (Chen et al., 2023) employs a combination of gradient anomaly checks and aggregation.
-It enables the model to interpret indeterminacy attributively, introducing channel-wise average abnormality and zero-deflation abnormality without prior knowledge to gauge the extent of data distribution variations.
-The OPNP  (Chen et al., 2024a) method discovers that the OOD detection capability of the model is highly sensitive to parameters and activated neurons that are close to zero.
-Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability.
+Gradient-based methods leverage model gradients to quantify uncertainty.
+Grad (Lee and AlRegib, 2020) and GradNorm (Huang et al., 2021) both utilize gradient magnitudes, with the latter specifically comparing KL-divergence-based gradients between ID and OOD samples.
+GradOrth (Behpour et al., 2023), instead of focusing on overall gradient norms, projects gradients onto low-rank subspaces to capture OOD features.
+GAIA (Chen et al., 2023) introduces channel-wise and zero-deflation abnormality scores to assess distribution shifts without prior knowledge.
+OPNP (Chen et al., 2024a) further improves OOD detection by pruning parameters and neurons near zero, thus enhancing generalization.
 In recent research, PRO (Chen et al., 2025a) introduces an adversarial score, with the core idea being to perturb the input data using gradient descent to search for the local minimum score near the original input.
@@
 Feature-based.
-Another avenue of research into OOD detection involves the impact of neural network intermediate variables relative to the final prediction.
-Inspired by adversarial examples, ODIN  (Liang et al., 2018a) experimentally discovered that adding small perturbations to the input image features can more effectively detect OOD data.
+Research in this direction explores the role of intermediate neural features for OOD detection.
+ODIN (Liang et al., 2018a), inspired by adversarial examples, demonstrates that small input perturbations can enhance OOD detection performance.
 The perturbation formula is as follows:
@@
 Due to the local neuroplasticity characteristic of Kolmogorov-Arnold Networks (Liu et al., 2025b), KANs distinguish between ID and OOD data by comparing the activation response differences between the trained and untrained KANs.
-SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison.
-It employs a storage mechanism to calculate the average logits for each class, preparing this simple storage pattern for subsequent OOD detection tasks.
-Drawing inspiration from experience, DDCS  (Yuan et al., 2024) adapts to select suitable channels for data classification after correcting each channel in the neural network.
-These channels are evaluated based on inter-class similarity and variance to measure their discriminative power for ID data.
-LINe  (Yong Hyun Ahn, 2023) similarly emphasizes neuron outputs at the feature level.
-It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs.
+SHE (Zhang et al., 2022) introduces a storage-based approach by leveraging Hopfield energy (Krotov and Hopfield, 2016) to facilitate OOD detection, averaging logits per class to form a reference pattern.
+Building on channel selection, DDCS (Yuan et al., 2024) evaluates and corrects neural network channels according to their discriminative ability, using inter-class similarity and variance.
+Similarly, LINe (Yong Hyun Ahn, 2023) focuses on neuron-level feature outputs, employing Shapley value pruning to retain only the most informative neurons for OOD detection.
 CADRef (Ling et al., 2025) decouples the features by using a strategy based on the sign alignment between the relative features and model weights, dividing the sample features into positive and negative errors.
@@
 Additionally, during the testing phase, ITP uses a right-tail Z-score test to assess whether any parameter exhibits overconfidence in the classification of a given test sample, further refining the model’s decision-making.
-Feature shaping is emerging as a prominent area of research focus.
-This methodology entails the refinement of intermediate representations—specifically, the intermediate features—within the forward propagation phase of the model.
-The salient advantage of this approach lies in its non-disruptive nature concerning the original classification results, coupled with its simplicity and efficacy.
-Building upon this foundation, ViM  (Wang et al., 2022c) simultaneously considers the roles of features, logits, and probabilities, constructing virtual logits to aggregate information for decision-making.
-By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection.
-Its computation of this score $S$ can be expressed in the following form:
+Feature shaping refines intermediate features during forward propagation, offering a simple and effective way to enhance OOD detection without affecting original classification results.
+Building on this, ViM (Wang et al., 2022c) integrates features, logits, and probabilities by constructing virtual logits, and leverages the penultimate layer’s null space—irrelevant to classification yet highly effective for OOD detection.
+Its computation of score $S$ can be expressed in the following form:
 (5) $S=-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
@@
 Therefore, NCI leverages the proximity between features and class weight vectors to effectively distinguish ID samples from OOD samples, while maintaining low computational cost.
-ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted.
-NAC (Liu et al., 2024b) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data.
-By quantifying this statistical property, NAC aims to distinguish between ID and OOD data.
-Based on the existing literature on feature shaping,  Zhao et al.
-(2024) utilize a piecewise constant shaping function to partition the feature domain into disjoint intervals, estimating a scalar within each interval as an approximation.
-As the interval width approaches zero, an approximation of the maximum logits can be obtained.
-The goal of BLOOD  (Jelenić et al., 2024) is to smooth the representations of intermediate layers in neural networks for predicting OOD data.
-This study found that, compared to OOD data, ID data exhibits smoother variations in the intermediate layer representations of neural networks.
-Leveraging this characteristic, new statistical measures can be designed to discriminate against anomalous data.
-In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH.
-SCALE  (Xu et al., 2024a) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data.
-Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.
+ASH (Djurisic et al., 2023) and neuron activation pruning focus on direct modification or removal of activations, with ASH offering a simple dynamic scheme and pruning methods building upon its foundation.
+In contrast, NAC (Liu et al., 2024b) introduces neuron activation coverage as a statistical indicator, aiming to capture OOD likelihood based on rarely activated neurons.
+BLOOD (Jelenić et al., 2024) and related methods leverage the smoother intermediate representations of ID data, using this property to design new statistical measures for OOD discrimination.
+Feature shaping approaches (Zhao et al., 2024) partition the feature space and estimate piecewise approximations to logits, while SCALE (Xu et al., 2024a) highlights the importance of scaling metrics and observes lower pruning rates for OOD samples.
+Overall, these methods differ in whether they reshape activations directly, exploit statistical properties, or design new metrics for improved OOD detection.
 Density-based.
-Recent advancements in density-based OOD detection models have demonstrated substantial performance gains, predicated on the models’ ability to accurately capture and understand the intrinsic characteristics of the true data distribution.
-For example, GEM  (Morteza and Li, 2022) models the feature space of ID data as a class-conditional multivariate Gaussian distribution.
-Under this assumption, it designs new statistical metrics to validate the model’s performance.
-Using a modeling approach based on Gaussian mixture models, GEM Score is aligned with the true log-likelihood for capturing OOD uncertainty.
-However, while GEM strictly relies on Gaussian assumptions, recent works ConjNorm  (Peng et al., 2024) introduce a novel framework based on Bregman divergence, extending the consideration of data distributions to encompass an exponential family of distributions.
-By redefining density functions for data, this model has a broader range of applications.
+Recent density-based OOD detection methods have achieved notable improvements by more effectively modeling the true data distribution.
+For example, GEM (Morteza and Li, 2022) employs class-conditional Gaussian distributions with statistical metrics for model validation.
+Building upon this, ConjNorm (Peng et al., 2024) extends the framework to exponential family distributions using Bregman divergence, enabling broader applications.
 4.2. Test-Time Adaptive OOD Detection Approaches
 Overview.
-Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize unlabeled test data, either the complete test set or a series of mini-batches, to enhance OOD detection performance through model adaptation.
+Recently, many works (Yu et al., 2024b, 2023; Sun et al., ; Yu et al., 2025; Liang et al., ) have focused on test-time adaptive methods, where a model, trained on the training set, adapts during the testing phase.
+In the context of OOD detection, these methods aim to leverage unlabeled test data—whether the complete test set or mini-batches—to improve performance through model adaptation.
 Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge can be challenging and potentially limited.
@@
 Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
+Figure 5.
+Overview and trends in LPM-based OOD Detection Methods.
+(a) presents statistical data on influential research works in recent years, which involves papers indexed in prominent journals or conferences in the field.
+In the bar chart, the lower and upper parts represent the total number of papers in the first and second halves of each year, respectively.
+(b) shows a trend pie chart illustrating the proportion of keyword categories related to the research works in that year.
 Model-optimization-based. Model optimization-based methods enhance the trained model by leveraging unlabeled data during the post-training phase.
-A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024a) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection.
-The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them.
-Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024a) explains how they help OOD detection from the lens of separability and learnability.
-Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model.
-In addition, the extra training requirement in WOODS, thought necessary, incurs significant costs and is generally undesirable.
-Another line of model-optimization-based approaches draws inspiration from the semi-supervised-learning (SSL) techniques  (Zhu, 2005), aiming for a more efficient, lightweight training process during the post-training phase.
-Pseudo-labeling  (Arazo et al., 2020) is a simple but effective way to label the test data, which enhances learning from test unlabeled data.
-The method proposed by  Yang et al.
-(2023a), termed AUTO, employs only pseudo-OOD data to refine the model.
-The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification.
-In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model.
-SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
+A line of methods (Katz-Samuels et al., 2022; Du et al., 2024a) leverage “wild data”—a mix of unlabeled ID and OOD samples—to enhance OOD detection.
+WOODS (Katz-Samuels et al., 2022) focuses on cleaning wild data for reliable OOD regularization, while SAL (Du et al., 2024a) analyzes its impact through separability and learnability.
+However, wild data may introduce unintended information if it differs from test OOD, and additional training requirements can increase computational costs.
+Another line of model-optimization-based methods draw inspiration from semi-supervised-learning techniques  (Zhu, 2005) for efficient post-training.
+While basic pseudo-labeling  (Arazo et al., 2020) labels test data directly, AUTO  (Yang et al., 2023a) uniquely focuses on pseudo-OOD data with semantic consistency.
+In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) utilize both pseudo-ID and pseudo-OOD data, with SODA implementing dual-loss optimization and ATTA applying differential weighting strategies.
 Model-optimization-free.
 Modifying the original trained model is infeasible in certain security-sensitive scenarios.
-Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” techniques, are increasingly garnering interest.
+Therefore, methods enabling test-time adaptation without requiring model updates, termed “model-optimization-free” techniques, are increasingly garnering interest.
 These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
-Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself.
-ETLT observes a linear correlation between the feature representation and the OOD score of a test input.
-In other words, for a given image, the pair consisting of its feature and OOD score (feature, OOD score) exhibits a linear correlation.
-Furthermore, the aforementioned pairs of ID and OOD data is linearly separable.
-Based on the observation, ETLT proposes to learn a linear regression module trained from the (feature, OOD score) pair.
-The availability of complete test datasets may not always be feasible, therefore,  Fan et al.
-(2022) also provide an online variant to ensure safer deployment.
-Similarly, GOODAT develops an add-on named graph masker, designed specifically for graph data.
-It integrates GIB-boosted losses and employs it as the metric for OOD scoring.
-In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach.
-Similarly, OODD (Yang et al., 2025) constructs a dynamic OOD dictionary, managed by a priority queue, which continuously updates with feature representations of test samples based on their OOD scores.
-The core principle behind AdaOOD and OODD is the maintenance of a memory bank, similar to the approach taken by AUTO.
+Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) preserve model integrity by training add-on modules for OOD score adjustment.
+ETLT leverages the linear correlation between features and OOD scores, implementing both offline and online variants, while GOODAT develops a graph-specific masker with GIB-boosted losses.
+In contrast, AdaOOD  (Zhang et al., 2023b) and OODD  (Yang et al., 2025) adopt non-parametric approaches, utilizing memory banks with k-nearest neighbours and dynamic dictionaries, respectively.
 Online v.s.
 Offline.
-Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment.
+Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) emphasize the offline scenario, where OOD detectors remain static and fixed after deployment.
 In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023a) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
@@
 The meta-learning approach is employed to swiftly adapt models in response to the complexities encountered in various scenarios in MOL.
-Figure 5.
-Overview and trends in LPM-based OOD Detection Methods.
-(a) presents statistical data on influential research works in recent years, which involves papers indexed in prominent journals or conferences in the field.
-In the bar chart, the lower and upper parts represent the total number of papers in the first and second halves of each year, respectively.
-(b) shows a trend pie chart illustrating the proportion of keyword categories related to the research works in that year.
 5. Problem: Large Pre-trained Model-based OOD Detection
@@
 The performance evaluations of several relevant competitive methods are summarized in Table  3 to provide an understanding of the performance level of OOD detection in this area.
+5.1. Zero-shot OOD Detection Approaches
+Overview.
+Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data.
+Note that we rely solely on textual category knowledge, without the need for access to ID images.
+It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training.
+Existing research on zero-shot OOD detection using vision-language models (VLMs-based) can be broadly categorized into two main approaches, depending on the underlying models employed: diffusion-based and CLIP-based methods.
 Figure 6.
@@
 None of these approaches use labels or images of OOD categories.
-5.1. Zero-shot OOD Detection Approaches
-Overview.
-Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data.
-Note that we rely solely on textual category knowledge, without the need for access to ID images.
-It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training.
-Existing research on zero-shot OOD detection using vision-language models (VLMs-based) can be broadly categorized into two main approaches, depending on the underlying models employed: diffusion-based and CLIP-based methods.
 Diffusion-based.
@@
 (6) $S_{MCM}(x^{\prime};Y_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}}.$
-where $x’$ ’ denotes the test image, $Y_{in}$ is the set of in-distribution labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
-Recent works have further enriched the textual context used by CLIP to improve OOD detection.
-For example, NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024) all leverage curated corpora to augment the set of candidate labels.
+where $x’$ ’ denotes the test image, $Y_{in}$ is the set of ID labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
+Several works have enriched the textual context in CLIP-based OOD detection by augmenting candidate labels with curated corpora, as seen in NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024).
 Specifically, NegLabel selects negative labels from a corpus that are semantically dissimilar to ID classes, thereby sharpening the model’s ability to distinguish OOD samples.
@@
 It is worth emphasizing that methods such as ZOC (Esmaeilpour et al., 2022), CLIPN (Wang et al., 2023c), NegLabel (Anonymous, 2024), and CLIPScope (Fu et al., 2024) each propose distinct OOD scoring mechanisms tailored to their respective architectures.
-In contrast, the MCM score stands out for its generalizability and can be seamlessly integrated into diverse CLIP-based frameworks.
-Recent advances in zero-shot OOD detection have moved beyond static negative sampling by leveraging more adaptive strategies within the joint vision-language space.
-For example, AdaNeg (Zhang and Zhang, 2024) dynamically generates negative proxies at test time to better match the semantics of OOD samples, while AdaND (Cao et al., 2025) focuses on mitigating the impact of noisy samples by decoupling noise detection from classification using CLIP-based pseudo-labels.
-In contrast, SimLabel (Zou et al., 2025) distinguishes OOD inputs by analyzing the consistency of image-label similarities among semantically related classes, automatically constructing similar class groups via hierarchical structures and large language models.
-These approaches thus differ in their strategies for negative sampling, noise handling, and semantic grouping, reflecting the diverse directions in recent OOD detection research.
-Complementing these proxy- and consistency-based advances, several works have proposed to enhance OOD detection by expanding the diversity and expressiveness of negative or outlier labels.
-OLE (Ding and Pang, 2024) exposes CLIP to a large set of outlier labels, selected from external corpora and clustered into prototypes, to better model the outlier space.
-Furthermore, OLE augments these prototypes through interpolation with ID representations, effectively filling the semantic gap between ID and OOD regions in the embedding space.
-Similarly, recent work on prompt learning for near-OOD detection (Jung et al., 2024) introduces novel scoring functions, such as the Empty-Class Score, which measures the alignment between an image and the context vector excluding any class name, and relative scores that subtract this alignment from standard logit-based OOD scores.
-This simple post-hoc adjustment provides substantial gains, especially for challenging near-OOD cases, without requiring retraining or architectural changes.
-Beyond label- and score-centric approaches, matrix approximation techniques have also been explored.
-SeTAR (Li et al., 2024b) applies selective low-rank approximation to CLIP’s weight matrices, aiming to retain the most informative components for OOD discrimination while filtering out spurious correlations.
-In a different vein, OT-DETECTOR (Liu et al., 2025a) formulates OOD detection as an optimal transport problem, quantifying the semantic and distributional discrepancy between test samples and ID labels, and further refining detection by semantic-aware content refinement.
-Collectively, these developments illustrate a clear methodological evolution: from static post-hoc scoring and negative label augmentation, to adaptive proxy generation, label consistency modeling, prototype-based outlier exposure, and low-rank or optimal transport-based representations.
-This progression not only improves OOD detection accuracy but also enhances robustness in multi-label, semantically rich, and highly dynamic deployment scenarios.
-Furthermore, the interplay between proxy-based, consistency-based, and low-rank approximation methods suggests promising directions for unified frameworks that leverage the strengths of each paradigm.
+In contrast, the MCM score can be seamlessly integrated into diverse CLIP-based frameworks.
+Recent zero-shot OOD detection methods mainly differ in negative sampling, noise handling, and semantic grouping.
+AdaNeg (Zhang and Zhang, 2024) and AdaND (Cao et al., 2025) refine negative sampling and noise separation, while SimLabel (Zou et al., 2025) focuses on automatic semantic grouping.
+For outlier representation, OLE (Ding and Pang, 2024) expands outlier label diversity through clustering and interpolation, and prompt learning (Jung et al., 2024) introduces new scoring functions for post-hoc detection without retraining.
+Matrix-based methods further diversify the field, with SeTAR (Li et al., 2024b) applying low-rank approximations and OT-DETECTOR (Liu et al., 2025a) using optimal transport to capture semantic gaps.
 LLM-based.
@@
 However, a significant challenge in employing LLMs for OOD detection is their tendency to generate hallucinations—false or misleading information—which can undermine the reliability of OOD identification.
-Recent works have sought to address these challenges by exploring the nuanced relationship between visual and semantic spaces.
-For instance, existing methods (Huang et al., 2024; Dai et al., 2023) observe that visually similar image categories may possess distinct properties in the semantic space, which can be exploited for more effective OOD detection.
+Recent works explore the relationship between visual and semantic spaces to improve OOD detection.
+Studies (Huang et al., 2024; Dai et al., 2023) leverage the distinct properties of visually similar categories in semantic space.
 Dai et al.
-(2023) introduce a consistency-based calibration method that incorporates an object detector to mitigate hallucinations, thereby leveraging the world knowledge of LLMs for more accurate ID class descriptions.
-In contrast, Huang et al.
-(2024) propose ODPC, which generates peer classes based on both ID labels and OOD samples across textual and visual modalities.
-Notably, ODPC requires ID images for training an MLP, indicating that it is not a zero-shot approach.
-In terms of OOD scoring, Dai et al.
-(2023) adopt the MSP, while ODPC utilizes a KNN-based scoring strategy.
+(2023) develops a consistency-based method with object detection to enhance LLM-based class descriptions, while Huang et al.
+(2024) proposes ODPC to generate peer classes using both ID labels and OOD samples.
+These approaches differ in their scoring mechanisms, with Dai et al.
+(2023) using MSP and ODPC employing KNN-based scoring.
 Building upon these foundations, subsequent studies have systematically investigated the behavior of LLMs in OOD detection.
@@
 (2024a) analyze the propensity of LLMs to detect near-OOD versus far-OOD samples, the impact of different fine-tuning strategies, and the suitability of various OOD scoring functions.
-Their findings reveal that LLMs are naturally adept at detecting far-OOD samples, that generative fine-tuning outperforms discriminative fine-tuning, and that cosine distance is an effective OOD score due to the anisotropic nature of LLM embedding spaces.
 More recently, LLM-based approaches such as CMA (Lee et al., 2025), COOD (Liu et al., 2024c), and EOE (Cao et al., 2024) have extended OOD detection into multi-label and concept-rich regimes.
@@
 As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels.
-However,  Fort et al.
-(2021); Ding and Pang (2024) argue that in deployment, labels for OOD are either readily available.
-They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance.
-However, access to OOD labels is rare in real-life scenarios and seems less reasonable.
+While some works (Fort et al., 2021; Ding and Pang, 2024) improve performance by using known OOD class names as candidate labels, this approach relies on the often unrealistic assumption that OOD labels are readily available.
+This setting is also distinct from recent unsupervised fine-tuning approaches (Liang et al., 2024), which address a more realistic scenario where the unlabeled data for adaptation is possibly contaminated with OOD samples.
 Table 3.
@@
 Studies.
-The direct way of adapting a large pre-trained model with several ID samples is fine-tuning.
-Ming and Li (2024) and  Dong et al.
-(2023) conduct studies on the impact of fine-tuning on OOD detection within the context of VLMs, with a special focus on CLIP.
-Ming and Li (2024) pay more attention to the role of parameter-efficient fine-tuning (PEFT) methods and the OOD score.
-The MCM score, introduced by  Ming et al.
-(2022b) as an innovative measure, alongside methods based on prompt learning for detecting OOD instances, is recognized for its effectiveness.
-Similarly,  Dong et al.
-(2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  Ming and Li (2024).
-Following this,  Kim et al.
-(2024) believe that Finetune Like You Pretrain (FLYP), a fine-tuning method, deserves attention due to its good performance on classification tasks.
-Specifically, FLYP is to mimic contrastive language-image pre-training as CLIP.
-After comparing the performance of FLYP and PEFT methods on zero-shot OOD detection,  Kim et al.
-(2024) find FLYP yields better OOD detection performance than PEFT.
-It should be noted that while  Fort et al.
-(2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
+Fine-tuning large pre-trained models with limited ID samples is a common adaptation strategy.
+Recent studies (Ming and Li, 2024; Dong et al., 2023) systematically evaluate fine-tuning and parameter-efficient fine-tuning (PEFT) methods for OOD detection in VLMs, particularly CLIP.
+Both works highlight the superior OOD detection performance of PEFT over traditional fine-tuning.
+The MCM score (Ming et al., 2022b) and prompt-based approaches are also shown to be effective.
+FLYP (Kim et al., 2024), which mimics CLIP-style pretraining, further outperforms PEFT in zero-shot OOD detection.
+Notably, few-shot OOD detection using outlier examples (Fort et al., 2021) is not included here, as it does not align with the focus on ID samples.
 Few-shot OOD Detection.
@@
 In addition, MMFT (Kim and Hwang, 2025) projects image and text features into a shared hyperspherical space and introduces a cross-modal alignment loss function to promote the alignment of image and text representations in the hyperspherical space.
-During training, the loss function brings the image and text embeddings closer together while effectively distinguishing OOD texts from ID texts.
 However, these prompt-only methods may underutilize the rich information contained in image features.
 Recent advances such as GalLoP (Lafon et al., 2024) and Local-Prompt (Zeng et al., 2025) address this limitation by jointly optimizing global and local prompts.
-GalLoP leverages both global and region-level features via prompt dropout and region-level regularization to improve generalization and robustness, whereas Local-Prompt introduces orthogonal local prompt optimization to better exploit limited image data.
-Beyond prompt learning, DSFG (Dong et al., 2023) highlights that conventional fine-tuning can erode critical OOD knowledge.
-To mitigate this, DSFG merges original and fine-tuned features before classifier training, thereby retaining broad OOD awareness and offering plug-and-play compatibility with other fine-tuning-based methods.
-Recently, SCT (Yu et al., 2024a) proposes a self-calibrated tuning framework that dynamically balances ID classification and OOD regularization by introducing modulation factors based on sample uncertainty, enabling more reliable adaptation with only a few ID samples.
-Complementary to this, GaCoOp (Tong et al., 2025) addresses the gradient conflict between ID classification and OOD regularization in prompt optimization.
-By decomposing and projecting gradients, GaCoOp aligns the optimization directions, thereby mitigating the adverse effects of background bias and improving few-shot OOD detection performance, especially when data is scarce.
-Despite these advancements, overfitting remains a challenge due to the scarcity of labeled data.
-To tackle this, Dual-Adapter (Chen et al., 2024c) proposes a training-free few-shot method that constructs positive and negative adapters in a plug-and-play manner, requiring only a handful of ID samples.
-Meanwhile, to address the intrinsic modality gap in VLMs, SUPREME (Wang et al., 2025a) introduces multi-modal prototypes and alignment modules, while CLIP-OS (Sun et al., 2024) regularizes the ID/OOD boundary by synthesizing OOD data through mixing ID features, achieving strong performance without external OOD supervision.
-Additionally, NegPrompt (Li et al., 2024c) further refines the boundary by optimizing negative prompts, demonstrating the continued evolution and diversification of few-shot OOD detection strategies.
+Beyond prompt learning, recent methods address the limitations of conventional fine-tuning, such as the loss of OOD awareness, by merging original and adapted features (Dong et al., 2023) or introducing self-calibrated tuning frameworks that balance ID and OOD objectives (Yu et al., 2024a).
+Other approaches improve prompt optimization through gradient-based techniques (Tong et al., 2025).
+To mitigate overfitting under limited supervision, strategies such as plug-and-play adapters (Chen et al., 2024c), multi-modal alignment (Wang et al., 2025a), and boundary regularization with synthesized OOD features (Sun et al., 2024) have been proposed.
+Further refinements include optimizing negative prompts to better distinguish ID/OOD boundaries (Li et al., 2024c).
+Collectively, these methods enhance few-shot OOD detection by improving adaptation, alignment, and boundary modeling.
 Meta-learning based.
-Meta-learning aims to devise a learning approach that enables rapid adaptation to new challenges  (Vilalta and Drissi, 2002; Hospedales et al., 2021).
+Meta-learning aims to devise a learning approach that enables rapid adaptation to new challenges  (Hospedales et al., 2021).
 OOD-MAML  (Jeong and Kim, 2020) adapt model-agnostic meta-learning (MAML) for few-shot OOD detection.
@@
 It assesses the balance between precision and recall, with a higher AUPR value indicating superior model performance.
-Therefore, a greater AUPR value is desirable.
 FPR@95 (False Positive Rate at 95% True Positive Rate).
@@
 Therefore, a lower value is desirable.
-These metrics provide comprehensive insights into the OOD detection performance, considering different aspects such as the ability to distinguish between ID and OOD samples, handling imbalanced class sensitivity towards ID samples, and controlling the false positive rate at a specific true positive rate threshold.
 6.2. Experimental Protocols
@@
 (2023b); Bai et al.
-(2023); Yu et al.
-(2024b) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously.
+(2023) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously.
 This is crucial to prevent the loss of model generalization.
@@
 To promote further research and enable standardized evaluation, commonly used benchmark datasets are systematically summarized in Table 4.
-To provide a comprehensive overview of recent advances in OOD detection, we first organize our discussion based on fundamental technical domains (Computer Vision, Natural Language Processing, and Beyond) as these domains have distinct methodological characteristics and theoretical foundations that shape how OOD detection is approached.
-This domain-based categorization allows us to systematically analyze the core technical innovations and challenges specific to each field.
-We then extend our analysis to an application-oriented perspective, as shown in Table 6.3, which bridges the gap between theoretical advances and practical implementations, illustrating how these domain-specific techniques are adapted and deployed in real-world scenarios.
+To comprehensively review recent advances in OOD detection, we first categorize our discussion by key technical domains (Computer Vision, Natural Language Processing, and Beyond), as each has distinct methods and theoretical bases shaping OOD approaches.
+This structure enables a systematic analysis of core innovations and challenges in each field.
+We then adopt an application-oriented perspective (see Table 6.3), highlighting how these techniques are adapted and implemented in real-world scenarios.
 Table 5. Representative applications of OOD detection across five major domains, highlighting key use cases in industrial, medical, safety, scientific, and general-purpose settings.
@@
 Industrial Applications Autonomous Driving Detecting unseen obstacles or traffic scenarios to ensure driving safety in autonomous vehicles. (tes, )
-3D Object Detection Identifying out-of-distribution objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
+3D Object Detection Identifying OOD objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
 Object Detection Finding anomalous or unknown items in industrial production lines. (Syed et al., 2024; Chen et al., 2024d; Zhang et al., 2025b)
@@
 Time Series Analysis Monitoring equipment by identifying abnormal patterns for predictive maintenance. (web, )
-Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Oh et al., 2024; Pokhrel et al., 2024; Smedsrud et al., 2021)
+Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Pokhrel et al., 2024; Smedsrud et al., 2021)
 Biological Classification Identifying previously unseen species and enhancing biodiversity monitoring. (Arslan et al., 2024)
-Security and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
+Securit and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
 Facial Expression Recognition Identifying unknown or abnormal facial expressions in surveillance. (Zhang et al., 2024b)
 Human Action Recognition Detecting anomalous or dangerous actions in monitored environments. (Sim et al., 2023; Mandal et al., 2019; Xu et al., 2024b)
-Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021; Gong et al., 2022)
-Scientific Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
-Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Giger and Csillaghy, 2024; Das et al., 2024)
-Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b; Jiang et al., 2025)
-Mathematical Reasoning Identifying out-of-distribution problem types in automated reasoning systems. (Wang et al., 2024b)
+Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021)
+Scientifi Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
+Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Das et al., 2024)
+Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b)
+Mathematical Reasoning Identifying OOD problem types in automated reasoning systems. (Wang et al., 2024b)
 Spiking Neural Networks Recognizing abnormal neural signals in bio-inspired computing models. (Terres-Escudero et al., ; Orchard et al., 2015)
-General-Purpose Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
+General-Purpos Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
 Semantic Segmentation Identifying unknown regions or objects in pixel-level segmentation tasks. (Gao et al., 2023a; Lis et al., 2019)
@@
   • Image Classification.
-The majority of OOD detection research targets image classification, utilizing ID datasets such as MNIST (Xiao et al., 2017), CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009).
-Common OOD datasets include iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017), and Textures (Cimpoi et al., 2014).
-Full-spectrum OOD evaluation often employs the DIGITS, OBJECTS, and COVID benchmarks (Yang et al., 2023b).
-In addition to the aforementioned methods, most current research classifies OOD datasets based on the difficulty of distinguishing them from ID datasets.
-The most popular approach is to use ImageNet-1k as the ID dataset, with SSB-hard (Vaze et al., 2022) and NINCO (Bitterwolf et al., 2023) categorized as near-OOD (representing categories that are difficult to distinguish), and iNaturalist (Van Horn et al., 2018), Textures (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022c) categorized as far-OOD (representing categories that are easier to distinguish
+Image classification represents the primary testbed for OOD detection research.
+Common ID datasets include MNIST (Xiao et al., 2017), CIFAR (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009), while OOD evaluation typically uses datasets like iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), and Textures (Cimpoi et al., 2014).
+Recent research categorizes OOD datasets based on their detection difficulty: near-OOD (e.g., SSB-hard (Vaze et al., 2022), NINCO (Bitterwolf et al., 2023)) for harder-to-distinguish categories, and far-OOD (e.g., iNaturalist, OpenImage-O (Wang et al., 2022c)) for more distinct categories.
   • Semantic Segmentation.
@@
     Object Detection.
-OOD detection in object detection is relatively new (Du et al., 2022b), with evaluations on datasets like PASCAL-VOC (Everingham et al., 2015) and Berkeley DeepDrive-100K (Yu et al., 2018).
-Novel approaches include Situation Monitor, which uses diversity-driven budding ensemble architectures for zero-shot OOD detection in Transformer-based models (Syed et al., 2024), Proto-OOD, which enhances OOD object detection with prototype feature similarity (Chen et al., 2024d), and VisTa, which employs visual-contextual and text-augmented methods for zero-shot object-level OOD detection (Zhang et al., 2025b).
+OOD detection for object detection tasks is a relatively recent research area (Du et al., 2022b), with evaluations commonly conducted on datasets such as PASCAL-VOC (Everingham et al., 2015) and BDD-100K (Yu et al., 2018).
+Recent methods explore diverse strategies, including ensemble architectures (Syed et al., 2024), prototype-based similarity (Chen et al., 2024d), and visual-contextual augmentation (Zhang et al., 2025b).
   • 3D Object Detection.
@@
 OOD detection is vital in medical imaging, using datasets like CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021) depending on the image category (Tan et al., 2024).
-Recent studies focus on reliable OOD detection methods for digital pathology (Oh et al., 2024) and test-time augmentation for improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024).
+Recent studies focus on reliable OOD detection methods for digital pathologyfor improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024).
 In practice, IDx-DR (Dis, ), the first FDA-approved autonomous AI diagnostic system, implements OOD detection by rejecting images that fail to meet quality standards or contain anomalies, deferring these cases to human experts for assessment.
@@
 By leveraging advanced techniques such as attention-based debiasing (Sim et al., 2023), generative feature synthesis (Mandal et al., 2019) and energy-based skeleton modeling (Xu et al., 2024b), recent research has significantly improved OOD detection performance.
-These methods are evaluated on benchmark datasets like UCF-101 (Soomro et al., 2012), HMDB-51 (Kuehne et al., 2011) and Kinetics-400 (Kay et al., 2017).
   • Solar Image Analysis.
 In space weather forecasting, OOD detection serves as a crucial tool for identifying solar anomalies.
-Recent unsupervised  (Giger and Csillaghy, 2024; Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.
+Recent unsupervised  (Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.
   • Facial Expression Recognition.
@@
   • Intent Detection.
-Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022; Akbari et al., 2025).
-Common datasets for evaluation include CLINC150 (Larson et al., 2019), Banking77 (Casanueva et al., 2020), and HWU64 (Liu et al., 2020b).
+Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022).
+Common datasets for evaluation include CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2020b).
 These datasets are specifically designed to evaluate OOD detection in intent classification tasks.
@@
 In text classification tasks, OOD detection helps identify texts that do not belong to any of the known categories.
-Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for in-distribution data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset.
+Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for ID data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset.
 Methods like VI-OOD (Zhan et al., 2024) enhance detection by optimizing joint distributions.
@@
 Question-answering systems require robust mechanisms to handle diverse user queries while maintaining response accuracy.
-The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches.
-Using standard in-distribution datasets (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), PubMedQA (Jin et al., 2019)) and OOD datasets (SQUAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2019)), recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
+The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches.Recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
   • Mathematical Reasoning.
 Mathematical problem-solving systems face challenges when encountering problem types not seen during training.
-Using MultiArith (Roy et al., 2016) as the primary in-distribution dataset, researchers evaluate OOD detection against diverse benchmarks including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), SVAMP (Patil et al., 2021), AddSub (Hill et al., 2015), SingleEq (Wang et al., 2024b), and SingleOp (Wang et al., 2024b).
+Using MultiArith (Roy et al., 2016) as the primary ID dataset, researchers evaluate OOD detection against diverse benchmarks.
 Recent advances, such as TV-Score (Wang et al., 2024b), analyze embedding trajectory variations to enhance OOD detection performance.
@@
   • Audio Recognition.
-In the field of audio recognition, OOD detection is essential for identifying inputs such as synthesized speech and dialects that diverge from the training distribution.
-Traditional probabilistic scoring methods often underperform on atypical audio samples (Lane et al., 2006), motivating the adoption of alternative approaches including confidence-based topic classification (Lane et al., 2006), neural sentence embeddings (Ryu et al., 2017), and deep nearest neighbor analysis (Bukhsh and Saeed, 2023).
-Autoencoder architectures built upon WavLM (Du et al., 2024b) have demonstrated improved performance, particularly in detecting synthetic speech.
-Standard datasets for evaluating OOD detection in this domain include MSCW (Mazumder et al., 2021) and Vocalsound (Gong et al., 2022), which provide both in-distribution and out-of-distribution samples for comprehensive benchmarking.
+OOD detection in audio recognition has evolved from traditional probabilistic scoring (Lane et al., 2006) to methods leveraging confidence-based classification (Lane et al., 2006), neural embeddings (Ryu et al., 2017), and deep nearest neighbor approaches (Bukhsh and Saeed, 2023).
+Recent advances include autoencoder models based on WavLM (Du et al., 2024b), which are particularly effective for synthetic speech detection.
+Standard benchmarks such as MSCW (Mazumder et al., 2021) support comprehensive evaluation in this domain.
   • Graph Neural Networks.
-Within graph neural networks (GNNs), OOD detection encompasses both node-level and graph-level tasks.
-Commonly explored techniques include uncertainty-based graph classification (Bazhenov et al., 2022), energy-based detection (Wu et al., 2023b; Jiang et al., 2025), and semi-supervised learning (Huang et al., 2022b).
-The HGOE framework (Junwei et al., 2024) enhances graph-level detection by incorporating anomalous samples during training, while PGR-MOOD (Shen et al., 2024) utilizes diffusion models to identify out-of-distribution molecular structures.
-GOLD (Wang et al., 2025b) generates pseudo-OOD data with significant differences from the ID embeddings by using a latent generative model to mimic the embeddings of training data and training it jointly with a GNN encoder and an OOD detector.Benchmarking typically relies on paired datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020), which facilitate systematic evaluation of OOD detection methods in GNNs.
+In GNNs, OOD detection encompasses node- and graph-level tasks, utilizing uncertainty estimation (Bazhenov et al., 2022), energy-based methods (Wu et al., 2023b), and semi-supervised learning (Huang et al., 2022b).
+Frameworks like HGOE (Junwei et al., 2024) and PGR-MOOD (Shen et al., 2024) introduce anomalous or pseudo-OOD samples to enhance detection.
+GOLD (Wang et al., 2025b) further advances this area by generating pseudo-OOD data through latent generative modeling.
+Evaluation commonly relies on datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020).
   • Reinforcement Learning.
-OOD detection plays a crucial role in reinforcement learning (RL) by enhancing system robustness when encountering unprecedented states or actions (Leonardos et al., 2023).
-Common approaches include uncertainty quantification through dropout techniques and bootstrapping methods (Sedlmeier et al., 2019), as well as systematic environment modifications (Mohammed and Valdenegro-Toro, 2021).
-In the context of offline RL, advanced methods such as ADAC (Chen et al., 2025b) systematically assess OOD actions to effectively address distribution shifts.
-Comprehensive evaluation frameworks for OOD detection in RL primarily concentrate on various forms of environmental perturbations (Mohammed and Valdenegro-Toro, 2021), providing standardized testing scenarios for robustness assessment.
+OOD detection is vital in reinforcement learning for improving robustness to novel states and actions (Leonardos et al., 2023).
+Approaches typically involve uncertainty quantification (Sedlmeier et al., 2019), environment modifications (Mohammed and Valdenegro-Toro, 2021), and offline RL techniques such as ADAC (Chen et al., 2025b).
+Evaluation frameworks often focus on environmental perturbations (Mohammed and Valdenegro-Toro, 2021) to benchmark robustness.
   • Time Series Analysis.
-In the domain of time series analysis, the inherent temporal dependencies present significant challenges for conventional OOD detection approaches.
-SRS (Belkhouja et al., 2023) addresses this by leveraging time series decomposition in conjunction with deep generative models to estimate conditional likelihoods, while CODiT (Kaur et al., 2023) introduces temporal equivariance to better capture sequential patterns.
-Recent studies have further evaluated modality-agnostic algorithms on multivariate time series datasets, outlining persistent challenges and future research directions (Gungor et al., 2025).
-In practice, OOD detection has been widely adopted in fraud detection services, such as the AWS Fraud Detector (web, ), which utilizes machine learning to identify transaction patterns that deviate from historical norms, effectively flagging potentially fraudulent activities in real time.
+Time series OOD detection faces unique challenges due to temporal dependencies.
+Recent research explores modality-agnostic approaches for multivariate time series (Gungor et al., 2025), with practical applications in fraud detection systems like AWS Fraud Detector (web, ).
   • Biological Classification.
-In biological classification, OOD detection plays a vital role in identifying previously unseen species or anomalies that extend beyond the categories present in training data.
-The integration of DNA barcodes has been shown to enhance image-based classification models through re-ranking strategies, thereby improving robustness and detection accuracy (Arslan et al., 2024).
-  • Spiking Neural Networks.
-Spiking Neural Networks (SNNs), recognized for their event-driven efficiency, present distinctive challenges for OOD detection due to their sparse latent representations.
-Recent methods have proposed combining feedforward learning with distance-based scoring and gradient-free attribution to address these challenges (Terres-Escudero et al., ).
+OOD detection identifies unseen species in biological classification.
+DNA barcodes enhance image-based classification through re-ranking, improving detection accuracy (Arslan et al., 2024).
+  • Spiking Neural Networks (SNN).
+SNNs’ sparse representations challenge OOD detection.
+Recent methods combine feedforward learning with distance-based scoring (Terres-Escudero et al., ).
 The N-MNIST dataset (Orchard et al., 2015) is commonly used for evaluation, with OOD samples generated through diverse event patterns.
-  • Astronomical Imaging.
-In astronomical imaging, OOD detection is essential for distinguishing between simulated and real galaxy images, where conventional statistical techniques often prove inadequate.
-Approaches such as Bayesian model comparison and sparse variational autoencoders (Zhou et al., 2024) have been utilized to uncover subtle discrepancies.
-The SDSS dataset is widely adopted for benchmarking, with simulation-induced biases serving as OOD samples.
+  • Astronomical Imaging. OOD detection distinguishes between simulated and real galaxy images. Bayesian comparison and sparse variational autoencoders (Zhou et al., 2024) uncover subtle differences.
   • Edge Computing.
-Within edge computing, efficient and privacy-preserving OOD detection is critical for devices with limited computational resources.
-SecDOOD (Li et al., 2025) addresses these constraints by employing a collaborative cloud-device architecture and HyperNetwork-based parameter generation, thereby eliminating the need for device-side backpropagation.
-CIFAR-10 and Tiny ImageNet are standard datasets for in-distribution and OOD evaluation in this context.
+Edge computing requires efficient, privacy-preserving OOD detection.
+SecDOOD (Li et al., 2025) uses cloud-device architecture and HyperNetwork generation to avoid device-side backpropagation.
+CIFAR-10 and Tiny ImageNet serve as evaluation datasets.
+The deployment of OOD detection in high-stakes applications introduces considerations beyond predictive accuracy, spanning regulatory compliance, operational safety, and algorithmic fairness.
+For compliance, OOD detection supports safety standards like ISO 21448111https://www.iso.org/standard/77490.html by identifying OOD scenarios, informing release decisions and enabling in-operation validation via shadow-mode monitoring (Henriksson et al., 2023).
+For safety, approaches (Henriksson et al., 2023; Lin et al., 2025) are advancing from runtime input flagging (Henriksson et al., 2023) toward monitoring system-level properties for predictive safety assurance with formal guarantees (Lin et al., 2025).
+For fairness, mitigating bias against minority subgroups involves developing fairness-aware algorithms (Shekhar et al., 2021) and constructing unbiased benchmarks that control for semantic ambiguity (Inkawhich et al., 2022).
 7. Emerging Trends And Open Challenges
@@
 Meta-Learning Adaptation.
-Rapid adaptation to novel data distributions at test time remains a significant challenge in OOD detection.
-Meta-learning algorithms, which follow a learning-to-learn paradigm, have emerged as promising solutions for enabling efficient adaptation to unseen test data.
-Additionally, the vast sample space of potential outliers in training-driven OOD detection motivates the development of improved sampling strategies for more effective utilization of outlier information (Ming et al., 2022a).
+Addressing the challenge of adapting to new data distributions at test time, meta-learning has gained attention for its ability to facilitate efficient OOD detection in novel scenarios.
+Recent work also highlights the importance of developing better sampling strategies to more effectively leverage outlier information during training (Ming et al., 2022a).
 Model Selection.
-Selecting the most suitable OOD detection approach for a given task is non-trivial, as detection performance often varies across data distributions.
-Automated model selection frameworks, such as MetaOOD (Qin et al., 2025), leverage meta-learning to facilitate zero-shot model selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) utilize proportion estimation and likelihood ratio tests to enhance ensemble and score fusion strategies.
-Interestingly, recent work has shown that partially trained, immature models can exhibit effective OOD detection capabilities (Montazeran and Köthe, 2025), suggesting new directions for leveraging model dynamics in detection.
+Choosing an appropriate OOD detection method for a specific task is challenging, as performance varies across data distributions.
+Automated frameworks like MetaOOD (Qin et al., 2025) use meta-learning for zero-shot selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) apply proportion estimation and likelihood ratio tests to improve ensemble and score fusion.
+Notably, recent studies show that partially trained models can also be effective for OOD detection (Montazeran and Köthe, 2025), highlighting new possibilities for leveraging model dynamics.
 Active Learning.
-Active learning strategies, which prioritize labeling the most informative data points, offer efficiency gains in data-scarce settings.
-In OOD detection, active learning can focus annotation efforts on ambiguous or high-uncertainty samples, thereby improving the model’s discrimination between in-distribution and out-of-distribution data.
-SISOM (Schmidt et al., 2024) exemplifies a unified approach that combines active learning and OOD detection by selecting samples based on feature space distances.
+Active learning prioritizes labeling informative data points for efficiency in data-scarce settings.
+In OOD detection, it focuses on ambiguous or uncertain samples to improve ID and OOD discrimination.
+SISOM (Schmidt et al., 2024) combines active learning and OOD detection using feature space distances.
 Continual and Incremental Learning.
-As real-world data is often encountered in a streaming and evolving manner, continual and incremental learning paradigms are increasingly relevant for OOD detection.
-These approaches enable models to adapt to new tasks and data distributions over time while retaining knowledge of previous ones.
-Recent efforts include continual unsupervised OOD detection (Doorenbos et al., 2024), systematic evaluation benchmarks like OpenCIL (Miao et al., 2024b), and hierarchical two-sample tests for continual OOD detection (Liu et al., 2025d).
+Continual and incremental learning enable OOD detection models to adapt to streaming data while preserving previous knowledge.
+Recent work includes unsupervised OOD detection (Doorenbos et al., 2024), OpenCIL benchmark (Miao et al., 2024b), and hierarchical two-sample tests (Liu et al., 2025d).
 Theoretically-Driven Score Design.
@@
 Federated Environments.
-In practical scenarios, data is often distributed across multiple clients with heterogeneous distributions.
-Federated learning (FL) enables collaborative model training while preserving privacy, but also brings unique OOD detection challenges due to covariate and semantic shifts among clients.
-Addressing these challenges is crucial for robust federated models.
-For instance, FOOGD (Liao et al., 2024) combines score matching and feature invariance to handle simultaneous OOD shifts in federated settings.
+Federated learning involves training models across decentralized clients, each with its own data distribution (Guo et al., ; Kuang et al., 2024).
+This distributional variability presents unique OOD detection challenges.
+Methods like FOOGD (Liao et al., 2024) address these issues by enhancing robustness to both covariate and semantic shifts in federated settings.
 Multi-Modal Detection.
-The proliferation of multi-modal data—ranging from images and text to audio and video—demands OOD detection methods capable of handling diverse inputs.
-Benchmarks such as MultiOOD (Dong et al., 2024) provide comprehensive testbeds for multi-modal OOD detection, introducing algorithms like A2D and NP-Mix for challenging scenarios.
-Further advancements, including DIAEF (Gao et al., 2024) and DPU (Li et al., 2024d), focus on cross-modal alignment and dynamic prototype updating, respectively.
-The integration of large language models (LLMs) into multi-modal OOD detection is poised to further enhance the field by combining advanced linguistic reasoning with other modalities.
+With the increasing prevalence of multi-modal data, OOD detection has expanded beyond single modalities.
+MultiOOD (Dong et al., 2024) established a comprehensive benchmark with algorithms like A2D and NP-Mix.
+Recent works (Gao et al., 2024) have advanced cross-modal alignment techniques.
 Multi-Label OOD Detection.
@@
 Detecting OOD samples in these scenarios is challenging due to increased dimensionality and label correlations.
-Recent approaches (Mei et al., 2024; Aguilar et al., 2025) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
+Recent approaches (Mei et al., 2024) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
 Open-Vocabulary Scenario.
-Traditional approaches often assume access to all in-distribution categories, an assumption that does not hold in open-vocabulary contexts.
-Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of in-distribution labels is available, thus paving the way for more generalizable models.
+Traditional approaches often assume access to all ID categories, an assumption that does not hold in open-vocabulary contexts.
+Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of ID labels is available, thus paving the way for more generalizable models.
 Noisy Label Environments.
@@
 Investigations into noisy label scenarios (Humblot-Renaux et al., 2024) offer insights and practical recommendations for improving OOD detection robustness under such conditions.
-Long-Sequence Processing.
-In NLP, the detection of OOD inputs becomes increasingly complex when language models process sequences that exceed their pre-trained context window.
-In such scenarios, models frequently encounter unfamiliar dependency structures, increased attention demands, and altered positional encoding behaviors, all of which deviate from patterns seen during training.
-These distributional shifts impede the model’s generalization capabilities, highlighting the importance of thoroughly understanding their effects.
-Advancing OOD detection in this context is essential for ensuring reliable model performance on sequences that extend beyond standard context boundaries (Liu et al., 2025c).
 7.3. New applications of OOD detection
@@
 Human-in-the-Loop Systems.
-The integration of human expertise into OOD detection is emerging as a critical direction, especially in high-stakes or safety-critical applications.
-Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) incorporate human feedback to refine detection thresholds and guide annotation, balancing algorithmic precision with human intuition.
-The AHA framework, for instance, strategically targets ambiguous regions for selective annotation, improving generalization and error correction under limited supervision.
+Human expertise integration in OOD detection is crucial for safety-critical applications.
+Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) leverage human feedback to optimize detection thresholds and guide annotation.
 Web Image Scraping.
 OOD detection is also finding novel applications in automating web image scraping.
-For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as in-distribution if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
+For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as ID if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
 8. Conclusion
@@
 9. Acknowledgements
-We sincerely appreciate the constructive comments and suggestions provided by all the reviewers and the editor, which have greatly improved the quality of our work.
-We are truly grateful for their valuable time and insightful feedback.
+We thank the editor and reviewers for their constructive feedback, and Zhen Jia and Yongcan Yu for their assistance with the manuscript.
+This work was supported by the National Natural Science Foundation of China (62276256, U2441251), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Young Scientists Fund of the State Key Laboratory of Multimodal Artificial Intelligence Systems (ES2P100117).